### v1

# 2024_subroc.bib

Abstract
Machine learning (ML) is increasingly employed in real-world applications like medicine or economics, thus, potentially affecting large populations. However, ML models often do not perform homogeneously across such populations resulting in subgroups of the population (e.g., sex=female $\wedge$ marital_status=married) where the model underperforms or, conversely, is particularly accurate. Identifying and describing such subgroups can support practical decisions on which subpopulation a model is safe to deploy or where more training data is required. The potential of identifying and analyzing such subgroups has been recognized, however, an efficient and coherent framework for effective search is missing. Consequently, we introduce SubROC, an open-source, easy-to-use framework based on Exceptional Model Mining for reliably and efficiently finding strengths and weaknesses of classification models in the form of interpretable population subgroups. SubROC incorporates common evaluation measures (ROC and PR AUC), efficient search space pruning for fast exhaustive subgroup search, control for class imbalance, adjustment for redundant patterns, and significance testing. We illustrate the practical benefits of SubROC in case studies as well as in comparative analyses across multiple datasets.
1 Introduction
Machine learning (ML) is widely adopted in high-stake application domains such as health care, loans, hiring decisions, or legal decision-making [mehrabi_survey_2021]. Thus, ML models and their decisions already influence large populations and everyday life. However, it has also repeatedly been shown that ML models can be highly heterogeneous in their performance [mehrabi_survey_2021]. Subgroups of a population may exist where a ML model underperforms (or performs remarkably well). For illustration, see Figure 1 where predictions are particularly bad for the subpopulation of women from rural areas (sex=female $\wedge$ housing=rural) and remarkably accurate for male teenagers (sex=male $\wedge$ $\leq$18). Knowing such subgroups may allow data scientists and practitioners to make informed decisions about model deployment and development. Here, the practitioners may decide to use the current model only for male teenagers, where the model performs well, and the data scientists may decide to collect more data for subgroups where the model underperforms.
Figure 1: In an exemplary application, binary classification could distinguish between healthy and diseased patients (left). Predictions of soft classifiers impose an ordering on the persons, which performs better the clearer healthy patients are separated from diseased patients. In this case there is a clear difference between the densities of healthy and diseased patients in this ordering, which are represented by the red and blue curves here (middle left). A dataset can be split up in different ways by patterns describing subsets. Our approach based on subgroup discovery evaluates these subsets systematically and returns those where the model predictions create the best or worst separation of the classes (middle right). Results may highlight arbitrary sensitive subgroups of a population, where the model under- or over-performs compared to the rest of the dataset (right).
Background and problem setting. The need and potential of identifying and analyzing subgroups has been widely recognized. For example, a wide array of work aims at alleviating consistent underperformance in known [sagawa_distributionally_2020] as well as previously unknown [sohoni_no_2020] subgroups. Similar observations have been made concerning the fairness of ML models [mehrabi_survey_2021]. Consequently, various approaches have been explored to study such atypical subgroups systematically, e.g., with interactive tools for manual discovery of subgroups [baylor_tfx_2017, kahng_visual_2016]. However, this is labor-intensive, inefficient, and poses the threat of missing essential subgroups. Thus, automated approaches have been proposed that enable the search for interpretable subgroups where ML models over- or underperform [chung_automated_2020, cabrera_fairvis_2019]. The most principled approach employs the framework of Exceptional Model Mining (EMM) with a custom performance measure for efficient exhaustive search [duivesteijn_understanding_2014]. However, none of these approaches combine the automated and exhaustive search of EMM, standard measures for soft classifiers like ROC AUC or PR AUC, efficient search, and handling of class imbalance and statistical significance testing, which are required for a routine and practical analysis.
Contribution. Consequently, we introduce SubROC, which builds on EMM to reliably and efficiently find strengths and weaknesses of classification models. For this, we (a) propose a new group of scoring functions based on the area-under-the-curve (AUC) performance statistic, (b) derive tight optimistic estimates that enable fast exhaustive subgroup search, (c) provide a well-rounded framework that includes control for class imbalance, pruning of redundant results, and significance testing, (d) show the practical benefits of our approach across multiple datasets, and (e) integrate SubROC in the easy-to-use, open source pysubgroup package and provide the code for our experiments.  [^1] Experiment code is available at https://github.com/bckrlab/subroc.
Overall, SubROC will enable practitioners to routinely evaluate their ML models by finding subgroups where they over- or underperform. This will not only have the potential to impact and improve quality control procedures and increase the confidence in deployed ML models but also lay the foundation for downstream approaches to systematically improve their performance.
2 Preliminaries
2.1 Soft Classifier Evaluation
SubROC focuses on binary soft classification. $D=(\mathcal{I},\mathcal{A})$ is a dataset with instances $c\in\mathcal{I}$ and attributes (features) $\mathcal{A}=(X_{1},\ldots,X_{m})$. $X_{i}(c)$ denotes the value of attribute $X_{i}$ in instance $c$. $Y(c)\in\{1,0\}$ is the ground truth label and $\hat{Y}(c)\in[0,1]$ a continuous (soft) prediction. Based on the ground truth labels, we call some instances positives ($\mathcal{P}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=1\}$) and some negatives ($\mathcal{N}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=0\}$). A classifier performs well if it returns high (low) values for instances with the ground truth $Y(c)=1$ ($Y(c)=0$).
While many evaluation metrics for soft classification exist, including binary cross-entropy or the Brier score, the area under the curve (AUC) of the Receiver Operating Characteristic (ROC) curve is among the most widely used. It assumes a ranking of instances according to their predicted probabilities from high to low. Then, it computes the probability of a randomly chosen positive instance being ranked above a randomly chosen negative instance.
A main reason for the popularity of AUC-based evaluation metrics is that they summarize across all potential thresholds that a user could apply to conduct a hard classification into two classes based on the soft classifier results, e.g., depending on misclassification costs. This makes the ROC AUC a natural choice for interpretable subgroup descriptions, contributing to the insights that subgroups from our framework may provide to experts.
2.2 Subgroup Discovery and Exceptional Model Mining
A selector $s\in\mathcal{S}$ is a boolean function based on a single attribute, e.g., $sex=female$ or $age\in[20,30)$. A pattern $p\in 2^{\mathcal{S}}$ is a conjunction of selectors, e.g., $sex=female\wedge age\in[20,30)$. In a given dataset, a pattern’s cover $sg(p)$ represents the set of all instances covered by it, i.e., all instances with $p=true$. A pattern and its cover are referred to as a subgroup. We call a subgroup $p_{spec}$ a specialization of another subgroup $p_{gen}$ (its generalization) if its pattern’s conjunction of selectors adds at least one additional selector: $p_{gen}\subset p_{spec}$.
Subgroup discovery (SD) [klosgen_explora_1996] finds subgroups that show an interesting deviation with respect to a target concept, which is traditionally given as a single data attribute. E.g., an interesting subgroup may contain a higher-than-average disease occurrence. EMM [leman_exceptional_2008] extends target concepts to model classes. Here, a subgroup is considered interesting if the model parameters fitted to the subgroup are significantly different from those fitted to the complete dataset.
SD and EMM formalize interestingness using scoring functions $\varphi:2^{\mathcal{S}}\rightarrow\mathbb{R}$ (also called interestingness measures or quality measures) which map a pattern $p\in 2^{\mathcal{S}}$ to a score given a dataset $D$. A search algorithm explores the search space of patterns ($2^{\mathcal{S}}$) to find the ones with the highest scores. The computational challenge of subgroup discovery lies in the exponentially growing search space $2^{\mathcal{S}}$ with the number of selectors $|\mathcal{S}|$. Beyond straightforward constraints like pattern length or a minimum number of covered instances, optimistic estimates $oe_{\varphi}$ are integral for optimizing runtimes. An $oe_{\varphi}$ provides an upper bound on the score $\varphi$ of all specializations $p_{spec}$ of a subgroup $p_{gen}$ [lemmerich_novel_2014]: $\forall p_{spec}\supset p_{gen}:\varphi(p_{spec})\leq oe_{\varphi}(p_{gen})$. For top-k search, this can be used to disregard all specializations of $p_{gen}$ if $oe_{\varphi}(p_{gen})$ is lower than the score of all current top-k patterns. Furthermore, tight optimistic estimates [grosskreutz_tight_2008] provide tighter upper bounds by guaranteeing that the upper bound can be reached by a hypothetical specialization $p_{spec}$, for an arbitrary subset of instances $I\subseteq sg(p_{gen})$ covered by $p_{gen}$. Optimistic estimates can be integrated into various search algorithms. However, they have to be derived for each scoring function individually.
2.3 SCaPE Model Class
For EMM, the Soft Classifier Performance Evaluation (SCaPE) model class has been proposed to discover subgroups with exceptional performance of soft classifiers [duivesteijn_understanding_2014]. It integrates the ground truth labels $Y$ and the model predictions $\hat{Y}$.
The original scoring function Relative Average Subranking Loss (RASL), written $\varphi^{rasl}$ here, proposed for SCaPE, measures the number of violations of the ranking induced by the predictions $\hat{Y}(c)$ of the given soft classifier against the ranking based on the ground truth labels $Y(c)$ within a subgroup $sg$ with pattern $p$: For each positive instance $c\in\mathcal{P}_{p}$ for a pattern $p$, it counts how many negative instances $c^{\prime}\in\mathcal{N}_{p}$ are ranked higher. This count is denoted as $PEN(c,p)$. The Average Ranking Loss (ARL) is then defined as the average of $PEN(c,p)$ across all positive instances. The final scoring function $\varphi^{rasl}$ is defined relative to the overall dataset: $\varphi^{rasl}(p):=ARL(p)-ARL(\mathcal{\emptyset})$.
Although the ARL score is normalized by the number of positive instances, it is heavily influenced by the number of overall instances covered by a pattern as the penalty score $PEN(c,p)$ counts the absolute number of violations. This favors large subgroups or subgroups with heavily biased class distributions without an option to manually trade off these aspects (see Section 4.4 for details).
3 The SubROC Framework
The SubROC framework consists of three components: defining novel scoring functions based on the widely employed performance measures (ROC AUC and Precision-Recall (PR) AUC) while accounting for class imbalance, size, and redundancy (Section 3.1), deriving optimistic estimates for efficient search space exploration and pruning (Section 3.2), and significance filtering for increasing the robustness and generalizability of the discovered subgroups (Section 3.3).
3.1 Scoring Functions
3.1.1 AUC-based subgroup interestingness
For a set of instances $I$, ROCAUC and PRAUC both map soft predictions $\hat{Y}(I)$ and corresponding ground truth labels $Y(I)$ to a performance metric, denoted as $ROCAUC(I)$ and $PRAUC(I)$, respectively. We define the corresponding relative AUC-based subgroup scores given a dataset $D=(\mathcal{I},\mathcal{A})$ and pattern $p$ as:
$\varphi^{rROCAUC}(p)=$ $ROCAUC(\mathcal{I})-ROCAUC(sg(p))$
$\varphi^{rPRAUC}(p)=$ $PRAUC(\mathcal{I})-PRAUC(sg(p))$
To guarantee well-defined scoring, we define constraints: at least one positive instance for ARL and PR AUC and at least one instance of both classes for ROC AUC. For PR AUC, we employ the widely used approximation via a linearly interpolated PR curve, with no practical consequences in our initial tests, cf. [davis_relationship_2006].
Subgroup scoring functions commonly include a factor for trading off size against the score to prevent the discovery of only small, overfitted subgroups (e.g., see [klosgen_explora_1996, leman_exceptional_2008]). We apply the same concept to our AUC-based scoring functions $\varphi^{\prime}$ and derive size-sensitive scoring functions $\varphi_{\alpha}$ as $\varphi_{\alpha}(p)=|sg(p)|^{\alpha}\cdot\varphi^{\prime}(p)$. Here, $\alpha\in\mathbb{R}_{\geq 0}$ is the trade-off parameter, where $\alpha=0$ removes the size influence.
3.1.2 Class Balance Factor
The performance measures ROC AUC and PR AUC are known to be sensitive to the Negative Class Ratio (NCR) [cook_when_2020, boyd_unachievable_2012]. This can lead to subgroups with high scores just due to high or low class imbalances rather than exceptional model performance. To counterbalance this, we introduce a user-weighted class balance factor $cb(I)$ for our scoring functions [canbek_binary_2017, datta_multiobjective_2019]. Given a set of instances $I$, it provides a maximum value close to 1 for perfectly balanced classes and lower values for more imbalanced data. We define $cb(I)$ as 00 if only one class occurs and as $\min\{\frac{|\mathcal{N}_{I}|}{|\mathcal{P}_{I}|},\frac{|\mathcal{P}_{I}|}{|\mathcal{N}_{I}|}\}$ otherwise. We combine $cb(I)$ with a scoring function $\varphi^{\prime}$ as an additional factor, forming a class balance weighted scoring function $\varphi_{\beta}(p)$ for a pattern $p$: $\varphi_{\beta}(p)=cb(sg(p))^{\beta}\cdot\varphi^{\prime}(p)$ with $\beta\in\mathbb{R}_{\geq 0}$. Scoring functions that include both the cover size weight with parameter $\alpha$ and class balance weight with parameter $\beta$ are denoted as $\varphi_{\alpha,\beta}$. This follows the intuition that ROC AUC can be deceptive for imbalanced data, and thus, balanced subgroups should be preferred [saito2015precision]. Alternatives such as the difference in class-ratios between the subgroup and the dataset could also be considered.
3.1.3 Generalization Awareness
It is a common issue in subgroup discovery that multiple subgroups in the result set have a high pairwise overlap in terms of their cover [lemmerich_difference-based_2013]. To reduce the intersections of subgroup covers in result sets, there is an adaptation to scoring functions available from literature, called generalization awareness. For this, a variant [lemmerich_difference-based_2013] of previous approaches[batal_concise_2010, lemmerich_local_2011, grosskreutz_subgroup_2010] extends a scoring function $\varphi^{\prime}$ by reducing the score $\varphi^{\prime}(p)$ of a subgroup with pattern $p$ by subtracting the maximum score across all its generalizations $h$ (making a specialization viable only if its score is substantially higher):
$\varphi_{\alpha}(p)=|sg(p)|^{\alpha}\cdot(\varphi^{\prime}(p)-\max_{h\subset p}\varphi^{\prime}(h))\text{, with }\alpha\in[0,1].$
We employ a variation of this which allows applying it to cover size and class balance weighted scoring functions $\varphi_{\alpha,\beta}^{\prime}$:
$\varphi_{\alpha,\beta}(p)=\varphi_{\alpha,\beta}^{\prime}(p)-\max_{h\subset p}\varphi_{\alpha,\beta}^{\prime}(h)$
3.2 Optimistic Estimates
For runtime optimization, we derive tight optimistic estimates for the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ as well as, separately, for the cover size and class balance factors. Proofs are given in Appendix A.
3.2.1 Average Ranking Loss
To date, no (tight) optimistic estimate was derived for the ARL based scoring function $\varphi^{rasl}$ [duivesteijn_understanding_2014]. The $ARL(p)$ is defined as the average over the ranking losses $PEN(c,p)$ of all covered positive instances $c\in\mathcal{P}_{p}$. The penalty $PEN(c,p)$ for each positive instance $c$ counts the negative instances ranked higher than $c$. Thus, removing positive instances with non-maximal $PEN(c,p)$ does not alter the $PEN$ of the remaining instances and increases the average by eliminating small values. The $ARL(p)$ is therefore maximized when only a single positive $c^{\prime}$ instance with maximal $PEN(c^{\prime},p)$ is left.
Theorem 1 (Tight Optimistic Estimate for $\varphi^{rasl}$).
Given a dataset $D=(\mathcal{I},\mathcal{A})$ and a pattern $p$, a tight optimistic estimate for $\varphi^{rasl}(p)$ is $oe_{\varphi^{rasl}}(p)=\max_{c\in\mathcal{P}_{p}}\{PEN(c,p)\}-ARL(\emptyset)$, with $\mathcal{P}_{p}=\{c\in sg(p)|Y(c)=1\}$.
3.2.2 Area Under the Receiver-Operating Curve
Next, we introduce a lower bound $b$ for ROC AUC that we will use to derive an optimistic estimate of $\varphi^{rROCAUC}$. We distinguish three cases: (a) If perfect separation of ground truth labels is possible with a threshold on the soft predictions, the same holds for every subset. Thus, the ROC AUC is always 1. (b) If the input additionally contains some instances that share the same prediction value (ties) but have different labels (positive and negative), then the ROC AUC on only these tied instances equals $\frac{1}{2}$. Any subset of the given instances containing both classes and at least one non-tied instance allows for the correct separation of some positives and negatives by at least one threshold. Thus, the ROC AUC is greater than $\frac{1}{2}$ for such subsets. (c) The input contains at least one incorrectly ranked pair of a positive and a negative instance. The subset containing exactly one such pair has an ROC AUC of 00, which is the lowest possible ROC AUC.
Lemma 2 (Tight Lower Bound $b_{ROCAUC}$).
For a set of instances $I$ containing at least one positive and one negative instance, a tight lower bound for the ROC AUC of all subsets of $I$ is given by
$b_{ROCAUC}(I)=\begin{cases}1,&\begin{aligned} \text{if }\forall c,c^{\prime}\in I:&(y<y^{\prime}\rightarrow\hat{y}<\hat{y}^{\prime})\wedge\\ &(y>y^{\prime}\rightarrow\hat{y}>\hat{y}^{\prime})\end{aligned}\\ \frac{1}{2},&\begin{aligned} \text{if }\forall c,c^{\prime}\in I:&(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge\\ &(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})\end{aligned}\\ 0,&\text{otherwise}\end{cases}$
with $y$ and $y^{\prime}$ ($\hat{y}$ and $\hat{y^{\prime}}$) denoting $Y(c)$ and $Y(c^{\prime})$ ($\hat{Y}(c)$ and $\hat{Y}(c^{\prime})$), respectively.
Theorem 3 (Tight Optimistic Estimate for $\varphi^{rROCAUC}$).
Given a dataset $D=(\mathcal{I},\mathcal{A})$ and a pattern $p$, a tight optimistic estimate for $\varphi^{rROCAUC}(p)$ is $oe_{\varphi^{rROCAUC}}(p)=ROCAUC(\mathcal{I})-b_{ROCAUC}(sg(p))$.
3.2.3 Area Under the Precision-Recall Curve
Lemma 4 introduces a lower bound for PR AUC which is used in Theorem 5 to derive a tight optimistic estimate for $\varphi^{rPRAUC}$. For Lemma 4, a subset $I_{worst}$ of the original set of instances $I$ is constructed to maximize classification errors while minimizing the positive class ratio. This makes it impossible for a threshold on any other subset to define a point in PR space below the PR curve of $I_{worst}$ (with some exceptions that have no impact on the AUC of the resulting PR curve). Proofs and an illustration are included in Section A.3.
Lemma 4 (Tight Lower Bound $b_{PRAUC}$).
For a set of instances $I$ containing at least one positive instance, a tight lower bound for the PR AUC of all subsets of $I$ is given by $b_{PRAUC}(I)=PRAUC(I_{worst})$, where $I_{worst}=\{\text{argmin}_{c\in\mathcal{P}_{I}}\{\hat{Y}(c)\}\}\cup\mathcal{N}_{I}$.
Theorem 5 (Tight Optimistic Estimate for $\varphi^{rPRAUC}$).
Given a dataset $D=(\mathcal{I},\mathcal{A})$ and a pattern $p$, a tight optimistic estimate for $\varphi^{rPRAUC}(p)$ is $oe_{\varphi^{rPRAUC}}(p)=PRAUC(\mathcal{I})-b_{PRAUC}(sg(p))$.
3.2.4 Cover Size and Class Balance Weighted Scoring Functions
The following Lemma 6 gives a tight upper bound for the weighting used in cover size and class balance weighted scoring functions when $\alpha\leq\beta$. This result translates to optimistic estimates for $\varphi_{\alpha,\beta}^{rasl}(\cdot)$, $\varphi_{\alpha,\beta}^{rROCAUC}(\cdot)$ and $\varphi_{\alpha,\beta}^{rPRAUC}(\cdot)$. It therefore enables the use of the optimistic estimates presented in theorems 1, 3 and 5 when cover size and class balance weighting is enabled.
Lemma 6 (Weighting Tight Upper Bound).
For a set of instances $I$, a tight upper bound of the overall weighting $w(I)=|I|^{\alpha}\cdot cb(I)^{\beta}$ with $\alpha\leq\beta\in\mathbb{R}_{>0}$ is given by $b_{w}(I)=(2\cdot\min\{|\mathcal{P}_{I}|,|\mathcal{N}_{I}|\})^{\alpha}$.
The product $b_{w}(sg(\cdot))\cdot\varphi^{\prime}$ with $\varphi^{\prime}\in\{oe_{\varphi^{rasl}}(\cdot),oe_{\varphi^{rROCAUC}}(\cdot),oe_{\varphi^{rPRAUC}}(\cdot)\}$ is an optimistic estimate of the cover size and class balance weighted scoring functions $\varphi_{\alpha,\beta}^{rasl}(\cdot)$, $\varphi_{\alpha,\beta}^{rROCAUC}(\cdot)$ and $\varphi_{\alpha,\beta}^{rPRAUC}(\cdot)$.
3.3 Significance Filtering
Subgroup discovery inspects an exponential search space of patterns and consequently requires correction for multiple hypothesis comparison. To reduce the number of compared subgroups we use a holdout procedure [webb_discovering_2007] splitting analyzed datasets into two parts, the search dataset to search subgroups and the validation dataset to test statistical significance. The tested statistic is the raw score of a subgroup (written $\varphi_{0,0}$) excluding cover size and class imbalance. The null hypothesis is that $\varphi_{0,0}(p)$ for a pattern $p$ is drawn from the sample distribution of $\varphi_{0,0}$ on random instances with the same size and class balance as the cover $sg(p)$ of $p$. This allows a test on performance differences, excluding the influence of cover size and class balance on the used performance measure.
We compute the empirical p-value of a subgroup $p$ by a randomization test (cf. [gionis_assessing_2007]), where $1000$ random subsets of the validation dataset are drawn without replacement and with fixed size and class balance. $\varphi_{0,0}$ is evaluated on each of the random subsets to construct an empirical distribution of exceptionality. The empirical p-value is the fraction of random subsets with interestingness $\varphi_{0,0}$ greater than or equal to $\varphi_{0,0}(p)$ divided by the total number of random subsets.
Multiple testing correction is done using Benjamini-Yekutieli (BY) [benjamini_control_2001] (controlling the false discovery rate (FDR)). BY has the necessary properties to be applied under any dependency between tested hypotheses. The correction approach can be customized by the user, e.g., with Bonferroni controlling familywise error rate (FWER). The default p-value threshold we use is $p\leq 0.05$.
4 Experiments
Datasets are from OpenML [vanschoren_openml_2013] and the UCI Machine Learning Repository [kelly_uci_nodate] (statistics in Supp. Table 3). Each was partitioned randomly into four equal parts for training, subgroup discovery, significance testing, and result evaluation.
4.1 Quality of SubROC Results Across Multiple Datasets
To illustrate the quality of subgroups discovered with SubROC, we applied xgboost on various datasets and searched for exceptionally underperforming subgroups with regard to the performance measures ROC AUC, PR AUC, and ARL. Table 1 compares two settings: a basic SubROC setting (b) and the full SubROC framework (f). For b, the search solely uses the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$, and $\varphi^{rPRAUC}$. The top-5 subgroups are returned and additionally tested for significance. The full framework (f) is configured with size and class imbalance weighting $\alpha=\beta=1$, generalization awareness, and statistical significance filtering. Here, the search returns the top-100 subgroups, which were filtered for significance, and finally, the top-5 subgroups are returned. In both cases, we used a minimum subgroup size of $\geq 20$, and a maximum pattern length of 4. Statistics were calculated for top-5 results. Note that exceptionality should not be compared between different performance measures because of their differing scales and responses to errors.
For b, cover sizes tend to be close to the minimum cover size limit, and class ratios tend to be close to 0 or 1. Despite their high difference in performance compared to the overall dataset (exceptionality), such subgroups do not necessarily point towards generalizable structural issues of the classifier but may be caused by artifacts (cf. Section 4.4). This is also apparent in the low number of subgroups passing statistical significance in the top-5 results of b. Similar patterns are observable in the top-100 results (see Supp. Table 5). While yielding less prominent exceptionality, we observe that the SubROC framework leads to larger and more balanced results. Additionally, substantially more subgroups are returned that pass the significance filtering step. This is apparent in the top-5 as well as the top-100 result (cf. Supp. Table 5).
We also inspect the pairwise Intersection over Union (IoU) of the discovered top-5 subgroups. A higher IoU refers to more overlapping and, thus, redundant subgroups. Values generally remain in a similar range to the baseline, but we observe occasional significant improvements by SubROC. This indicates that our framework does not come at a cost of and in some cases improves result diversity.
The statistical filtering step of SubROC saves resources for practitioners by discarding a considerable amount of initially found, but unreliable subgroups. This effect is particularly prominent for small datasets with less than $2,500$ search instances, where the significance filtering discarded almost all subgroups. It is less prominent for ARL based scoring functions, which generally produce smaller result sets with a greater share of significant subgroups. Additional results can be found in Section B.2. For concrete result sets from a case study on the Adult dataset see Section B.3.
Table 1: Overview of top-5 result set properties from a basic SubROC setting (b) and the full SubROC framework (f) for a range of datasets and performance measures. SubROC managed to find diverse and exceptional subgroups, which are often much larger and balanced than the baseline. (Continued)
Mean Cover Size
Significant/ Top-5
Top-5/ Significant/ Original
Dataset Metric     b f     b f
Census n=199523 NCR=0.06 ROC     25 6753    1/5 5/99/100
PR     23 864    0/5 5/99/100
ARL     8000 11945    5/5 5/27/27
Adult n=48842 NCR=0.24 ROC     27 6492    0/5 5/77/100
PR     27 2803    2/5 5/70/100
ARL     4614 6930    5/5 5/5/5
Bank n=45211 NCR=0.88 ROC     22 311    0/5 5/22/100
PR     157 72    0/5 2/2/100
ARL     - -    0/0 0/0/0
Credit n=30000 NCR=0.78 ROC     24 1506    0/5 5/38/100
PR     75 1895    0/5 5/17/100
ARL     - -    0/0 0/0/0
Approval n=690 NCR=0.55 ROC     29 -    0/5 0/0/41
PR     41 80    2/5 1/1/25
ARL     60 87    3/5 3/3/3
Table 1: Continued from previous columns. (Rows continued in Supp. Table 4, containing results for more datasets)
Mean Exceptionality
Mean NCR
Mean Pairwise IoU
Dataset Metric     b f     b f     b f
Census n=199523 NCR=0.06 ROC     0.95 0.05    0.04 0.16    0.10 0.31
PR     0.92 0.14    0.93 0.45    0.37 0.45
ARL     47 31    0.15 0.11    0.99 0.50
Adult n=48842 NCR=0.24 ROC     0.92 0.06    0.04 0.37    0.48 0.61
PR     0.94 0.08    0.96 0.46    0.37 0.19
ARL     148 55    0.42 0.34    0.85 0.61
Bank n=45211 NCR=0.88 ROC     0.83 0.16    0.95 0.48    0.10 0.12
PR     0.68 0.10    0.99 0.48    0.80 0.51
ARL     - -    - -    - -
Credit n=30000 NCR=0.78 ROC     0.75 0.12    0.96 0.61    0.01 0.09
PR     0.51 0.29    0.99 0.85    0.18 0.34
ARL     - -    - -    - -
Approval n=690 NCR=0.55 ROC     0.64 -    0.96 -    0.69 -
PR     0.89 0.85    0.98 0.95    0.75 -
ARL     26 10    0.96 0.79    0.68 0.69
4.2 Optimistic Estimates Substantially Speed Up Runtimes
To investigate the efficiency improvements enabled by SubROC, we measure runtimes of subgroup discovery with and without pruning based on the introduced optimistic estimates 10 times for each dataset (3 times for Census due to long runtimes). The searches run on predictions from a Gaussian Naive Bayes classifier, using the best-first-search enumeration strategy, with a maximum pattern length of 4. Different size and class imbalance weightings ($\alpha,\beta$) are compared. Table 2 gives an overview of the measured speedups in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning. Absolute runtimes and experiments with different pattern lengths are reported in Section B.5.
The results show substantial runtime reductions. The most prominent improvements were observed for ARL with speedups up to $538.9$ on the Credit dataset, reducing runtime from $\sim 1.25$ hours to less than 10 seconds. For PR AUC, the highest speedup of $22.6$ was also on Credit with runtime reductions from $\sim 1.25$ hours to $\sim 4$ minutes. Finally, for ROC AUC, a speedup of $101.0$ on Census was the highest with an improvement from $\sim 20$ hours to $\sim 12$ minutes. Overall, larger datasets (both in terms of number of instances and features) benefit the most from optimistic estimates across all performance measures. For the smallest datasets (with generally small runtimes) like Approval, where the search split contains less than 250 instances, the speedups substantially decrease across all measures, leading to an overall slowdown in some cases. With regard to parameterization, a very low emphasis on size and class imbalance compared to the performance measure ($\alpha=\beta=0$) as well as a high emphasis ($\alpha=\beta=1$) seems to produce the highest speedups. Further investigations are required, however, an explanation for this may be that the search can more strongly focus on one or the other optimistic estimate factor
Table 2: Speedups of median runtimes using optimistic estimates. Speedups are given in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning. Substantial runtime reductions were achieved on all but the smallest dataset. (Cont. in Supp. Table 12, containing results for more datasets)
Performance Measure
Weight ($\alpha=\beta$)
Census n/m=199523/41
Adult n/m=48842/14
Bank n/m=45211/16
Credit n/m=30000/23
Approval n/m=690/15
ARL 0    18.4 47.4 9.7 538.9 1.7
0.1    18.6 39.4 7.7 385.6 0.8
0.3    19.9 40.4 7.6 383.1 0.7
1    23.3 44.7 7.6 388.4 0.6
PR AUC 0    2.9 2.3 1.7 2.5 1.2
0.1    1.9 1.2 0.7 1.1 0.4
0.3    2.7 2.1 0.8 2.9 0.4
1    11.4 19.9 2.2 22.6 0.4
ROC AUC 0    101.0 0.9 2.1 83.8 0.8
0.1    1.0 0.8 0.8 1.0 0.5
0.3    1.3 1.0 0.9 1.1 0.5
1    6.4 8.4 2.1 3.9 0.5
4.3 Discovering Injected Subgroups
To investigate the validity of SubROC and its ability to discover expected subgroups, we injected an artificially underperforming subgroup into a dataset based on the predictions of an xgboost classifier on the Adult dataset. The injected subgroup was selected randomly from all subgroups with pattern length $\leq 3$ in a cover size range from $0.4\%$ to $0.6\%$ of the search split. It covers 54 instances, has an NCR of 0.2, and an ARL of 1.07, PR AUC of 0.97, and ROC AUC of 0.90 before and 9.93, 0.62, and 0.1, respectively, after injection. For the injection, we multiplied the predicted scores of instances in the subgroup cover by $-1$. Thereby, the injected subgroup forms a known weakness in the predictions, and we can evaluate the subgroup search by whether this subgroup is included in the result set. The subgroup discovery was performed up to depth 4 under different configurations (baseline without the SubROC framework and three different weightings of cover size and class balance in the SubROC framework). Result sets were top-10 filtered to reflect practically interpretable result set sizes.
Figure 2 shows the IoU as an overlap measure between found subgroups and the injected subgroup for the ROC AUC. In contrast to the baseline, the SubROC framework successfully discovers the injected subgroup and reports it at the highest ranks in the result set. Note that multiple patterns can describe the same set of instances, so the injected errors can be found several times.
While ROC AUC robustly identifies the subgroup, similar experiments on ARL or the PR AUC did not reproduce these findings. The result sets contain only subgroups with IoU s with the injected subgroup below $0.1$. See Section B.4 for detailed results. Differences in terms of cover size and class balance skew (see Section 4.4) between the performance measures have a large influence on the results of these experiments. The injected subgroup contains more negatives than positives, which interacts with the class balance skews of ARL and PR AUC. Furthermore, the injected subgroup is small, making it inherently hard to capture with ARL, which has an embedded size skew. This highlights the importance of the size and class balance terms introduced in SubROC to counter these skews, but also shows that they do not completely remove the effects. Future work, for example, employing a normalized PR AUC [boyd_unachievable_2012] may be promising.
Figure 2: Intersection over Union (IoU) of an injected subgroup with subgroups in a top-10 ROC AUC result set, ordered decreasingly by interestingness. “baseline” was obtained searching in a base SubROC setting. The other results were obtained using the full SubROC framework with corresponding cover size weighting $\alpha$ and class balance weighting $\beta$, successfully discovering the injected subgroup. The full SubROC framework accurately found the injected subgroup for any weighting configuration by returning two highly overlapping subgroups in the highest ranks of all result sets.
4.4 Importance of Cover Size and Class Imbalance Weighting
To illustrate the effect of the interaction between model performance, size and class ratio on the proposed scoring functions, we conduct synthetic experiments (cf. [cook_when_2020]). Data was generated from a bivariate normal distribution with means 00 and variances $1$. The first variable $z$ was used to derive a ground truth label, the second $\hat{z}$ represents the prediction values. Varying their correlation thus varies the “prediction performance” (the $y$-axes). Dataset sizes (cover of a subgroup) are varied in Figures 3(a), 3(b), 3(e), 3(f), 3(i) and 3(j) ($x$ axis) and set to $100$ in Figures 3(c), 3(d), 3(g), 3(h), 3(k) and 3(l). Class balance (NCR) $q$ was varied by thresholding $z$ at the first $q$-quantile in Figures 3(c), 3(d), 3(g), 3(h), 3(k) and 3(l) ($x$ axis) and fixed at $q=0.5$ for Figures 3(a), 3(b), 3(e), 3(f), 3(i) and 3(j). Since scoring functions are relative to the overall dataset, we equivalently derive a dataset with correlation 00, size $100$ and $q=0.5$. The generation process was repeated $20$ times followed by an averaging of scores to reduce noise. Note that this procedure allows to evaluate practically impossible subgroups. When the cover size or NCR of a subgroup is close to that of the overall dataset, then the exceptionality (i.e. performance difference) is limited in practice.
4.4.1 Skew from Cover Size
Figure 3(a) shows the dependence of $\varphi^{rasl}$ on the number of instances in its input, yielding larger values for larger inputs even if predictive performance is fixed constant (see the horizontal gradient at fixed correlations). This can lead to low values of $\varphi^{rasl}$ on small subgroups, which do not necessarily correspond to a performance improvement over the complete dataset. Underperforming subgroups can therefore not be identified by a fixed threshold on the metric (e.g., $\geq 0$). This means that inherently smaller subgroups have to have a higher error rate to be considered important for ARL, and thus the approach returns results with a higher rate of passing the statistical test (Table 1). However, the strength of this effect can not be influenced by the user.
Neither $\varphi^{rPRAUC}$ nor $\varphi^{rROCAUC}$ change with cover size (3(e), 3(i)). Furthermore these figures show increased noise on the lower end of the cover size axes, indicating increased random effects on these scoring functions for subgroups with few instances which underlines the necessity of the statistical tests introduced by SubROC. Additionally, Figures 3(f) and 3(j) show that the introduced cover size weighting successfully leads to a dependence on input size, encoding a preference for larger inputs for both of these scoring functions.
4.4.2 Skew from Class Ratio
Figures 3(c) and 3(k) show that the ARL as well as the PR AUC depend on the class ratio. The PR AUC skew is known to be related to the unachievable area in PR space [boyd_unachievable_2012]. This skew has a similar influence on the interpretation of the scores as has the cover size skew of the ARL. Figure 3(g) shows a more subtle effect for the ROC AUC through its contour lines, that was also found by [cook_when_2020]. Figures 3(d), 3(h) and 3(l) show that the class balance weighting successfully introduces a preference towards subgroups with balanced classes. In particular, the score of subgroups with highly imbalanced classes is close to zero. Because cover size also has only a minor effect on the ROC AUC (cf. Figure 3(e)), searches utilizing the ROC AUC within SubROC can be expected to yield the most consistently underperforming subgroups over the widest range of cover sizes and class balances compared to the ARL and the PR AUC.
(a) $\varphi^{rasl}_{0,0}$
(b) $\varphi^{rasl}_{1,0}$
(c) $\varphi^{rasl}_{0,0}$
(d) $\varphi^{rasl}_{0,1}$
(e) $\varphi^{rROCAUC}_{0,0}$
(f) $\varphi^{rROCAUC}_{1,0}$
(g) $\varphi^{rROCAUC}_{0,0}$
(h) $\varphi^{rROCAUC}_{0,1}$
(i) $\varphi^{rPRAUC}_{0,0}$
(j) $\varphi^{rPRAUC}_{1,0}$
(k) $\varphi^{rPRAUC}_{0,0}$
(l) $\varphi^{rPRAUC}_{0,1}$
Figure 3: Skewed response of cover size and class balance weighted scoring functions depending on cover size, NCR and correlation (model performance) on synthetic data. Contour lines show smoothed paths along which interestingness is equal.
5 Related Work
We summarize work on producing descriptions for parts of a dataset where a soft classifier over- or underperforms and describe the differences to our approach.
[sagadeeva_sliceline_2021] only consider residuals-based loss functions to rank subgroups in contrast to our ranking-based measures. Similarly, [zhang_identifying_2017] does not use AUC metrics or the subgroup discovery framework. Other research allows for user-specified performance measures for subgroup ranking, but they do not provide specific support such as optimistic estimates or generalization-awareness[chung_automated_2020, zhang_sliceteller_2017]. [cabrera_fairvis_2019, chung_automated_2020, zhang_sliceteller_2017] propose automatic search approaches as part of generic interactive evaluation tools. None of these works applied the difference between the subgroup AUC to the overall dataset AUC for the ROC or PR curve in a subgroup discovery setting and investigated its properties in detail.
The following works are concerned with the ROC curve of subgroups in detail. [gardner_evaluating_2019] propose to use the area between the ROC curves of a model on two subgroups in the context of fairness of data slicing but do not discuss the mining of respective patterns. [carrington_deep_2023] propose to evaluate specific parts of the ROC curve and call these parts subgroups. Considered parts in that work are intervals in the x-axis or the y-axis of the ROC space as well as intervals in the thresholds used to compute the ROC points from. In addition to this differing notion of subgroups, they do not investigate an automated search for subgroups. [tafvizi_attributing_2022] presents methods for computing the attribution of subgroups to the overall AUC. They do this to identify parts of the data that contribute to a reduction in the overall AUC. The contribution of a subgroup to the overall dataset AUC is highly influenced by how the subgroup is embedded into the ordering of the overall dataset. Therefore finding parts of the dataset with an exceptional AUC would simply not contribute to their goal. Our goal is to find under which circumstances a model is most likely to fail or succeed, so we take a different approach with AUC-based exceptionality.
For subgroup discovery, many (tight) optimistic estimates (OE) have been proposed, primarily for settings with traditional target concepts (binary, numeric, or both) [grosskreutz_fast_2010, lemmerich_fast_2016]. In a few settings, tight OE have also been derived for more complex model classes in the EMM framework [kalofolias_discovering_2019, belfodil_identifying_2020]. An OE for a scoring function that computes the ROC AUC is given in [lemmerich_fast_2016]. However, there the ROC curve is based on different attributes than what is studied in this work, such that the OE is not applicable here. To the best of our knowledge, none of the previously presented OEs can be applied to the scoring functions used in this work.
6 Conclusion
We introduced SubROC, a novel framework for reliable and efficient discovery of interpretable subgroups with exceptional model performance with respect to soft classifiers. Our experiments demonstrate the necessity for such a framework as well as the effectiveness of SubROC when using the performance measures ARL, ROC AUC and PR AUC. We identified the side effects of a subgroup’s cover size and class balance on these performance measures and their intricate influence on the results. By providing optimistic estimates, we enable fast mining. To summarize, SubROC has the potential to enable routine evaluation of ML models, improve quality control procedures, inform model performance enhancements, and increase confidence in deployed ML models overall.
In this work, we focused on ranking-based soft classifier evaluation measures due to their widespread use and their specific challenges to Exceptional Model Mining. Other popular performance measures, such as cross-entropy or the Brier score, can be integrated into classic subgroup discovery. Their in-depth evaluation is an avenue for future research. Also, directly integrating statistical tests into the search may help to further improve the significance of the discovered subgroups. Finally, the investigation of alternatives to account for class imbalances is a promising direction for future work.
7 Acknowledgements
The funding for this study was in part provided by the BMBF (01IS22077).
[…]
Appendix A Proofs of Tight Optimistic Estimates
Firstly this section generalizes and formalizes the interestingness measure $\varphi^{rasl}$. Generic interestingness measures are defined in the following, that allow specialization by soft classifier performance measures. Here soft classifier performance measures are functions that assign a real value to ground truth and prediction pairs of soft classifiers.
Our formalizations make use of standard multiset notation as presented in [syropoulos_mathematics_2001]. To handle values of instances in a dataset $D=(\mathcal{I},\mathcal{A})$ we define a database as the multiset $d(D)=[(A_{1}(c),\dots,A_{|\mathcal{A}|}(c))|c\in\mathcal{I}]$. A shorthand notation for ground truth and prediction values is $I^{\{y,\hat{y}\}}=[(Y(c),\hat{Y}(c))|c\in I]$ on a set of instances $I$, the item values of which are referenced by $y_{i}$ and $\hat{y}_{i}$ with index $i$.
For the start, the aim of the presented interestingness measures is to assign the highest interestingness to the subgroups on which the model exhibits the worst performance. The type of a performance measure characterizes whether a higher value indicates better (score-type) or worse (loss-type) performance. A generic interestingness measure is defined for each performance measure type, negating score-type performance measure values such that higher values indicate worse performance.
Definition 1 (Soft Classifier Loss Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $m_{l}$ be a loss-type soft classifier performance measure. The soft classifier loss interestingness measure for $m_{l}$ is defined as
$\varphi^{m_{l}}(p)=m_{l}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
Definition 2 (Soft Classifier Score Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $m_{s}$ be a score-type soft classifier performance measure. The soft classifier score interestingness measure for $m_{s}$ is defined as
$\varphi^{m_{s}}(p)=-m_{s}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
These interestingness measures both increase with decreasing exhibited performance on the input data. It remains unclear at which interestingness value a subgroup exceeds the performance (positively or negatively) compared to the overall analyzed dataset. An example application of that information is to prune the search for subgroups exhibiting exceptionally good model performance, when the aim is to find subgroups exhibiting exceptionally bad model performance. The following interestingness measure definitions build on the previous definitions to include the comparison to the overall dataset performance.
Definition 3 (Relative Soft Classifier Loss Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $\varphi^{m_{l}}$ be a soft classifier loss interestingness measure for the loss-type soft classifier performance measure $m_{l}$. The relative soft classifier loss interestingness measure is defined as
$\varphi^{rm_{l}}(p)=\varphi^{m_{l}}(p)-m_{l}(\mathcal{I}^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
Definition 4 (Relative Soft Classifier Score Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $\varphi^{m_{s}}$ be a soft classifier score interestingness measure for the score-type soft classifier performance measure $m_{s}$. The relative soft classifier score interestingness measure is defined as
$\varphi^{rm_{s}}(p)=\varphi^{m_{s}}(p)+m_{s}(\mathcal{I}^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
The main interestingness measure $\varphi^{rasl}$ from [duivesteijn_understanding_2014] is defined with alternative signatures (compared to the main text) for use in the proofs as
$PEN_{i}([c_{1},\dots,c_{m}]):=$ $\sum_{j=1}^{m}(\mathbbm{1}_{\{(y,\hat{y})\mid y=0\wedge\hat{y}>\hat{y}_{i}\}}(y_{j},\hat{y}_{j})$
$+\frac{1}{2}\cdot\mathbbm{1}_{\{(y,\hat{y})\mid y=0\wedge\hat{y}=\hat{y}_{i}\}}(y_{j},\hat{y}_{j}))$
$RL_{i}([c_{1},\dots,c_{m}]):=$ $\mathbbm{1}_{\{1\}}(y_{i})\cdot PEN_{i}([c_{1},\dots,c_{m}])$
$ARL([c_{1},\dots,c_{m}]):=$ $\frac{\sum_{i=1}^{m}\mathbbm{1}_{\{1\}}(y_{i})\cdot PEN_{i}([c_{1},\dots,c_{m}])}{\sum_{i=1}^{m}\mathbbm{1}_{\{1\}}(y_{i})}$
$\varphi^{rasl}(p):=$ $ARL(sg(p)^{\{y,\hat{y}\}})-ARL(\mathcal{I}^{\{y,\hat{y}\}})$
for multisets of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$. $\varphi^{rasl}$ is a special case of Definition 3 with $m_{l}=ARL$. The definitions from the main text can be formalized on this basis as
$PEN(c,p):=$ $PEN_{index(c)}(sg(p)^{\{y,\hat{y}\}})$
$ARL(p):=$ $ARL(sg(p)^{\{y,\hat{y}\}})$
for patterns $p$ and instances $c$.
$\varphi^{ROCAUC}$ and $\varphi^{PRAUC}$ are both instances of Definition 2 with the ROC AUC and the PR AUC performance measures substituted in place of $m_{s}$ respectively. $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ are the same for Definition 4.
Note that when referring to the AUC of the PR curve in this work, we mean the area under the linearly interpolated PR curve. Therefore ROC AUC and PR AUC are defined as follows.
We define the count of positives and negatives in $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$) as
$P(C):=|[(y_{i},\hat{y}_{i})\in C\mid y_{i}=1]|$
and
$N(C):=|[(y_{i},\hat{y}_{i})\in C\mid y_{i}=0]|$
respectively.
Furthermore we define multisets of true positives, false positives, true negatives and false negatives in $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$) under threshold $t\in\mathbb{R}$ as
$TP(C,t)$ $:=[(y_{i},\hat{y}_{i})\in C\mid y_{i}=1\wedge\hat{y}_{i}>t]$
$FP(C,t)$ $:=[(y_{i},\hat{y}_{i})\in C\mid y_{i}=0\wedge\hat{y}_{i}>t]$
$TN(C,t)$ $:=[(y_{i},\hat{y}_{i})\in C\mid y_{i}=0\wedge\hat{y}_{i}\leq t]$
$FN(C,t)$ $:=[(y_{i},\hat{y}_{i})\in C\mid y_{i}=1\wedge\hat{y}_{i}\leq t]$
respectively.
We also define the set of ROC/PR supporting points as
$S(C):=\{(|TP(C,t)|,|FP(C,t)|)\mid t\in\{\hat{y}_{i}\mid(y_{i},\hat{y}_{i})\in C\}\cup\{-\infty\}\},$
which consists of (true positives, false positives) pairs constructed from applying unique $\hat{y}_{i}$ from $C$ as classification thresholds on $C$.
With that we define ROC AUC and PR AUC as follows.
Definition 5 (ROCAUC [fawcett_introduction_2006]).
The ROC AUC is defined as
$ROCAUC(C):=\frac{\sum_{i=2}^{|S(C)|}\mid FP_{i}-FP_{i-1}\mid\cdot\frac{TP_{i}+TP_{i-1}}{2}}{P(C)\cdot N(C)}$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$) with $S(C)$ indexed by the decreasing order of thresholds.
Definition 6 (PRAUC).
The (linear!) PR AUC is defined as
$PRAUC(C):=\frac{\sum_{i=2}^{|S(C)|}\mid TP_{i}-TP_{i-1}\mid\cdot\frac{\frac{TP_{i}}{TP_{i}+FP_{i}}+\frac{TP_{i-1}}{TP_{i-1}+FP_{i-1}}}{2}}{P(C)}$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$) with $S(C)$ indexed by the decreasing order of thresholds.
Both of these definitions are derived from the trapezoidal rule, which corresponds to the area under the linear interpolation of the set of points. The underlying performance measures Precision and Recall as well as a general AUC definition based on the trapezoidal rule read as follows.
We define the Precision and Recall performance measures on $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$) under threshold $t\in\mathbb{R}$ as
$Prec(C,t)$ $:=\frac{|TP(C,t)|}{|TP(C,t)|\cdot|FP(C,t)|}$
$Rec(C,t)$ $:=\frac{|TP(C,t)|}{P(C)}$
respectively.
Definition 7 (AUC).
The AUC of a set $\mathfrak{P}$ of at least two two-dimensional points $x_{i}=(x_{i,1},x_{i,2})$ (with indices $i$ such that points are in increasing order, primarily for the first dimension and secondarily for the second dimension) is defined using the trapezoidal rule as
$AUC(\mathfrak{P}):=\sum_{i=2}^{|\mathfrak{P}|}\left|x_{i,1}-x_{i-1,1}\right|\cdot\frac{x_{i,2}+x_{i-1,2}}{2}.$
The ROC AUC and PR AUC functions as they are used in the main text have different signatures. They can be formalized based on the definitions here as
$ROCAUC(I):=$ $ROCAUC(I^{\{y,\hat{y}\}})$
$PRAUC(I):=$ $PRAUC(I^{\{y,\hat{y}\}})$
for sets of instances $I$.
On a side note, the class balance can now be defined on multisets of ground truth and prediction values $C=I^{\{y,\hat{y}\}}$ for sets of instances $I$. By definition $P(C)=|\mathcal{P}_{I}|$ and $N(C)=|\mathcal{N}_{I}|$. We define $cb(C)$ as 0 if only one class occurs and as $\min\{\frac{N(C)}{P(C)},\frac{P(C)}{N(C)}\}$ otherwise.
Definitions of (tight) lower bounds of soft classifier performance measures are now given. They allow to structure the construction of some of the tight optimistic estimates that are presented in this section into a statement about a performance measure, considering a bound for the performance measure, and an adaptation step of that bound to interestingness measures, yielding an optimistic estimate. The lower bound definition (Definition 8) requires for a soft classifier performance measure an additional soft classifier performance measure that does not return smaller values than the reference performance measure for any input. The tightness criterion of Definition 9 requires a lower bound to be equal to the reference performance measure for at least one input. The performance measure lower bound definition and the bound tightness definition are defined analogously to optimistic estimates and tight optimistic estimates respectively.
Definition 8 (Soft Classifier Performance Measure Lower Bound).
Let $m$ and $b_{m}$ be soft classifier performance measures. $b_{m}$ is the soft classifier performance measure lower bound of $m$ if and only if
$\forall C^{\prime}\subseteq C:m(C^{\prime})\geq b_{m}(C)$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
Definition 9 (Tight Soft Classifier Performance Measure Bound).
Let $m$ be a soft classifier performance measure with soft classifier performance measure lower bound $b_{m}$. $b_{m}$ is called tight if and only if
$\exists C^{\prime}\subseteq C:m(C^{\prime})=b_{m}(C)$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
Definition 10 (Soft Classifier Performance Measure Upper Bound).
Let $m$ and $b_{m}$ be soft classifier performance measures. $b_{m}$ is the soft classifier performance measure upper bound of $m$ if and only if
$\forall C^{\prime}\subseteq C:m(C^{\prime})\leq b_{m}(C)$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
Lemma 7 (Result Transition from Loss to Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $m_{l}$ be a loss-type soft classifier performance measure with tight soft classifier performance measure upper bound $b_{m_{l}}$. Let $\varphi^{m_{l}}$ be a soft classifier loss interestingness measure based on $m_{l}$. A tight optimistic estimate for $\varphi^{m_{l}}$ is given by
$oe_{\varphi^{m_{l}}}(p)=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
Proof.
Let $p$ be a pattern that is defined on $D$.
What has to be shown is that $oe_{\varphi^{m_{l}}}$ is an optimistic estimate of $\varphi^{m_{l}}$:
$\forall p^{\prime}\supset p:\varphi^{m_{l}}(p^{\prime})\leq oe_{\varphi^{m_{l}}}(p)$ (1)
and that this optimistic estimate is tight:
$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi^{m_{l}}(p^{\prime})=oe_{\varphi^{m_{l}}}(p).$ (2)
Proof for Equation 1
By definition of $\varphi^{m_{l}}$ and $oe_{\varphi^{m_{l}}}$ this equation is equivalent to
$\forall p^{\prime}\supset p:m_{l}(sg(p^{\prime})^{\{y,\hat{y}\}})\leq b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (3)
The right side of this equation is constant since $D$ and $p$ are fixed.
[lemmerich_novel_2014] states that for all two rules $p_{gen}$ and $p_{spec}$ the following statement holds
$p_{gen}\subset p_{spec}\Rightarrow sg(p_{gen})\supseteq sg(p_{spec}).$
It follows
$\forall p^{\prime}\supset p:sg(p^{\prime})\subseteq sg(p).$
From the definition of databases follows
$\forall p^{\prime}\supset p:sg(p^{\prime})^{\{y,\hat{y}\}}\subseteq sg(p)^{\{y,\hat{y}\}}.$ (4)
Because $b_{m_{l}}$ is defined to be a soft classifier performance measure upper bound of $m_{l}$ we know that
$\forall C^{\prime}\subseteq C:m_{l}(C^{\prime})\leq b_{m_{l}}(C)$ (5)
holds for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
Equation 3 follows from Equation 4 and Equation 5, concluding the proof for Equation 1.
Proof for Equation 2
By definition of $\varphi^{m_{l}}$ and $oe_{\varphi^{m_{l}}}$ this equation is equivalent to
$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge m_{l}(sg(p^{\prime})^{\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$
The equivalence of $sg(p^{\prime})$ and $sg^{\prime}$ as well as the triviality of $sg^{\prime}=sg^{\prime}$ imply that the following equation is equivalent:
$\exists sg^{\prime}\subseteq sg(p):m_{l}(sg^{\prime\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (6)
Because $b_{m_{l}}$ is defined to be a tight soft classifier performance measure upper bound of $m_{l}$ we know that
$\exists C^{\prime}\subseteq C:m_{l}(C^{\prime})=b_{m_{l}}(C)$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
By fixing $C=sg(p)^{\{y,\hat{y}\}}$ we get
$\exists C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}:m_{l}(C^{\prime})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (7)
The definition of databases states that for any dataset $D^{\prime}=(\mathcal{I}^{\prime},\mathcal{A}^{\prime})$ every instance in $\mathcal{I}^{\prime}$ creates exactly one element in $d(D^{\prime})$.
It follows
$\forall C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}\exists\mathcal{I}^{\prime}\subseteq sg(p):C^{\prime}=\mathcal{I^{\prime}}^{\{y,\hat{y}\}}.$ (8)
From Equation 7 and Equation 8 follows
$\exists C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}\exists\mathcal{I}^{\prime}\subseteq sg(p):C^{\prime}=\mathcal{I^{\prime}}^{\{y,\hat{y}\}}\wedge m_{l}(C^{\prime})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (9)
The indifference of the order in a series of the same quantifiers, the equality of $C^{\prime}$ and $\mathcal{I^{\prime}}^{\{y,\hat{y}\}}$ as well as simplification rules imply that the following is equivalent to Equation 9
$\exists\mathcal{I}^{\prime}\subseteq sg(p):m_{l}(\mathcal{I^{\prime}}^{\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$
This is trivially equivalent to Equation 6, concluding the proof for Equation 2. ∎
Lemma 8 (Result Transition from Score to Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $m_{s}$ be a score-type soft classifier performance measure with tight soft classifier performance measure lower bound $b_{m_{s}}$. Let $\varphi^{m_{s}}$ be a soft classifier score interestingness measure based on $m_{s}$. A tight optimistic estimate for $\varphi^{m_{s}}$ is given by
$oe_{\varphi^{m_{s}}}(p)=-b_{m_{s}}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
Proof.
Let $m_{s}^{\prime}(C)=-m_{s}(C)$ and $b_{m_{s}^{\prime}}(C)=-b_{m_{s}}(C)$ for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$). $m_{s}^{\prime}$ is a loss-type soft classifier performance measure by definition and $b_{m_{s}^{\prime}}$ is a tight soft classifier performance measure upper bound for $m_{s}^{\prime}$.
It follows from Lemma 7 that the soft classifier loss interestingness measure $\varphi^{m_{s}^{\prime}}$ based on $m_{s}^{\prime}$ has the following tight optimistic estimate
$oe_{\varphi^{m_{s}^{\prime}}}(p)=b_{m_{s}^{\prime}}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
By definition of $\varphi^{m_{s}}$ and $\varphi^{m_{s}^{\prime}}$ we know that
$\varphi^{m_{s}}(p)=-m_{s}(sg(p)^{\{y,\hat{y}\}})=\varphi^{m_{s}^{\prime}}(p)$
for any pattern $p$ that is defined on $D$.
It follows that $oe_{\varphi^{m_{s}^{\prime}}}$ is also a tight optimistic estimate for $\varphi^{m_{s}}$.
$oe_{\varphi^{m_{s}^{\prime}}}$ and $oe_{\varphi^{m_{s}}}$ define the same function. ∎
The following lemma allows to translate an optimistic estimate for a soft classifier loss (or score) interestingness measure to an optimistic estimate for a relative soft classifier loss (or score) interestingness measure.
Lemma 9 (Result Transition to Relative Interestingness Measure).
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $\varphi$ be an interestingness measure with tight optimistic estimate $oe_{\varphi}$. Let $c\in\mathbb{R}$ be a constant. A tight optimistic estimate for the interestingness measure $\varphi^{\prime}(p)=\varphi(p)+c$ is given by
$oe_{\varphi^{\prime}}(p)=oe_{\varphi}(p)+c$
for any pattern $p$ that is defined on $D$.
Proof.
Let $p$ be a pattern that is defined on $D$.
What has to be shown is that $oe_{\varphi^{\prime}}$ is an optimistic estimate of $\varphi^{\prime}$:
$\forall p^{\prime}\supset p:\varphi^{\prime}(p^{\prime})\leq oe_{\varphi^{\prime}}(p)$ (10)
and that this optimistic estimate is tight:
$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi^{\prime}(p^{\prime})=oe_{\varphi^{\prime}}(p).$ (11)
Proof for Equation 10
By definition of $\varphi^{\prime}$ and $oe_{\varphi^{\prime}}$ this equation is equivalent to
$\forall p^{\prime}\supset p:\varphi(p^{\prime})+c\leq oe_{\varphi}(p)+c.$ (12)
Because $oe_{\varphi}$ is defined to be an optimistic estimate of $\varphi$ we know that
$\forall p^{\prime}\supset p:\varphi(p^{\prime})\leq oe_{\varphi}(p)$ (13)
Equation 12 is true if and only if Equation 13 is true, concluding the proof for Equation 10.
Proof for Equation 11
By definition of $\varphi^{\prime}$ and $oe_{\varphi^{\prime}}$ this equation is equivalent to
$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi(p^{\prime})+c=oe_{\varphi}(p)+c.$ (14)
Because $oe_{\varphi}$ is defined to be a tight optimistic estimate of $\varphi$ we know that
$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi(p^{\prime})=oe_{\varphi}(p).$ (15)
Equation 14 is true if and only if Equation 15 is true, concluding the proof for Equation 11. ∎
All in all by sequentially applying Lemma 7 (respectively Lemma 8) and then Lemma 9 to a soft classifier performance measure upper (respectively lower) bound for a performance measure $m$ we obtain a tight optimistic estimate for the relative soft classifier loss (respectively score) interestingness measure based on $m$. This proving scheme is used to prove tight optimistic estimates for the relative ARL, ROC AUC and PR AUC based interestingness measures.
A.1 Average Ranking Loss
Lemma 10 (Tight Soft Classifier Performance Measure Upper Bound $b_{ARL}$).
A tight soft classifier performance measure upper bound for the average ranking loss is given by the function
$b_{ARL}(C)=\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
Proof.
Let $C=[c_{1},\dots,c_{n}]$ be a multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$.
What has to be shown is that $b_{ARL}$ is a soft classifier performance measure upper bound for the average ranking loss:
$\forall C^{\prime}\subseteq C:ARL(C^{\prime})\leq b_{ARL}(C)$ (16)
and that this soft classifier performance measure upper bound is tight:
$\exists C^{\prime}\subseteq C:ARL(C^{\prime})=b_{ARL}(C).$ (17)
Proof for Equation 16
By definition of the average ranking loss (see main text) and $b_{ARL}$ this equation is equivalent to
$\forall C^{\prime}\subseteq C:\frac{\sum_{i=1}^{|C^{\prime}|}RL_{i}(C^{\prime})}{\sum_{i=1}^{|C^{\prime}|}\mathbbm{1}_{\{1\}}(y^{\prime}_{i})}\leq\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}.$ (18)
The ranking loss $RL_{i}$ for the instance at position $i$ in any given multiset of instance values $c_{j}\in\{0,1\}\times\mathbb{R}$ defines a weighted sum of the negative instances with rank at least as high as $i$ with weights that do not change for sub-multisets of the input. It is trivial that the maximum ranking loss of instances in $C$ does not increase for sub-multisets $C^{\prime}$ of $C$:
$\forall C^{\prime}\subseteq C:\max_{i\in\{1,\dots,|C^{\prime}|\}}\{RL_{i}(C^{\prime})\}\leq\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ (19)
It follows that the following expression implies Equation 18.
$\forall C^{\prime}\subseteq C:\frac{\sum_{i=1}^{|C^{\prime}|}RL_{i}(C^{\prime})}{\sum_{i=1}^{|C^{\prime}|}\mathbbm{1}_{\{1\}}(y^{\prime}_{i})}\leq\max_{i\in\{1,\dots,|C^{\prime}|\}}\{RL_{i}(C^{\prime})\}$ (20)
It remains to show that Equation 20 holds.
It is trivial that the average of a multiset $[m_{1},\dots,m_{|M|}]=M$ of real numbers $m_{i}\in\mathbb{R}$ is not greater than the maximum of $M$:
$\frac{\sum_{i=1}^{|M|}m_{i}}{|M|}\leq\max M$ (21)
For any $C^{\prime}=[c_{1}^{\prime},\dots,c_{n}^{\prime}]$ (multiset of instance values $c_{i}^{\prime}\in\{0,1\}\times\mathbb{R}$) and corresponding multiset of ranking loss values of positive instances $M_{C^{\prime}}=[m_{i}=RL_{i}(C^{\prime})\mid i\in\{1,\dots,|C^{\prime}|\}\wedge(y^{\prime}_{i},\hat{y}^{\prime}_{i})\in C^{\prime}\wedge y^{\prime}_{i}=1]$ the $ARL$ of $C^{\prime}$ is equal to the average of the values in $M_{C^{\prime}}$:
$\frac{\sum_{i=1}^{|C^{\prime}|}RL_{i}(C^{\prime})}{\sum_{i=1}^{|C^{\prime}|}\mathbbm{1}_{\{1\}}(y^{\prime}_{i})}=\frac{\sum_{i=1}^{|M_{C^{\prime}}|}m_{i}}{|M_{C^{\prime}}|}$ (22)
Equation 20 follows from Equation 21 and Equation 22, concluding the proof for Equation 16.
Proof for Equation 17
By definition of the average ranking loss and $b_{ARL}$ this equation is equivalent to
$\exists C^{\prime}\subseteq C:\frac{\sum_{i=1}^{|C^{\prime}|}RL_{i}(C^{\prime})}{\sum_{i=1}^{|C^{\prime}|}\mathbbm{1}_{\{1\}}(y^{\prime}_{i})}=\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}.$ (23)
The ranking loss $RL_{i}$ for the instance at position $i$ in any given multiset of instance values $c_{j}\in\{0,1\}\times\mathbb{R}$ defines a weighted sum of the negative instances with rank at least as high as $i$ with weights that do not change for sub-multisets of the input.
Let $i=\arg\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ and $C^{\prime}=(C\circledast\{(y,\hat{y})\in C\mid y=0\wedge\hat{y}\geq\hat{y}_{i}\})\cup[c_{i}]$. $C^{\prime}$ consists of the positive instance $c_{i}$ and exactly the elements of $C$ that define the non-zero summands in the definition of $PEN_{i}$ (see main text). By definition of the ranking loss, $C^{\prime}$ demonstrates the existence of a sub-multiset of $C$ for which Equation 23 holds.
This concludes the proof for Equation 17. ∎
See 1
Proof.
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset.
Lemma 10 states that a tight soft classifier performance measure upper bound for the (loss-type) average ranking loss is given by the function
$b_{ARL}(C)=\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$
for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
It follows from Lemma 7 that a tight optimistic estimate for $\varphi^{asl}$ is given by
$oe_{\varphi^{asl}}(p)=\max_{i\in\{1,\dots,|sg(p)|\}}\{RL_{i}(sg(p)^{\{y,\hat{y}\}})\}$
for any pattern $p$ that is defined on $D$.
$c=-ARL(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rasl}$ is given by
$oe_{\varphi^{rasl}}(p)=\max_{i\in\{1,\dots,|sg(p)|\}}\{RL_{i}(sg(p)^{\{y,\hat{y}\}})\}-ARL(\mathcal{I}^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
This result is equivalent to the function in the theorem. ∎
A.2 Area Under the ROC Curve
See 2
Proof.
Let $I$ be a set of instances such that $C=I^{\{y,\hat{y}\}}=[c_{1},\dots,c_{n}]$ is a multiset of instance values $(y_{i},\hat{y}_{i})\in\{0,1\}\times\mathbb{R}$ that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each.
We show that $b_{ROCAUC}$ is a tight lower bound using an alternative signature:
$b_{ROCAUC}(C)=\begin{cases}1,&\text{if }\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C:\\ &(y<y^{\prime}\rightarrow\hat{y}<\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}>\hat{y}^{\prime})\\ \frac{1}{2},&\text{if }\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C:\\ &(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})\\ 0,&\text{otherwise}\end{cases}$
The three cases $b_{ROCAUC}(C)=1$, $b_{ROCAUC}(C)=\frac{1}{2}$ and $b_{ROCAUC}(C)=0$ are treated separately.
The case $b_{ROCAUC}(C)=1$ is present if and only if $C$ contains perfect predictions, meaning that all positive instances in $C$ have strictly greater predicted values than all negative instances in $C$. All sub-multisets $C^{\prime}$ of $C$ such that $ROCAUC(C^{\prime})\neq undefined$ must contain at least one positive and one negative instance. It follows that all eligible sub-multisets of $C$ also contain perfect predictions. The ROC AUC is equal to $1$ for perfect predictions. $b_{ROCAUC}$ consequently meets the criterion of a soft classifier performance measure lower bound for $ROCAUC$ on perfect predictions. The tightness criterion is also met since $ROCAUC(C)=1$.
The case $b_{ROCAUC}(C)=\frac{1}{2}$ is present if and only if $C$ contains perfect predictions with ties, meaning that predictions in $C$ are similar to the case $b_{ROCAUC}(C)=1$ except for the negative instances with the highest predicted value (compared to other negative instances) and the positive instances with the lowest predicted value (compared to other positive instances), which have equal predicted values (i.e. they are tied). First we show that $b_{ROCAUC}$ satisfies the tightness criterion in this case and then that it is a soft classifier performance measure lower bound for ROC AUC in this case.
The sub-multiset of $C$ that only contains ties has a ROC AUC of $\frac{1}{2}$. Therefore the tightness criterion is fulfilled in this case.
From the definition of $b_{ROCAUC}$ we know that in this case
$\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C:(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$ (24)
holds. This condition also holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined), because of the universal quantifier. Let $C^{\prime}$ be such a sub-multiset. We continue with a proof by contradiction, showing that the premise $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that Equation 24 is not true. Firstly the diagonal ROC curve with only supporting points $(0,0)$ and $(1,1)$ has an AUC of $\frac{1}{2}$. Clearly any ROC curve with only supporting points $(0,0)$, $(1,1)$ and additional points above or on the diagonal ROC curve has an AUC of $\frac{1}{2}+x$ with $x\in\mathbb{R}_{\geq 0}$. Therefore $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that at least one supporting point of the ROC curve of $C^{\prime}$ lies below the diagonal ROC curve. This is equivalently expressed by
$\exists(TP,FP)\in S(C^{\prime}):$ $\frac{TP}{P(C^{\prime})}$ $<\frac{FP}{N(C^{\prime})}$
$\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}\cup\{-\infty\}:$ $\frac{TP(C^{\prime},t)}{P(C^{\prime})}$ $<\frac{FP(C^{\prime},t)}{N(C^{\prime})}$
$\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}:$ $\frac{1}{P(C^{\prime})}\cdot|[(y,\hat{y})$ $\in C^{\prime}\mid y=1\wedge\hat{y}>t]|<$
$\frac{1}{N(C^{\prime})}\cdot|[(y,\hat{y})$ $\in C^{\prime}\mid y=0\wedge\hat{y}>t]|.$
Because the FDR and TPR are both limited to the range $[0,1]$, it follows
$\Rightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}:TP(C^{\prime},t)\neq P(C^{\prime})\wedge FP(C^{\prime},t)\neq 0.$
This implies that there exists a positive instance in $C^{\prime}$ that is not counted in $TP(C^{\prime},t)$ and that there exists a negative instance in $C^{\prime}$ that is counted in $FP(C^{\prime},t)$. Now from the definitions of $TP(C^{\prime},t)$ and $FP(C^{\prime},t)$ follows
$\Rightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}:$
$(\exists(y,\hat{y})\in C^{\prime}:y=1\wedge\neg(\hat{y}>t))\wedge(\exists(y,\hat{y})\in C^{\prime}:y=0\wedge\hat{y}>t)$
$\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}:y=1\wedge\hat{y}\leq t\wedge y^{\prime}=0\wedge\hat{y}^{\prime}>t$
$\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}:y=1\wedge y^{\prime}=0\wedge\hat{y}\leq t<\hat{y}^{\prime}$
If there exist instances $(y,\hat{y})$ and $(y^{\prime},\hat{y}^{\prime})$ in $C^{\prime}$ that fulfill this expression, then they also fulfill the following expressions:
$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime}$
$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $(y<y^{\prime}\wedge\hat{y}>\hat{y}^{\prime})\vee(y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime})$
The negation of Equation 24 is equivalent:
$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $\neg(y\geq y^{\prime}\wedge\hat{y}\leq\hat{y}^{\prime})\vee\neg(y\leq y^{\prime}\wedge\hat{y}\geq\hat{y}^{\prime})$
$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $\neg((y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime}))$
$\Leftrightarrow$ $\neg\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$
Since in case $b_{ROCAUC}(C)=\frac{1}{2}$ we know that Equation 24 holds, we can conclude that $ROCAUC(C^{\prime\prime})\geq\frac{1}{2}$ holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined). Therefore $b_{ROCAUC}$ is a soft classifier performance measure lower bound for ROC AUC in this case.
The case $b_{ROCAUC}(C)=0$ is present if and only if $C$ contains erroneous predictions, meaning that there exists at least two instances $(y,\hat{y})$ and $(y^{\prime},\hat{y}^{\prime})$ in $C$ such that $y=0$, $y^{\prime}=1$ and $\hat{y}>\hat{y}^{\prime}$. The sub-multiset of $C$ that only contains these two instances (once each) has an ROC AUC of 00. This is the overall minimum of the ROC AUC. $b_{ROCAUC}$ consequently meets the criteria of a tight soft classifier performance measure lower bound for $ROCAUC$ in case $b_{ROCAUC}(C)=0$. ∎
See 3
Proof.
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset.
Lemma 2 states that a tight soft classifier performance measure lower bound for the (score-type) area under the ROC curve is given by $b_{ROCAUC}$.
It follows from Lemma 8 that a tight optimistic estimate for $\varphi^{ROCAUC}$ is given by
$oe_{\varphi^{ROCAUC}}(p)=-b_{ROCAUC}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
$c=ROCAUC(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rROCAUC}$ is given by
$oe_{\varphi^{rROCAUC}}=ROCAUC(\mathcal{I}^{\{y,\hat{y}\}})-b_{ROCAUC}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
This result is equivalent to the function in the theorem. ∎
A.3 Area Under the Precision-Recall Curve
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111recallprecisionborderapproximation 1approximation 2
Figure 4: Lower border of the achievable area in PR space and approximations for a positive class ratio $\frac{1}{2}$. The curve named “border” is the actual border of the achievable area. The curves named “approximation 1” and “approximation 2” are the (linearly interpolated) PR curves for the multisets of instance values $[(1,0),(1,0),(0,1),(0,1)]$ and $[(1,0),(1,0.5),(0,1),(0,1)]$ respectively. Both include only erroneous predictions, meaning that all negative instances have greater predicted values than all positive instances.
See 4
Proof.
Let $I$ be a set of instances such that $C=I^{\{y,\hat{y}\}}$ is a multiset of instance values $(y,\hat{y})\in\{0,1\}\times\mathbb{R}$ that contains at least one instance with $y=1$. Let
$(1,\hat{y}_{minpos})=c_{minpos}=\arg\min_{(y,\hat{y})\in[(y^{\prime},\hat{y}^{\prime})\in C\mid y^{\prime}=1]}\{\hat{y}\}$
and
$C_{worst}=[c_{minpos}]\cup(C\circledast\{(y,\hat{y})\in C\mid y=0\}).$
It holds $I_{worst}^{\{y,\hat{y}\}}=C_{worst}$, therefore $b_{PRAUC}(I)=PRAUC(C_{worst})$. Define for convenience $b_{PRAUC}(C)=PRAUC(C_{worst})$.
The following three cases (27, 28 and 29) are treated separately:
$\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C:(y<y^{\prime}\rightarrow\hat{y}<\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}>\hat{y}^{\prime})$ (25)
$\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C:(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$ (26)
[…]
Assuming case 27.
This case is present if and only if $C$ contains perfect predictions, meaning that all positive instances in $C$ have strictly greater predicted values than all negative instances in $C$. All sub-multisets $C^{\prime}$ of $C$ such that $PRAUC(C^{\prime})\neq undefined$ must contain at least one positive instance. It follows that all eligible sub-multisets of $C$ also contain perfect predictions or only positives. The PR AUC is equal to $1$ for perfect predictions and for only positives. In particular, this also holds for $C_{worst}$. $b_{PRAUC}$ consequently meets the criterion of a soft classifier performance measure lower bound for $PRAUC$ on perfect predictions. The tightness criterion is also met since $PRAUC(C)=1$.
Assuming case 28.
Define
$T(C)$ :=$:=\{(-\infty,\hat{y}_{1})\}\cup\{(\hat{y}_{i-1},\hat{y}_{i})\mid i\in\{2,\dots,|C|\}\wedge\hat{y}_{i-1}\neq\hat{y}_{i}\}$
$T_{<\hat{y}_{minpos}}(C)$ :=$:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}<\hat{y}_{minpos}\}$
$T_{\geq\hat{y}_{minpos}}(C)$ :=$:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}\geq\hat{y}_{minpos}\}$
From the definition of $\hat{y}_{minpos}$ follows
$\forall t\in\{\hat{y}\mid(y,\hat{y})\in C\}:$ $t<\hat{y}_{minpos}$
$\Rightarrow TP(C,t)=P(C)$
$\Rightarrow Rec(C,t)=\frac{P(C)}{P(C)}=1,$
which implies
$\forall(t_{lo},t_{hi})\in T_{<\hat{y}_{minpos}}(C):$
$\quad\left|Rec(C,t_{hi})-Rec(C,t_{lo})\right|\cdot\frac{Prec(C,t_{hi})+Prec(C,t_{lo})}{2}$
$\quad=\left|1-1\right|\cdot\frac{Prec(C,t_{hi})+Prec(C,t_{lo})}{2}=0.$
From the definitions of $PRAUC(C)$, $T(C)$, $T_{<\hat{y}_{minpos}}(C)$ and $T_{\geq\hat{y}_{minpos}}(C)$ follows
$PRAUC(C)$
$=\sum_{(t_{lo},t_{hi})\in T(C)}\left|Rec(C,t_{hi})-Rec(C,t_{lo})\right|\cdot\frac{Prec(C,t_{hi})+Prec(C,t_{lo})}{2}$
$=\sum_{(t_{lo},t_{hi})\in T_{<\hat{y}_{minpos}}(C)}\left|Rec(C,t_{hi})-Rec(C,t_{lo})\right|\cdot\frac{Prec(C,t_{hi})+Prec(C,t_{lo})}{2}$
$+\sum_{(t_{lo},t_{hi})\in T_{\geq\hat{y}_{minpos}}(C)}\left|Rec(C,t_{hi})-Rec(C,t_{lo})\right|\cdot\frac{Prec(C,t_{hi})+Prec(C,t_{lo})}{2}.$
Therefore
$PRAUC(C)$
$=\sum_{(t_{lo},t_{hi})\in T_{\geq\hat{y}_{minpos}}(C)}\left|Rec(C,t_{hi})-Rec(C,t_{lo})\right|\cdot\frac{Prec(C,t_{hi})+Prec(C,t_{lo})}{2}$
$=PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}})$ (30)
with
$C_{\hat{y}\geq\hat{y}_{minpos}}:=$ $\ C\ominus(C\circledast\{(y,\hat{y})\in C\mid\hat{y}<\hat{y}_{minpos}\})$
$=$ $\ C\circledast\{(y,\hat{y})\in C\mid\hat{y}\geq\hat{y}_{minpos}\}.$
Define $\mathcal{C}_{\hat{y}\subseteq\hat{y}_{minpos}}$ similarly for any multiset $\mathcal{C}$ of instance values $(y,\hat{y})\in\{0,1\}\times\mathbb{R}$ that contains at least one instance with $y=1$.
In particular, Equation 30 also holds for any sub-multiset $C^{\prime}$ of $C$ with corresponding $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$. Therefore
$\forall C^{\prime}\subseteq C:PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})=PRAUC(C^{\prime})$
and
$b_{PRAUC}(C)=b_{PRAUC}(C_{\hat{y}\geq\hat{y}_{minpos}}).$
Therefore the lower bound condition on $C$
$\forall C^{\prime}\subseteq C:PRAUC(C^{\prime})\geq b_{PRAUC}(C)$
is true if and only if the lower bound condition holds on $C_{\hat{y}\geq\hat{y}_{minpos}}$
$\forall C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}:PRAUC(C^{\prime})\geq b_{PRAUC}(C_{\hat{y}\geq\hat{y}_{minpos}}).$
By definition of $C_{\hat{y}\geq\hat{y}_{minpos}}$
$\forall C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}:C^{\prime}=C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}.$
Therefore the lower bound condition on $C_{\hat{y}\geq\hat{y}_{minpos}}$ is equivalent to
$\forall C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}:PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})\geq b_{PRAUC}(C_{\hat{y}\geq\hat{y}_{minpos}}).$
For any sub-multiset $C^{\prime}$ of $C_{\hat{y}\geq\hat{y}_{minpos}}$ to have
$PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})<PRAUC(C_{worst})$
there must be at least one point on the PR curve of $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ below that of $C_{worst}.$ Consider any
$t\in\{\hat{y}\mid(y,\hat{y})\in C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}\wedge\hat{y}\geq\hat{y}_{minpos}\}.$
From the condition of case 28 follows
$FP(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime},t)=0.$
This implies
$Prec(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime},t)=1.$
Therefore for any sub-multiset $C^{\prime}$ of $C_{\hat{y}\geq\hat{y}_{minpos}}$ no supporting point on the PR curve of $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ with such a threshold can lie below the PR curve of $C_{worst}$. This also holds for the interpolated points between the supporting points because of the linear interpolation.
It remains to show that the point with $t=-\infty$ for $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ is not below the PR curve of $C_{worst}$. Clearly
$Rec(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime},-\infty)=1$
and
$Prec(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime},-\infty)=\frac{P(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})}{P(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})+N(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})}.$
By definition
$P(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})\geq 1=P(C_{worst})$
and
$N(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})\leq N(C_{\hat{y}\geq\hat{y}_{minpos}})=N(C_{worst}).$
Therefore
$Prec(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime},-\infty)\geq Prec(C_{worst},-\infty).$
In conclusion, for this case we know that all points on the PR curve of any sub-multiset of $C$ must either not lie below the PR curve of $C_{worst}$ or have no influence on the AUC. Therefore the lower bound condition on $C$ is fulfilled by $b_{PRAUC}$ in this case.
Assuming case 29.
We show that no $C^{\prime}\subseteq C$ produces a PR point below the PR curve of $C_{worst}$, which implies
$\nexists C^{\prime}\subseteq C:PRAUC(C^{\prime})<b_{PRAUC}(C_{\hat{y}\geq\hat{y}_{minpos}}).$
We know that $PRAUC(C)=PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}})$, so it suffices to show
$\nexists C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}:PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})<b_{PRAUC}(C_{\hat{y}\geq\hat{y}_{minpos}}).$ (31)
In this case (with $C=C_{\hat{y}\geq\hat{y}_{minpos}}$) the supporting points of the PR curve of $C_{worst}$ are
$\left\{(0,1),(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
=$=\left\{(0,1),(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
Let
$\mathfrak{P}_{worst}^{\prime}$ =$=\left\{(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
=$=\left\{(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
Clearly the AUC between the points $(0,1)$ and $(0,0)$ is 00, so
$PRAUC(C_{worst})=AUC(\mathfrak{P}_{worst}^{\prime}).$
$\mathfrak{P}_{worst}^{\prime}$ are the supporting points of the minimum PR curve described in [boyd_unachievable_2012]. As $\mathfrak{P}_{worst}^{\prime}$ shows, the minimum PR curve only depends on the positive class ratio. By definition of $C_{worst}$ there is no $C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}$ such that the positive class ratio of $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ is strictly lower than that of $C_{worst}$. Also no point on the linear interpolation of $\mathfrak{P}_{worst}^{\prime}$ is strictly above the correct PR interpolation of $\mathfrak{P}_{worst}^{\prime}$ as shown in the following.
The points on the correct interpolation between the two points in $\mathfrak{P}_{worst}^{\prime}$ according to [boyd_unachievable_2012] are
$\left(r_{i},\frac{\pi r_{i}}{(1-\pi)+\pi r_{i}}\right)$
with $r_{i}=\frac{i}{1+N(C_{worst})}$ and $\pi=\frac{1}{1+N(C_{worst})}$ for $0\leq i\leq 1$.
The points on the linear interpolation between the two points in $\mathfrak{P}_{worst}^{\prime}$ are
$\left(r_{i},\pi r_{i}\right)$
with $r_{i}=\frac{i}{1+N(C_{worst})}$ and $\pi=\frac{1}{1+N(C_{worst})}$ for $0\leq i\leq 1$.
For the denominator of the second component in the points of the correct interpolation holds
$(1-\pi)+\pi r_{i}$
$=$ $\left(1-\frac{1}{1+N(C_{worst})}\right)+\frac{1}{1+N(C_{worst})}\cdot\frac{i}{1+N(C_{worst})}$
$=$ $\frac{1+N(C_{worst})-1}{1+N(C_{worst})}+\frac{1\cdot i}{(1+N(C_{worst}))^{2}}$
$\leq$ $\frac{N(C_{worst})}{1+N(C_{worst})}+\frac{1}{(1+N(C_{worst}))^{2}}$
$=$ $\frac{1+N(C_{worst})\cdot(1+N(C_{worst}))}{(1+N(C_{worst}))^{2}}$
$=$ $\frac{1+N(C_{worst})+N(C_{worst})^{2}}{1+2N(C_{worst})+N(C_{worst})^{2}}$
$\leq$ $1$
Therefore
$\frac{\pi r_{i}}{(1-\pi)+\pi r_{i}}\geq\pi r_{i}.$
It follows that there is no $C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}$ such that the PR curve of $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ has a point that lies strictly below the PR curve of $C_{worst}$. Thus 31 is fulfilled and we can conclude that $b_{PRAUC}$ is a lower bound for the PR AUC in case 29. It is also tight since $b_{PRAUC}(C)=PRAUC(C_{worst})$ and $C_{worst}\subseteq C$. ∎
See 5
Proof.
Let $D=(\mathcal{I},\mathcal{A})$ be a dataset.
Lemma 4 states that a tight soft classifier performance measure lower bound for the (score-type) area under the PR curve is given by $b_{PRAUC}$.
It follows from Lemma 8 that a tight optimistic estimate for $\varphi^{PRAUC}$ is given by
$oe_{\varphi^{PRAUC}}(p)=-b_{PRAUC}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
$c=PRAUC(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rPRAUC}$ is given by
$oe_{\varphi^{rPRAUC}}=PRAUC(\mathcal{I}^{\{y,\hat{y}\}})-b_{PRAUC}(sg(p)^{\{y,\hat{y}\}})$
for any pattern $p$ that is defined on $D$.
This result is equivalent to the function in the theorem. ∎
See 6
Proof.
Let $I$ be a set of instances such that $C=I^{\{y,\hat{y}\}}$ is a multiset of instance values $(y,\hat{y})\in\{0,1\}\times\mathbb{R}$ and $\alpha\leq\beta\in\mathbb{R}_{>0}$ fixed real numbers.
From the equalities $P(C)=|\mathcal{P}_{I}|$, $N(C)=|\mathcal{N}_{I}|$ and $cb(I)=cb(C)$, follows $w(I)=|C|^{\alpha}\cdot cb(C)^{\beta}$. Therefore we simply write $w(C)$. Similarly we use $b_{w}(C)=(2\cdot\min{\{P(C),N(C)\}})^{\alpha}$ in place of $b_{w}(I)$.
By definition
$|C|=P(C)+N(C).$
Therefore
$w(C)$ $=(P(C)+N(C))^{\alpha}\cdot cb(C)^{\beta}.$
Assume $P(C)>N(C)$ and $N(C)>0$ (Equivalent to the first case of the minimum in the $cb$ definition).
By definition of $cb(C)$
$w(C)=(P(C)+N(C))^{\alpha}\cdot\left(\frac{N(C)}{P(C)}\right)^{\beta}.$
Let $c\in[(y,\hat{y})\in C\mid y=1]$ be any positive instance in $C$.
We show that the weighting $w$ does not decrease when $c$ is removed from $C$.
$w(C\ominus[c])$ $\geq w(C)$
$\Rightarrow$ $(P(C)-1+N(C))^{\alpha}\left(\frac{N(C)}{P(C)-1}\right)^{\beta}$ $\geq(P(C)+N(C))^{\alpha}\left(\frac{N(C)}{P(C)}\right)^{\beta}$
$\Leftrightarrow$ $\frac{P(C)-1+N(C)}{P(C)+N(C)}$ $\geq\left(\frac{N(C)}{P(C)}\right)^{\frac{\beta}{\alpha}}\cdot\left(\frac{P(C)-1}{N(C)}\right)^{\frac{\beta}{\alpha}}$
$\Leftrightarrow$ 11 11 $\geq\left(1-\frac{1}{P(C)}\right)^{\frac{\beta}{\alpha}}+\frac{1}{P(C)+N(C)}$ (32)
In case $\alpha=\beta$ follows
11 11 $\geq\left(1-\frac{1}{P(C)}\right)^{1}+\frac{1}{P(C)+N(C)}$
$\Leftrightarrow$ $N(C)$ $\geq 0$
This is true because of the assumption $N(C)>0$.
From
$0\leq 1-\frac{1}{P(C)}\leq 1$
follows
$\left(1-\frac{1}{P(C)}\right)^{c}\leq 1-\frac{1}{P(C)}$
for any constant $c\geq 1$.
Therefore Equation 32 holding in case $\alpha=\beta$ implies that it also holds in case $\alpha\leq\beta$. We conclude that the weighting $w$ does not decrease when $c$ is removed from $C$.
Note that the case $P(C)=N(C)+1$ works out here even though $cb(C\ominus[c])$ is not defined by the first case of the minimum in the $cb$ definition anymore. This follows from $P(C\ominus[c])=N(C\ominus[c])$ in this case.
Let $c^{\prime}\in[(y,\hat{y})\in C\mid y=0]$ be any negative instance in $C$. Clearly $w(C\ominus[c^{\prime}])\leq w(C)$, so removing any negative instance from $C$ does not increase the weighting $w$.
In conclusion the weighting $w$ is maximized in the first case of the minimum in the $cb$ definition by sequentially removing a positive instance until the thereby created sub-multiset $C^{\prime}$ of $C$ does not fulfill this case anymore, but fulfills $P(C^{\prime})=N(C^{\prime})$.
For such a sub-multiset $C^{\prime}$ of $C$ holds $cb(C^{\prime})=1$. Therefore
$w(C^{\prime})=(P(C^{\prime})+N(C^{\prime}))^{\alpha}.$
From the definition of $C^{\prime}$ follows
$P(C^{\prime})=N(C^{\prime})=\min\{P(C),N(C)\}.$
Therefore
$w(C^{\prime})=(2\cdot\min\{P(C),N(C)\})^{\alpha}.$
Assume $N(C)>P(C)$ and $P(C)>0$ (Equal to the second case of the minimum in the $cb$ definition).
It is analogous to the above to proof that removing any negative instance in $C$ does not decrease the weighting $w$ and that removing any positive instance in $C$ does not increase the weighting $w$.
Therefore the weighting $w$ is maximized in the second case of the the minimum in the $cb$ definition by sequentially removing a negative instance until the thereby created sub-multiset $C^{\prime}$ of $C$ does not fulfill $N(C^{\prime})>P(C^{\prime})$ anymore, but fulfills $N(C^{\prime})=P(C^{\prime})$.
It is shown above that for such a sub-multiset $C^{\prime}$ of $C$ the following holds:
$w(C^{\prime})=(2\cdot\min\{P(C),N(C)\})^{\alpha}$
Assume $N(C)=0\vee P(C)=0$ (Equal to the case in the $cb$ definition returning 00).
In this case
$\forall C^{\prime}\subseteq C:cb(C^{\prime})=0.$
Therefore
$\forall C^{\prime}\subseteq C:w(C^{\prime})=|C^{\prime}|^{\alpha}\cdot 0^{\beta}=0.$
$b_{w}$ is a tight upper bound in this case since it also returns 00:
$b_{w}(C)=(2\cdot\min\{P(C),N(C)\})^{\alpha}=(2\cdot 0)^{\alpha}=0$
Appendix B Experimental Results
B.1 Dataset statistics.
Table 3: Dataset metadata
Abbreviation Name #Rows NCR #Features
Census UCI Census-Income (KDD) 199523 0.0621 41
Adult OpenML Adult 48842 0.2393 14
Bank UCI Bank Marketing 45211 0.8830 16
Credit UCI Credit Card Clients 30000 0.7788 23
Mushroom UCI Mushroom 8124 0.5180 22
Statlog Statlog (German Credit Data) 1000 0.3000 20
Cancer UCI Breast Cancer Wisconsin 699 0.6552 9
Approval UCI Credit Approval 690 0.5551 15
B.2 Quality Parameters of Results Across Multiple Datasets
Table 4: Overview of result set properties for combinations of datasets and performance measures for an xgboost classifier. (Continuation of Table 1 in the main text)
Mean Exceptionality
Mean Cover Size
Mean NCR
Mean Pairwise IoU
Significant/ Top-5
Top-5/ Significant/ Original
Dataset Metric     b f     b f     b f     b f     b f
Mushroom n=8124 NCR=0.52 ROC AUC     0.00 -    203 -    0.76 -    0.28 -    0/5 0/0/100
PR AUC     0.00 0.00    35 57    0.17 0.49    0.14 -    0/5 1/1/100
ARL     - -    - -    - -    - -    0/0 0/0/0
Statlog n=1000 NCR=0.30 ROC AUC     0.74 -    23 -    0.04 -    0.36 -    0/5 0/0/100
PR AUC     0.50 -    21 -    0.63 -    0.95 -    0/5 0/0/100
ARL     0.16 -    241 -    0.28 -    - -    0/1 0/0/1
Cancer n=699 NCR=0.66 ROC AUC     0.03 -    41 -    1.00 -    0.17 -    0/5 0/0/5
PR AUC     0.55 -    78 -    1.00 -    0.80 -    0/5 0/0/7
ARL     0.18 -    83 -    1.00 -    - -    0/1 0/0/1
Table 5: Overview of result set properties for combinations of datasets and performance measures for an xgboost classifier. No cover size weighting, class balance weighting, generalization awareness or significance filtering was applied. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. Here, the search returns the top-100 subgroups, which were filtered for significance, and finally, the top-5 subgroups are returned.
Mean Exceptionality
Mean Pairwise IoU
Mean Cover Size
Mean NCR
Result Set Size (Top-5/ Significant/ Original)
Dataset Interestingness Measure
UCI Census-Income (KDD) ROC AUC 0.95 0.46 33 0.03 5/11/100
PR AUC 0.84 0.60 20 0.94 5/13/100
ARL 46.72 0.99 8000 0.15 5/100/100
OpenML Adult ROC AUC 0.80 0.16 25 0.06 5/14/100
PR AUC 0.93 0.21 26 0.95 5/14/100
ARL 147.66 0.85 4614 0.42 5/82/82
UCI Bank Marketing ROC AUC - - - - 0/0/100
PR AUC - - - - 0/0/100
ARL - - - - 0/0/0
UCI Credit Card Clients ROC AUC - - - - 0/0/100
PR AUC - - - - 0/0/100
ARL - - - - 0/0/0
UCI Mushroom ROC AUC - - - - 0/0/100
PR AUC - - - - 0/0/100
ARL - - - - 0/0/0
Statlog (German Credit Data) ROC AUC - - - - 0/0/100
PR AUC - - - - 0/0/100
ARL - - - - 0/0/1
UCI Breast Cancer Wisconsin ROC AUC - - - - 0/0/8
PR AUC - - - - 0/0/15
ARL - - - - 0/0/1
UCI Credit Approval ROC AUC 0.58 - 49 0.98 1/1/100
PR AUC 0.89 0.81 40 0.97 5/12/100
ARL 20.40 0.68 49 0.97 5/8/100
B.3 Case Study: Adult
The following are full result sets of the searches on the Adult dataset, that are also reported in Table 1 in the main text in terms of aggregated statistics. Note in particular the amount of overlap between patterns in the result sets from the base SubROC setting compared to the subgroups in the results of the full SubROC framework. Moreover values of the main text table are reflected in the extreme performance results, cover sizes and negative class ratios.
Table 6: Result set of basic SubROC setting with ARL.
Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
Table 7: Result set of basic SubROC setting with ROC AUC.
Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
Table 8: Result set of basic SubROC setting with PR AUC.
Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
Table 9: Result set of full SubROC framework with ARL.
Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
Table 10: Result set of full SubROC framework with ROC AUC.
Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
Table 11: Result set of full SubROC framework with PR AUC.
Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
B.4 Searching for an Injected Subgroup
For details on the experimental setup, see Section 4.3 in the main text.
Figure 5: IoU of an injected subgroup with subgroups in a top-10 ARL result set, ordered decreasingly by interestingness. “baseline” was obtained searching in a base SubROC setting, the other results were obtained using the full SubROC framework with corresponding cover size weighting $\alpha$ and class balance weighting $\beta$.
Figure 6: IoU of an injected subgroup with subgroups in a top-10 PR AUC result set, ordered decreasingly by interestingness. “baseline” was obtained searching in a base SubROC setting, the other results were obtained using the full SubROC framework with corresponding cover size weighting $\alpha$ and class balance weighting $\beta$.
B.5 Optimistic Estimates
Table 12: Speedups of median runtimes using optimistic estimates. (Continuation of Table 2 in the main text)
Performance Measure
Weight ($\alpha=\beta$)
Mushroom n/m=8124/22
Statlog n/m=1000/20
Cancer n/m=699/9
ARL 0    381.7 2.6 2.8
0.1    292.1 1.7 1.5
0.3    283.2 1.7 1.1
1    280.6 1.6 1.2
PR AUC 0    11.8 0.6 2.2
0.1    5.2 0.5 0.6
0.3    5.2 0.5 0.6
1    5.2 0.5 0.6
ROC AUC 0    29.8 0.8 0.8
0.1    3.6 0.6 0.5
0.3    3.7 0.6 0.5
1    3.6 0.7 0.5
Table 13: Median runtimes in seconds for ARL with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 3. 10 measurements were performed for each configuration.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 6899 723.8 6795 712.9 6877 667.3 6872 592.3
OpenML Adult 177.7 11.6 178.0 13.9 177.9 12.9 178.8 11.5
UCI Bank Marketing 28.8 5.9 28.9 6.9 29.0 6.9 29.2 6.8
UCI Credit Card Clients 523.3 7.5 525.2 10.1 530.3 10.0 531.4 10.0
UCI Mushroom 17.7 0.2 17.8 0.3 17.9 0.3 17.8 0.3
Statlog (German Credit Data) 3.3 1.6 3.3 2.2 3.3 2.2 3.3 2.2
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.1 0.2 0.1 0.1 0.1
UCI Credit Approval 0.7 0.5 0.7 0.9 0.7 1.0 0.7 1.0
Table 14: Median runtimes in seconds for ROC AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 3. 10 measurements were performed for each configuration.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 5469 4785 5499 5391 5609 4814 5477 1429
OpenML Adult 187.2 192.5 187.8 197.7 186.4 182.9 187.6 44.1
UCI Bank Marketing 39.6 21.2 39.8 43.3 39.9 40.6 40.0 26.6
UCI Credit Card Clients 633.1 18.8 647.0 658.6 641.0 633.9 645.6 367.6
UCI Mushroom 20.5 1.0 20.7 7.7 20.6 7.6 20.8 7.7
Statlog (German Credit Data) 5.7 6.1 5.6 7.0 5.6 7.0 5.6 6.9
UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.3 0.2 0.3 0.2 0.3
UCI Credit Approval 1.0 1.2 1.0 1.6 1.0 1.6 1.0 1.6
Table 15: Median runtimes in seconds for PR AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 3. 10 measurements were performed for each configuration.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 7007 3645 6923 4149 7014 3502 7147 1522
OpenML Adult 192.4 121.7 194.2 173.6 192.2 114.4 191.8 23.9
UCI Bank Marketing 35.8 28.2 35.9 39.5 35.5 37.1 35.6 23.2
UCI Credit Card Clients 595.4 457.1 599.8 578.2 598.9 405.8 597.4 110.5
UCI Mushroom 24.2 3.3 24.3 7.4 24.3 7.4 24.3 7.2
Statlog (German Credit Data) 4.7 6.0 4.7 7.1 4.8 7.1 4.8 6.7
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.3 0.2 0.3 0.2 0.3
UCI Credit Approval 0.9 0.9 0.9 1.6 0.9 1.7 0.9 1.7
Table 16: Speedups of median runtimes reported in Tables 13, 14 and 15.
Performance Measure
Weight ($\alpha=\beta$)
UCI Census-Income (KDD)
OpenML Adult
UCI Bank Marketing
UCI Credit Card Clients
UCI Mushroom
Statlog (German Credit Data)
UCI Breast Cancer Wisconsin
UCI Credit Approval
ARL 0 9.5 15.3 4.9 69.4 98.9 2.1 2.5 1.4
0.1 9.5 12.8 4.2 52.2 64.3 1.5 1.6 0.8
0.3 10.3 13.8 4.2 52.8 64.4 1.5 1.1 0.7
1 11.6 15.6 4.3 52.9 64.8 1.5 1.1 0.7
PR AUC 0 1.9 1.6 1.3 1.3 7.3 0.8 1.7 1.0
0.1 1.7 1.1 0.9 1.0 3.3 0.7 0.6 0.6
0.3 2.0 1.7 1.0 1.5 3.3 0.7 0.6 0.6
1 4.7 8.0 1.5 5.4 3.4 0.7 0.5 0.5
ROC AUC 0 1.1 1.0 1.9 33.7 21.6 0.9 0.9 0.9
0.1 1.0 0.9 0.9 1.0 2.7 0.8 0.6 0.6
0.3 1.2 1.0 1.0 1.0 2.7 0.8 0.6 0.6
1 3.8 4.3 1.5 1.8 2.7 0.8 0.6 0.7
Table 17: Share of remaining evaluated subgroups in the same searches reported in Tables 13, 14 and 15
Performance Measure
Weight ($\alpha=\beta$)
UCI Census-Income (KDD)
OpenML Adult
UCI Bank Marketing
UCI Credit Card Clients
UCI Mushroom
Statlog (German Credit Data)
UCI Breast Cancer Wisconsin
UCI Credit Approval
ARL 0 4.51% 2.47% 6.34% 0.41% 0.21% 28.23% 20.76% 48.52%
0.1 4.10% 2.37% 6.34% 0.41% 0.21% 28.23% 20.76% 53.70%
0.3 3.74% 2.13% 6.34% 0.41% 0.21% 28.23% 20.76% 61.64%
1 3.12% 1.81% 6.34% 0.41% 0.21% 28.23% 20.76% 67.35%
PR AUC 0 42.38% 48.05% 45.50% 59.99% 8.48% 100.00% 20.76% 51.82%
0.1 48.67% 76.41% 89.06% 87.05% 16.68% 100.00% 51.18% 88.98%
0.3 39.40% 42.41% 75.48% 47.02% 16.70% 100.00% 51.18% 94.74%
1 11.98% 4.73% 24.25% 7.79% 15.56% 89.35% 62.44% 91.12%
ROC AUC 0 83.00% 97.49% 47.80% 3.10% 2.74% 100.00% 71.65% 94.74%
0.1 91.62% 97.49% 98.05% 99.03% 22.29% 100.00% 71.65% 94.74%
0.3 79.95% 87.01% 82.80% 94.49% 22.29% 100.00% 71.65% 94.74%
1 13.92% 11.33% 28.98% 36.66% 22.29% 94.27% 71.65% 92.11%
Figure 7: Histogram of standard deviations divided by medians for all runtimes of Tables 13, 14 and 15. Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.
Table 18: Median runtimes in seconds for ARL with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 98310 5346 98206 5265 98876 4958 98284 4223
OpenML Adult 978.1 20.6 995.1 25.2 969.7 24.0 983.5 22.0
UCI Bank Marketing 119.5 12.4 119.2 15.5 119.9 15.7 119.4 15.6
UCI Credit Card Clients 4510 8.4 4479 11.6 4449 11.6 4491 11.6
UCI Mushroom 109.5 0.3 110.8 0.4 109.8 0.4 109.9 0.4
Statlog (German Credit Data) 9.8 3.8 10.0 5.9 9.9 5.9 9.8 6.0
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.2 0.2 0.2 0.1 0.1
UCI Credit Approval 1.7 1.0 1.8 2.2 1.8 2.5 1.8 2.8
Table 19: Median runtimes in seconds for ROC AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 73835 731.2 72673 73204 74144 57059 74759 11694
OpenML Adult 961.8 1076 965.9 1138 971.8 1014 976.0 116.3
UCI Bank Marketing 150.1 69.9 150.5 199.2 149.8 170.8 148.0 72.0
UCI Credit Card Clients 4907 58.6 5000 5256 4954 4569 4967 1267
UCI Mushroom 107.2 3.6 107.6 29.7 108.3 29.6 106.9 29.9
Statlog (German Credit Data) 15.4 18.2 15.4 25.0 15.2 25.2 15.3 23.4
UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.4 0.2 0.4 0.2 0.4
UCI Credit Approval 2.3 3.1 2.3 4.8 2.3 4.8 2.3 4.7
Table 20: Median runtimes in seconds for PR AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$. Subgroups were required to cover at least 20 instances and the search stopped at depth 4. 10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
0 0.1 0.3 1
Dataset Off On Off On Off On Off On
UCI Census-Income (KDD) 99395 34225 101272 52252 101602 37638 101372 8896
OpenML Adult 1044 444.9 1060 917.9 1032 483.9 1031 52.0
UCI Bank Marketing 137.4 79.2 137.4 189.0 137.7 162.6 138.8 62.3
UCI Credit Card Clients 4779 1899 4823 4199 4842 1681 4731 209.3
UCI Mushroom 146.4 12.5 149.2 29.0 149.0 28.9 148.7 28.5
Statlog (German Credit Data) 13.2 20.8 13.3 28.1 13.3 27.9 13.4 25.3
UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.3 0.2 0.3 0.2 0.4
UCI Credit Approval 2.1 1.8 2.1 5.1 2.1 5.4 2.1 5.2
Table 21: Share of remaining evaluated subgroups in the same searches reported in Tables 18, 19 and 20
Performance Measure
Weight ($\alpha=\beta$)
UCI Census-Income (KDD)
OpenML Adult
UCI Bank Marketing
UCI Credit Card Clients
UCI Mushroom
Statlog (German Credit Data)
UCI Breast Cancer Wisconsin
UCI Credit Approval
ARL 0 2.39% 0.68% 2.66% 0.04% 0.03% 19.99% 18.97% 27.64%
0.1 2.17% 0.66% 2.66% 0.04% 0.03% 19.99% 18.97% 32.30%
0.3 1.97% 0.61% 2.66% 0.04% 0.03% 19.99% 18.97% 39.50%
1 1.54% 0.53% 2.66% 0.04% 0.03% 19.99% 18.97% 47.98%
PR AUC 0 25.80% 27.77% 22.54% 22.70% 5.31% 99.60% 18.97% 30.55%
0.1 42.09% 66.38% 84.34% 65.46% 10.43% 99.60% 47.47% 83.66%
0.3 27.81% 28.28% 65.27% 17.37% 10.49% 98.95% 47.47% 92.61%
1 4.32% 1.57% 11.72% 1.23% 9.85% 87.02% 57.76% 88.14%
ROC AUC 0 0.99% 96.32% 38.06% 1.05% 2.10% 99.60% 68.17% 93.91%
0.1 90.86% 96.32% 97.18% 92.96% 16.28% 99.60% 68.17% 93.91%
0.3 67.43% 83.35% 74.42% 75.78% 16.28% 99.60% 68.17% 93.91%
1 8.41% 4.76% 15.70% 11.74% 16.28% 92.08% 68.17% 89.20%
Figure 8: Histogram of standard deviations divided by medians for all runtimes of Tables 18, 19 and 20. Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.
For the speedups of median runtimes reported in Tables 18, 19 and 20, see Table 2 in the main text.

### v2

@@
-2024_subroc.bib
+SubROC: AUC-Based Discovery of Exceptional Subgroup Performance for Binary Classifiers
 Abstract
 Machine learning (ML) is increasingly employed in real-world applications like medicine or economics, thus, potentially affecting large populations.
-However, ML models often do not perform homogeneously across such populations resulting in subgroups of the population (e.g., sex=female $\wedge$ marital_status=married) where the model underperforms or, conversely, is particularly accurate.
-Identifying and describing such subgroups can support practical decisions on which subpopulation a model is safe to deploy or where more training data is required.
-The potential of identifying and analyzing such subgroups has been recognized, however, an efficient and coherent framework for effective search is missing.
+However, ML models often do not perform homogeneously, leading to underperformance or, conversely, unusually high performance in certain subgroups (e.g., sex=female $\wedge$ marital_status=married).
+Identifying such subgroups can support practical decisions on which subpopulation a model is safe to deploy or where more training data is required.
+However, an efficient and coherent framework for effective search is missing.
 Consequently, we introduce SubROC, an open-source, easy-to-use framework based on Exceptional Model Mining for reliably and efficiently finding strengths and weaknesses of classification models in the form of interpretable population subgroups.
@@
 1 Introduction
-Machine learning (ML) is widely adopted in high-stake application domains such as health care, loans, hiring decisions, or legal decision-making [mehrabi_survey_2021].
-Thus, ML models and their decisions already influence large populations and everyday life.
-However, it has also repeatedly been shown that ML models can be highly heterogeneous in their performance [mehrabi_survey_2021].
-Subgroups of a population may exist where a ML model underperforms (or performs remarkably well).
-For illustration, see Figure 1 where predictions are particularly bad for the subpopulation of women from rural areas (sex=female $\wedge$ housing=rural) and remarkably accurate for male teenagers (sex=male $\wedge$ $\leq$18).
-Knowing such subgroups may allow data scientists and practitioners to make informed decisions about model deployment and development.
-Here, the practitioners may decide to use the current model only for male teenagers, where the model performs well, and the data scientists may decide to collect more data for subgroups where the model underperforms.
-Figure 1: In an exemplary application, binary classification could distinguish between healthy and diseased patients (left).
-Predictions of soft classifiers impose an ordering on the persons, which performs better the clearer healthy patients are separated from diseased patients.
-In this case there is a clear difference between the densities of healthy and diseased patients in this ordering, which are represented by the red and blue curves here (middle left).
-A dataset can be split up in different ways by patterns describing subsets.
-Our approach based on subgroup discovery evaluates these subsets systematically and returns those where the model predictions create the best or worst separation of the classes (middle right).
+[…]
+Figure 1: Illustrative example.
+Predictions of a binary soft classifier produce a ranking of patients, from healthy to sick.
+The clearer this separation, the better the classifier.
+Here, the density distributions of healthy and diseased patients (red and blue curves, middle left) are clearly distinguishable but there is no perfect separation.
+SubROC automatically discovers interpretable subgroups where the model predictions create the best or worst separation of the classes (middle right).
 Results may highlight arbitrary sensitive subgroups of a population, where the model under- or over-performs compared to the rest of the dataset (right).
@@
 The need and potential of identifying and analyzing subgroups has been widely recognized.
-For example, a wide array of work aims at alleviating consistent underperformance in known [sagawa_distributionally_2020] as well as previously unknown [sohoni_no_2020] subgroups.
-Similar observations have been made concerning the fairness of ML models [mehrabi_survey_2021].
-Consequently, various approaches have been explored to study such atypical subgroups systematically, e.g., with interactive tools for manual discovery of subgroups [baylor_tfx_2017, kahng_visual_2016].
+For example, a wide array of work aims at alleviating consistent underperformance in known [2] as well as previously unknown [3] subgroups.
+Similar observations have been made concerning the fairness of ML models [1].
+Consequently, various approaches explore such atypical subgroups systematically, e.g., via interactive tools for manual discovery [4, 5].
 However, this is labor-intensive, inefficient, and poses the threat of missing essential subgroups.
-Thus, automated approaches have been proposed that enable the search for interpretable subgroups where ML models over- or underperform [chung_automated_2020, cabrera_fairvis_2019].
-The most principled approach employs the framework of Exceptional Model Mining (EMM) with a custom performance measure for efficient exhaustive search [duivesteijn_understanding_2014].
-However, none of these approaches combine the automated and exhaustive search of EMM, standard measures for soft classifiers like ROC AUC or PR AUC, efficient search, and handling of class imbalance and statistical significance testing, which are required for a routine and practical analysis.
+Thus, automated approaches enable the search for subgroups where ML models over- or underperform [6, 7].
+The most principled approach employs Exceptional Model Mining (EMM) with a custom performance measure for efficient exhaustive search [8].
+However, none of these approaches combines the automated and exhaustive search of EMM, standard measures for soft classifiers like ROC AUC or PR AUC, efficient search, and handling of class imbalance and statistical significance testing, which are required for a routine and practical analysis.
 Contribution.
-Consequently, we introduce SubROC, which builds on EMM to reliably and efficiently find strengths and weaknesses of classification models.
-For this, we (a) propose a new group of scoring functions based on the area-under-the-curve (AUC) performance statistic, (b) derive tight optimistic estimates that enable fast exhaustive subgroup search, (c) provide a well-rounded framework that includes control for class imbalance, pruning of redundant results, and significance testing, (d) show the practical benefits of our approach across multiple datasets, and (e) integrate SubROC in the easy-to-use, open source pysubgroup package and provide the code for our experiments.
-[^1] Experiment code is available at https://github.com/bckrlab/subroc.
-Overall, SubROC will enable practitioners to routinely evaluate their ML models by finding subgroups where they over- or underperform.
-This will not only have the potential to impact and improve quality control procedures and increase the confidence in deployed ML models but also lay the foundation for downstream approaches to systematically improve their performance.
+Consequently, we introduce SubROC, which builds on EMM to reliably and efficiently find strengths and weaknesses of classification models: We (a) propose a new group of scoring functions based on the area-under-the-curve (AUC) statistic, (b) derive tight optimistic estimates enabling fast exhaustive subgroup search, (c) provide a well-rounded framework that includes control for class imbalance, pruning of redundant results, and significance testing, (d) show the practical benefits of our approach across multiple datasets, and (e) integrate SubROC in the easy-to-use, open source pysubgroup package and provide the code for our experiments.
+[^1] https://github.com/bckrlab/subroc, archival version at [9]
+SubROC will enable practitioners to routinely evaluate ML models by finding subgroups where they over- or underperform.
+This will not only improve quality control procedures and increase the confidence in deployed ML models but also lay the foundation for downstream approaches to systematically improve their performance.
 2 Preliminaries
-2.1 Soft Classifier Evaluation
-SubROC focuses on binary soft classification.
+2.1 Notation
 $D=(\mathcal{I},\mathcal{A})$ is a dataset with instances $c\in\mathcal{I}$ and attributes (features) $\mathcal{A}=(X_{1},\ldots,X_{m})$.
@@
 $Y(c)\in\{1,0\}$ is the ground truth label and $\hat{Y}(c)\in[0,1]$ a continuous (soft) prediction.
-Based on the ground truth labels, we call some instances positives ($\mathcal{P}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=1\}$) and some negatives ($\mathcal{N}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=0\}$).
-A classifier performs well if it returns high (low) values for instances with the ground truth $Y(c)=1$ ($Y(c)=0$).
-While many evaluation metrics for soft classification exist, including binary cross-entropy or the Brier score, the area under the curve (AUC) of the Receiver Operating Characteristic (ROC) curve is among the most widely used.
+We call instances with $Y(c)=1$ positives ($\mathcal{P}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=1\}$) and the rest negatives ($\mathcal{N}_{\mathcal{I}}=\{c\in\mathcal{I}|Y(c)=0\}$).
+A classifier performs well if it returns high (low) values for positives $\mathcal{P}_{\mathcal{I}}$ (negatives $\mathcal{N}_{\mathcal{I}}$).
+2.2 Soft Classifier Evaluation
+Traditional subgroup discovery on $Y\neq\hat{Y}$ is applicable to hard classification but ignores classifier confidence.
+Since in practice most models provide probability estimates, SubROC focuses on soft classification with the performance measures Receiver Operating Characteristic (ROC) area under the curve (AUC) and Precision-Recall (PR) AUC that require labels $Y$ as well as prediction scores $\hat{Y}$.
+This leads to conceptually different, not necessarily better results.
+While many evaluation measures for soft classification exist, including binary cross-entropy or the Brier score, the AUC of the ROC curve is among the most widely used.
 It assumes a ranking of instances according to their predicted probabilities from high to low.
-Then, it computes the probability of a randomly chosen positive instance being ranked above a randomly chosen negative instance.
-A main reason for the popularity of AUC-based evaluation metrics is that they summarize across all potential thresholds that a user could apply to conduct a hard classification into two classes based on the soft classifier results, e.g., depending on misclassification costs.
-This makes the ROC AUC a natural choice for interpretable subgroup descriptions, contributing to the insights that subgroups from our framework may provide to experts.
-2.2 Subgroup Discovery and Exceptional Model Mining
+It can be interpreted as the probability of a randomly chosen positive instance being ranked above a randomly chosen negative instance.
+One reason for the popularity of AUC-based performance measures is that they summarize classifier performance across all potential thresholds that a user can apply on soft classifier probabilities to separate samples into two classes e.g., depending on misclassification costs.
+This makes the ROC AUC a natural choice for interpretable subgroup descriptions, contributing to the insights our framework may provide to experts.
+Nevertheless, ROC AUC has also been criticized, for example, for being misleading under strong class imbalance and for treating all thresholds as equally relevant.
+2.3 Subgroup Discovery / Exceptional Model Mining
 A selector $s\in\mathcal{S}$ is a Boolean function based on a single attribute, e.g., $sex=female$ or $age\in[20,30)$.
 A pattern $p\in 2^{\mathcal{S}}$ is a conjunction of selectors, e.g., $sex=female\wedge age\in[20,30)$.
-In a given dataset, a pattern’s cover $sg(p)$ represents the set of all instances covered by it, i.e., all instances with $p=true$.
+Given a dataset, a pattern’s cover $sg(p)$ represents the set of all instances covered by it, i.e., all instances with $p=true$.
 A pattern and its cover are referred to as a subgroup.
-We call a subgroup $p_{spec}$ a specialization of another subgroup $p_{gen}$ (its generalization) if its pattern’s conjunction of selectors adds at least one additional selector: $p_{gen}\subset p_{spec}$.
-Subgroup discovery (SD) [klosgen_explora_1996] finds subgroups that show an interesting deviation with respect to a target concept, which is traditionally given as a single data attribute.
+A subgroup $p_{spec}$ is a specialization of another subgroup $p_{gen}$ (its generalization) if its set of selectors is a proper superset of those of the other subgroup: $p_{gen}\subset p_{spec}$.
+Subgroup discovery (SD) [10] finds subgroups that show an interesting deviation with respect to a target concept, which is traditionally given as a single data attribute.
 E.g., an interesting subgroup may contain a higher-than-average disease occurrence.
-EMM [leman_exceptional_2008] extends target concepts to model classes.
+EMM [11] extends target concepts to model classes.
 Here, a subgroup is considered interesting if the model parameters fitted to the subgroup are significantly different from those fitted to the complete dataset.
-SD and EMM formalize interestingness using scoring functions $\varphi:2^{\mathcal{S}}\rightarrow\mathbb{R}$ (also called interestingness measures or quality measures) which map a pattern $p\in 2^{\mathcal{S}}$ to a score given a dataset $D$.
+SD and EMM formalize interestingness using scoring functions $\varphi:2^{\mathcal{S}}\rightarrow\mathbb{R}$ (also called interestingness measures or quality measures) which map a pattern $p\in 2^{\mathcal{S}}$ to a score, given a dataset $D$.
 A search algorithm explores the search space of patterns ($2^{\mathcal{S}}$) to find the ones with the highest scores.
 The computational challenge of subgroup discovery lies in the exponentially growing search space $2^{\mathcal{S}}$ with the number of selectors $|\mathcal{S}|$.
-Beyond straightforward constraints like pattern length or a minimum number of covered instances, optimistic estimates $oe_{\varphi}$ are integral for optimizing runtimes.
-An $oe_{\varphi}$ provides an upper bound on the score $\varphi$ of all specializations $p_{spec}$ of a subgroup $p_{gen}$ [lemmerich_novel_2014]: $\forall p_{spec}\supset p_{gen}:\varphi(p_{spec})\leq oe_{\varphi}(p_{gen})$.
-For top-k search, this can be used to disregard all specializations of $p_{gen}$ if $oe_{\varphi}(p_{gen})$ is lower than the score of all current top-k patterns.
-Furthermore, tight optimistic estimates [grosskreutz_tight_2008] provide tighter upper bounds by guaranteeing that the upper bound can be reached by a hypothetical specialization $p_{spec}$, for an arbitrary subset of instances $I\subseteq sg(p_{gen})$ covered by $p_{gen}$.
+Beyond straightforward constraints like the number of selectors in a pattern or the minimum support, in terms of the number of covered instances, optimistic estimates are integral for optimizing runtimes.
+An optimistic estimate on a score $\varphi$}, $oe_{\varphi}$ provides an upper bound on the score $\varphi$ of all specializations $p_{spec}$ of a subgroup $p_{gen}$ [12]: $\forall p_{spec}\supset p_{gen}:\varphi(p_{spec})\leq oe_{\varphi}(p_{gen})$.
+When searching for the top-k patterns with the highest interestingness, this can be used to prune all specializations of $p_{gen}$ if $oe_{\varphi}(p_{gen})$ is lower than the score of all current top-k patterns.
+Furthermore, tight optimistic estimates [13] provide tighter upper bounds by guaranteeing that the upper bound can be reached by a hypothetical specialization $p_{spec}$, for an arbitrary subset of instances $I\subseteq sg(p_{gen})$ covered by $p_{gen}$.
 Optimistic estimates can be integrated into various search algorithms.
@@
 2.4 SCaPE Model Class
-For EMM, the Soft Classifier Performance Evaluation (SCaPE) model class has been proposed to discover subgroups with exceptional performance of soft classifiers [duivesteijn_understanding_2014].
+For EMM, the Soft Classifier Performance Evaluation (SCaPE) model class has been proposed to discover subgroups with exceptional performance of soft classifiers [8].
 It integrates the ground truth labels $Y$ and the model predictions $\hat{Y}$.
-The original scoring function Relative Average Subranking Loss (RASL), written $\varphi^{rasl}$ here, proposed for SCaPE, measures the number of violations of the ranking induced by the predictions $\hat{Y}(c)$ of the given soft classifier against the ranking based on the ground truth labels $Y(c)$ within a subgroup $sg$ with pattern $p$: For each positive instance $c\in\mathcal{P}_{p}$ for a pattern $p$, it counts how many negative instances $c^{\prime}\in\mathcal{N}_{p}$ are ranked higher.
+The original scoring function Relative Average Subranking Loss (RASL) $\varphi^{rasl}$ proposed for SCaPE, measures the number of violations of the ranking induced by the predictions $\hat{Y}(c)$ of the given soft classifier, against the ranking based on the ground truth labels $Y(c)$: For each positive instance $c\in\mathcal{P}_{p}$ for a pattern $p$, it counts how many negative instances $c^{\prime}\in\mathcal{N}_{p}$ are ranked higher.
 This count is denoted as $PEN(c,p)$.
 The Average Ranking Loss (ARL) is then defined as the average of $PEN(c,p)$ across all positive instances.
-The final scoring function $\varphi^{rasl}$ is defined relative to the overall dataset: $\varphi^{rasl}(p):=ARL(p)-ARL(\mathcal{\emptyset})$.
-Although the ARL score is normalized by the number of positive instances, it is heavily influenced by the number of overall instances covered by a pattern as the penalty score $PEN(c,p)$ counts the absolute number of violations.
-This favors large subgroups or subgroups with heavily biased class distributions without an option to manually trade off these aspects (see Section 4.4 for details).
+The final scoring function $\varphi^{rasl}$ is defined relative to the overall dataset: $\varphi^{rasl}(p):=ARL(p)-ARL(\mathcal{\emptyset})$, where the empty pattern $\emptyset$ corresponds to the subgroup containing all instances.
+Although the ARL score is normalized by the number of positive instances, it is heavily influenced by the number of overall instances (of either class) covered by a pattern as the penalty score $PEN(c,p)$ counts the absolute number of violations.
+This favors large subgroups or heavily biased class distributions without an option to manually trade off these aspects (see Section˜4.4 for details).
+In practice, ARL is not a common score to measure classifier performance and the corresponding RASL measure [8] does not provide optimistic estimates for efficient subgroup search.
+ROC and PR AUC, which are widely applied to evaluate soft classifiers, currently are not available for the SCaPE model class.
+However, due to their widespread adoption, they are more readily interpretable and have technical advantages like their independence of the number of evaluated samples (also see Section˜4.4).
+Thus, SubROC introduces ROC and PR AUC for the SCaPE model class and provides optimistic estimates for ROC and PR AUC as well as RASL.
 3 The SubROC Framework
-The SubROC framework consists of three components: defining novel scoring functions based on the widely employed performance measures (ROC AUC and Precision-Recall (PR) AUC) while accounting for class imbalance, size, and redundancy (Section 3.1), deriving optimistic estimates for efficient search space exploration and pruning (Section 3.2), and significance filtering for increasing the robustness and generalizability of the discovered subgroups (Section 3.3).
+The SubROC framework consists of three components: defining novel scoring functions based on ROC AUC, while accounting for class imbalance, size, and redundancy (Section˜3.1), deriving optimistic estimates for efficient search space exploration and pruning (Section˜3.2), and significance filtering for increasing the robustness and generalizability of the discovered subgroups (Section˜3.3).
+The same construction is also applied to PR AUC by replacing ROC ’s true and false positive rates with precision and recall.
 3.1 Scoring Functions
@@
 For PR AUC, we employ the widely used approximation via a linearly interpolated PR curve, with no practical consequences in our initial tests, cf.
-[davis_relationship_2006].
-Subgroup scoring functions commonly include a factor for trading off size against the score to prevent the discovery of only small, overfitted subgroups (e.g., see [klosgen_explora_1996, leman_exceptional_2008]).
+[14].
+Subgroup scoring functions commonly include a factor for trading off size against the score to prevent the discovery of only small, overfitted subgroups (e.g., see [10, 11]).
 We apply the same concept to our AUC-based scoring functions $\varphi^{\prime}$ and derive size-sensitive scoring functions $\varphi_{\alpha}$ as $\varphi_{\alpha}(p)=|sg(p)|^{\alpha}\cdot\varphi^{\prime}(p)$.
@@
 3.1.2 Class Balance Factor
-The performance measures ROC AUC and PR AUC are known to be sensitive to the Negative Class Ratio (NCR) [cook_when_2020, boyd_unachievable_2012].
-This can lead to subgroups with high scores just due to high or low class imbalances rather than exceptional model performance.
-To counterbalance this, we introduce a user-weighted class balance factor $cb(I)$ for our scoring functions [canbek_binary_2017, datta_multiobjective_2019].
+The performance measures ROC AUC and PR AUC are known to be sensitive to the Negative Class Ratio (NCR) [15, 16].
+This can lead to subgroups with high scores just due to high class imbalances rather than exceptional model performance.
+To counterbalance this, we introduce a user-weighted class balance factor $cb(I)$ for our scoring functions [17, 18].
 Given a set of instances $I$, it provides a maximum value close to 1 for perfectly balanced classes and lower values for more imbalanced data.
-We define $cb(I)$ as 00 if only one class occurs and as $\min\{\frac{|\mathcal{N}_{I}|}{|\mathcal{P}_{I}|},\frac{|\mathcal{P}_{I}|}{|\mathcal{N}_{I}|}\}$ otherwise.
+We define $cb(I)$ as 0 if only one class occurs and as $\min\{\frac{|\mathcal{N}_{I}|}{|\mathcal{P}_{I}|},\frac{|\mathcal{P}_{I}|}{|\mathcal{N}_{I}|}\}$ otherwise.
 We combine $cb(I)$ with a scoring function $\varphi^{\prime}$ as an additional factor, forming a class balance weighted scoring function $\varphi_{\beta}(p)$ for a pattern $p$: $\varphi_{\beta}(p)=cb(sg(p))^{\beta}\cdot\varphi^{\prime}(p)$ with $\beta\in\mathbb{R}_{\geq 0}$.
 Scoring functions that include both the cover size weight with parameter $\alpha$ and class balance weight with parameter $\beta$ are denoted as $\varphi_{\alpha,\beta}$.
-This follows the intuition that ROC AUC can be deceptive for imbalanced data, and thus, balanced subgroups should be preferred [saito2015precision].
+This follows the intuition that ROC AUC can be deceptive for imbalanced data, and thus, balanced subgroups should be preferred [19].
 Alternatives such as the difference in class-ratios between the subgroup and the dataset could also be considered.
 3.1.3 Generalization Awareness
-It is a common issue in subgroup discovery that multiple subgroups in the result set have a high pairwise overlap in terms of their cover [lemmerich_difference-based_2013].
-To reduce the intersections of subgroup covers in result sets, there is an adaptation to scoring functions available from literature, called generalization awareness.
-For this, a variant [lemmerich_difference-based_2013] of previous approaches[batal_concise_2010, lemmerich_local_2011, grosskreutz_subgroup_2010] extends a scoring function $\varphi^{\prime}$ by reducing the score $\varphi^{\prime}(p)$ of a subgroup with pattern $p$ by subtracting the maximum score across all its generalizations $h$ (making a specialization viable only if its score is substantially higher):
+It is a common issue in subgroup discovery that multiple subgroups in the result set have a high pairwise overlap in terms of their cover [20].
+To reduce redundancy along the generalization-specialization hierarchy, adaptations to scoring functions have been proposed under the term generalization awareness.
+In this approach, a variant [20] of previous approaches [21, 22, 23] extends a scoring function $\varphi^{\prime}$ by reducing the score $\varphi^{\prime}(p)$ of a subgroup with pattern $p$ by subtracting the maximum score across all its generalizations $h$ (making a specialization viable only if its score is substantially higher):
 $\varphi_{\alpha}(p)=|sg(p)|^{\alpha}\cdot(\varphi^{\prime}(p)-\max_{h\subset p}\varphi^{\prime}(h))\text{, with }\alpha\in[0,1].$
@@
 For runtime optimization, we derive tight optimistic estimates for the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ as well as, separately, for the cover size and class balance factors.
-Proofs are given in Appendix A.
+Proofs are given in Appendix˜A.
 3.2.1 Average Ranking Loss
-To date, no (tight) optimistic estimate was derived for the ARL based scoring function $\varphi^{rasl}$ [duivesteijn_understanding_2014].
-The $ARL(p)$ is defined as the average over the ranking losses $PEN(c,p)$ of all covered positive instances $c\in\mathcal{P}_{p}$.
+To date, no (tight) optimistic estimate was derived for the ARL-based scoring function $\varphi^{rasl}$ [8].
+$ARL(p)$ is defined as the average over the ranking losses $PEN(c,p)$ of all covered positive instances $c\in\mathcal{P}_{p}$.
 The penalty $PEN(c,p)$ for each positive instance $c$ counts the negative instances ranked higher than $c$.
@@
 Next, we introduce a lower bound $b$ for ROC AUC that we will use to derive an optimistic estimate of $\varphi^{rROCAUC}$.
-We distinguish three cases: (a) If perfect separation of ground truth labels is possible with a threshold on the soft predictions, the same holds for every subset.
+We distinguish three cases: (a) If perfect separation of ground truth labels is possible with a threshold on the soft predictions, the same holds for every subset of the cover.
 Thus, the ROC AUC is always 1.
-(b) If the input additionally contains some instances that share the same prediction value (ties) but have different labels (positive and negative), then the ROC AUC on only these tied instances equals $\frac{1}{2}$.
+(b) If the input contains some instances that share the same prediction value (ties) but have different labels (positive and negative), then the ROC AUC on only these tied instances equals $\frac{1}{2}$.
 Any subset of the given instances containing both classes and at least one non-tied instance allows for the correct separation of some positives and negatives by at least one threshold.
 Thus, the ROC AUC is greater than $\frac{1}{2}$ for such subsets.
-(c) The input contains at least one incorrectly ranked pair of a positive and a negative instance.
-The subset containing exactly one such pair has an ROC AUC of 00, which is the lowest possible ROC AUC.
+(c) If the input contains at least one incorrectly ranked pair of a positive and a negative instance.
+The subset containing exactly one such pair has a ROC AUC of 0, which is the lowest possible ROC AUC.
 Lemma 2 (Tight Lower Bound $b_{ROCAUC}$).
@@
 3.2.3 Area Under the Precision-Recall Curve
-Lemma 4 introduces a lower bound for PR AUC which is used in Theorem 5 to derive a tight optimistic estimate for $\varphi^{rPRAUC}$.
-For Lemma 4, a subset $I_{worst}$ of the original set of instances $I$ is constructed to maximize classification errors while minimizing the positive class ratio.
-This makes it impossible for a threshold on any other subset to define a point in PR space below the PR curve of $I_{worst}$ (with some exceptions that have no impact on the AUC of the resulting PR curve).
-Proofs and an illustration are included in Section A.3.
+Lemma˜4 introduces a lower bound for PR AUC which is used in Theorem˜5 to derive a tight optimistic estimate for $\varphi^{rPRAUC}$.
+For Lemma˜4, a subset $I_{worst}$ of the original set of instances $I$ is constructed to maximize classification errors while minimizing the positive class ratio.
+This makes it impossible for a threshold on any other subset to define a point in PR space below the PR curve of $I_{worst}$ (with some exceptions that have no impact on the AUC).
+Proofs and an illustration are included in Section˜A.3.
 Lemma 4 (Tight Lower Bound $b_{PRAUC}$).
@@
 Given a dataset $D=(\mathcal{I},\mathcal{A})$ and a pattern $p$, a tight optimistic estimate for $\varphi^{rPRAUC}(p)$ is $oe_{\varphi^{rPRAUC}}(p)=PRAUC(\mathcal{I})-b_{PRAUC}(sg(p))$.
-3.2.4 Cover Size and Class Balance Weighted Scoring Functions
-The following Lemma 6 gives a tight upper bound for the weighting used in cover size and class balance weighted scoring functions when $\alpha\leq\beta$.
+3.2.4 Cover Size and Class Balance Weighted Scoring
+The following Lemma˜6 gives a tight upper bound for the weighting used in cover size and class balance weighted scoring functions when $\alpha\leq\beta$.
 This result translates to optimistic estimates for $\varphi_{\alpha,\beta}^{rasl}(\cdot)$, $\varphi_{\alpha,\beta}^{rROCAUC}(\cdot)$ and $\varphi_{\alpha,\beta}^{rPRAUC}(\cdot)$.
@@
 3.3 Significance Filtering
-Subgroup discovery inspects an exponential search space of patterns and consequently requires correction for multiple hypothesis comparison.
-To reduce the number of compared subgroups we use a holdout procedure [webb_discovering_2007] splitting analyzed datasets into two parts, the search dataset to search subgroups and the validation dataset to test statistical significance.
+Subgroup discovery inspects an exponential search space of patterns and consequently requires correction for multiple hypothesis comparison [6].
+To reduce the number of compared subgroups, we use a holdout procedure [24]: We split the dataset into two parts, the search dataset to search subgroups and the validation dataset to test statistical significance.
 The tested statistic is the raw score of a subgroup (written $\varphi_{0,0}$) excluding cover size and class imbalance.
@@
 We compute the empirical p-value of a subgroup $p$ by a randomization test (cf.
-[gionis_assessing_2007]), where $1000$ random subsets of the validation dataset are drawn without replacement and with fixed size and class balance.
-$\varphi_{0,0}$ is evaluated on each of the random subsets to construct an empirical distribution of exceptionality.
+[25]), where $1000$ random subsets of the validation dataset are drawn without replacement.
+To prevent significance due to size and class balance, we keep these parameters constant during sampling.
 The empirical p-value is the fraction of random subsets with interestingness $\varphi_{0,0}$ greater than or equal to $\varphi_{0,0}(p)$ divided by the total number of random subsets.
-Multiple testing correction is done using Benjamini-Yekutieli (BY) [benjamini_control_2001] (controlling the false discovery rate (FDR)).
+Multiple testing correction is done using Benjamini-Yekutieli (BY) [26] (controlling the false discovery rate (FDR)).
 BY has the necessary properties to be applied under any dependency between tested hypotheses.
 The correction approach can be customized by the user, e.g., with Bonferroni controlling familywise error rate (FWER).
-The default p-value threshold we use is $p\leq 0.05$.
+The default p-value threshold we use is $p\leq\alpha$.
+Here $\alpha=0.05$.
+This enables significance filtering, focusing the search on significant subgroups: To find top-$k$ significant subgroups (e.g., $k=5$), search for the top-$k^{\prime}$ subgroups with $k^{\prime}\gg k$ (e.g., $k^{\prime}=100$), calculate p-values, and only include a subgroup in the final top-5 if their corrected p-value passes the significance threshold $\alpha$.
 4 Experiments
-Datasets are from OpenML [vanschoren_openml_2013] and the UCI Machine Learning Repository [kelly_uci_nodate] (statistics in Supp.
-Table 3).
-Each was partitioned randomly into four equal parts for training, subgroup discovery, significance testing, and result evaluation.
-4.1 Quality of SubROC Results Across Multiple Datasets
-To illustrate the quality of subgroups discovered with SubROC, we applied xgboost on various datasets and searched for exceptionally underperforming subgroups with regard to the performance measures ROC AUC, PR AUC, and ARL.
-Table 1 compares two settings: a basic SubROC setting (b) and the full SubROC framework (f).
-For b, the search solely uses the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$, and $\varphi^{rPRAUC}$.
-The top-5 subgroups are returned and additionally tested for significance.
-The full framework (f) is configured with size and class imbalance weighting $\alpha=\beta=1$, generalization awareness, and statistical significance filtering.
-Here, the search returns the top-100 subgroups, which were filtered for significance, and finally, the top-5 subgroups are returned.
-In both cases, we used a minimum subgroup size of $\geq 20$, and a maximum pattern length of 4.
-Statistics were calculated for top-5 results.
+Datasets are from OpenML [27] and the UCI Machine Learning Repository [28] (statistics in Table˜4).
+Each dataset was partitioned randomly into three equal parts: (i) a training set for fitting the prediction model; (ii) a search set where the trained model is applied to yield predictions $\hat{Y}$ on unseen data and the subgroup discovery is performed; (iii) a validation set for significance testing of the found subgroups (see Section˜3.3).
+4.1 Quality of SubROC Across Multiple Datasets
+Table 1: Overview of top-5 results from a basic SubROC setting (b) and the full SubROC framework (f) for a range of datasets and performance measures.
+Full SubROC managed to find diverse and exceptional subgroups, which are often much larger and balanced than the baseline.
+Numbers are bold for the better result when comparing b and f.
+(Cont.
+in Table˜5, containing results for more datasets)
+Mean Cover Size
+Significant / Top-5
+Filtered-5 / Significant / Top-100
+Mean Exception- ality
+Mean NCR
+Mean Pairwise IoU
+Dataset Metric     b f     b f     b f     b f     b f
+Census n=199523 NCR=0.06 ROC     25 6753    1 / 5 5 / 99 / 100    0.95 0.05    0.04 0.16    0.10 0.31
+PR     23 864    0 / 5 5 / 99 / 100    0.92 0.14    0.93 0.45    0.37 0.45
+ARL     8000 11945    5 / 5 5 / 27 / 27    47 31    0.15 0.11    0.99 0.50
+Adult n=48842 NCR=0.24 ROC     27 6492    0 / 5 5 / 77 / 100    0.92 0.06    0.04 0.37    0.48 0.61
+PR     27 2803    2 / 5 5 / 70 / 100    0.94 0.08    0.96 0.46    0.37 0.19
+ARL     4614 6930    5 / 5 5 / 5 / 5    148 55    0.42 0.34    0.85 0.61
+Bank n=45211 NCR=0.88 ROC     22 311    0 / 5 5 / 22 / 100    0.83 0.16    0.95 0.48    0.10 0.12
+PR     157 72    0 / 5 2 / 2 / 100    0.68 0.10    0.99 0.48    0.80 0.51
+ARL     - -    0 / 0 0 / 0 / 0    - -    - -    - -
+Credit n=30000 NCR=0.78 ROC     24 1506    0 / 5 5 / 38 / 100    0.75 0.12    0.96 0.61    0.01 0.09
+PR     75 1895    0 / 5 5 / 17 / 100    0.51 0.29    0.99 0.85    0.18 0.34
+ARL     - -    0 / 0 0 / 0 / 0    - -    - -    - -
+Approval n=690 NCR=0.55 ROC     29 -    0 / 5 0 / 0 / 41    0.64 -    0.96 -    0.69 -
+PR     41 80    2 / 5 1 / 1 / 25    0.89 0.85    0.98 0.95    0.75 -
+ARL     60 87    3 / 5 3 / 3 / 3    26 10    0.96 0.79    0.68 0.69
+Table 2: Example filtered-5 result set from the full SubROC framework using ROC AUC on Adult data with xgboost predictions.
+Pattern ROC AUC Cover NCR
+$\text{marital-status}=\text{Married-civ-spouse}$
+0.8508
+5280
+0.4496
+$\text{relationship}=\text{Husband}$
+0.8485
+4638
+0.4495
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}$
+0.8053
+4628
+0.4114
+$\text{sex}=\text{Male}$
+0.9053
+7581
+0.3109
+$\text{capital-gain}\in[0.0,114.0)$
+0.9015
+10331
+0.2126
+$\emptyset$
+0.9241
+11305
+0.2482
+To illustrate the effectiveness of the full SubROC framework, we compare a basic SubROC setting (b) and the full SubROC framework (f) for finding the top-5 underperforming subgroups (Table˜1).
+Both cases train xgboost [29] as the binary classifier on (i), chosen because it is a widely adopted, strong baseline method.
+Subgroups are then exhaustively [^2] Results are the same for any non-heuristic (exhaustive) search algorithm.
+searched on (ii) with a minimum subgroup size of 20, and a maximum pattern length of 4.
+The basic setting (b) solely uses the bare scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$, and $\varphi^{rPRAUC}$ and returns the top-5 subgroups.
+Their significance is tested on (iii) to underpin their tendency for random artifacts (e.g.
+very small subgroups, high class imbalance), see column Significant / Top-5.
+The full framework (f) is configured with size and class imbalance weighting $\alpha=\beta=1$, generalization awareness, and statistical significance filtering ($k^{\prime}=100$).
+To illustrate the effectiveness of significance filtering, we report how many subgroups were found for a top-100 search on (ii), how many of these passed the statistical test on (iii), and how many finally remain after significance filtering (see column Filtered-5 / Significant / Top-100).
+Statistics were calculated for the final top-5 results on (ii) in both cases.
 Note that exceptionality should not be compared between different performance measures because of their differing scales and responses to errors.
 For b, cover sizes tend to be close to the minimum cover size limit, and class ratios tend to be close to 0 or 1.
-Despite their high difference in performance compared to the overall dataset (exceptionality), such subgroups do not necessarily point towards generalizable structural issues of the classifier but may be caused by artifacts (cf.
-Section 4.4).
+Despite their high difference in performance compared to the overall dataset (exceptionality), such subgroups do not necessarily point towards structural issues of the classifier but may be caused by artifacts (cf.
+Section˜4.4).
 This is also apparent in the low number of subgroups passing statistical significance in the top-5 results of b.
-Similar patterns are observable in the top-100 results (see Supp.
-Table 5).
+Similar patterns are observable in the top-100 results (see Table˜6).
 While yielding less prominent exceptionality, we observe that the SubROC framework leads to larger and more balanced results.
-Additionally, substantially more subgroups are returned that pass the significance filtering step.
-This is apparent in the top-5 as well as the top-100 result (cf.
-Supp.
-Table 5).
-We also inspect the pairwise Intersection over Union (IoU) of the discovered top-5 subgroups.
-A higher IoU refers to more overlapping and, thus, redundant subgroups.
-Values generally remain in a similar range to the baseline, but we observe occasional significant improvements by SubROC.
-This indicates that our framework does not come at a cost of and in some cases improves result diversity.
-The statistical filtering step of SubROC saves resources for practitioners by discarding a considerable amount of initially found, but unreliable subgroups.
-This effect is particularly prominent for small datasets with less than $2,500$ search instances, where the significance filtering discarded almost all subgroups.
+Additionally, returned subgroups are substantially more likely to pass the significance filtering step.
+Mean pairwise Intersection over Union (IoU) of the discovered subgroups was computed to assess redundancy.
+Values for SubROC remain comparable to the baseline, with some significant improvements, indicating that our framework does not sacrifice diversity.
+The significance filtering step of SubROC saves resources for practitioners by discarding a considerable amount of initially found, but unreliable subgroups.
+This effect is particularly prominent for small datasets with less than $2,500$ instances in the search set, where the significance filtering discarded almost all subgroups.
 It is less prominent for ARL based scoring functions, which generally produce smaller result sets with a greater share of significant subgroups.
-Additional results can be found in Section B.2.
-For concrete result sets from a case study on the Adult dataset see Section B.3.
-Table 1: Overview of top-5 result set properties from a basic SubROC setting (b) and the full SubROC framework (f) for a range of datasets and performance measures.
-SubROC managed to find diverse and exceptional subgroups, which are often much larger and balanced than the baseline.
-(Continued)
-Mean Cover Size
-Significant/ Top-5
-Top-5/ Significant/ Original
-Dataset Metric     b f     b f
-Census n=199523 NCR=0.06 ROC     25 6753    1/5 5/99/100
-PR     23 864    0/5 5/99/100
-ARL     8000 11945    5/5 5/27/27
-Adult n=48842 NCR=0.24 ROC     27 6492    0/5 5/77/100
-PR     27 2803    2/5 5/70/100
-ARL     4614 6930    5/5 5/5/5
-Bank n=45211 NCR=0.88 ROC     22 311    0/5 5/22/100
-PR     157 72    0/5 2/2/100
-ARL     - -    0/0 0/0/0
-Credit n=30000 NCR=0.78 ROC     24 1506    0/5 5/38/100
-PR     75 1895    0/5 5/17/100
-ARL     - -    0/0 0/0/0
-Approval n=690 NCR=0.55 ROC     29 -    0/5 0/0/41
-PR     41 80    2/5 1/1/25
-ARL     60 87    3/5 3/3/3
-Table 1: Continued from previous columns. (Rows continued in Supp. Table 4, containing results for more datasets)
-Mean Exceptionality
-Mean NCR
-Mean Pairwise IoU
-Dataset Metric     b f     b f     b f
-Census n=199523 NCR=0.06 ROC     0.95 0.05    0.04 0.16    0.10 0.31
-PR     0.92 0.14    0.93 0.45    0.37 0.45
-ARL     47 31    0.15 0.11    0.99 0.50
-Adult n=48842 NCR=0.24 ROC     0.92 0.06    0.04 0.37    0.48 0.61
-PR     0.94 0.08    0.96 0.46    0.37 0.19
-ARL     148 55    0.42 0.34    0.85 0.61
-Bank n=45211 NCR=0.88 ROC     0.83 0.16    0.95 0.48    0.10 0.12
-PR     0.68 0.10    0.99 0.48    0.80 0.51
-ARL     - -    - -    - -
-Credit n=30000 NCR=0.78 ROC     0.75 0.12    0.96 0.61    0.01 0.09
-PR     0.51 0.29    0.99 0.85    0.18 0.34
-ARL     - -    - -    - -
-Approval n=690 NCR=0.55 ROC     0.64 -    0.96 -    0.69 -
-PR     0.89 0.85    0.98 0.95    0.75 -
-ARL     26 10    0.96 0.79    0.68 0.69
-4.2 Optimistic Estimates Substantially Speed Up Runtimes
-To investigate the efficiency improvements enabled by SubROC, we measure runtimes of subgroup discovery with and without pruning based on the introduced optimistic estimates 10 times for each dataset (3 times for Census due to long runtimes).
-The searches run on predictions from a Gaussian Naive Bayes classifier, using the best-first-search enumeration strategy, with a maximum pattern length of 4.
-Different size and class imbalance weightings ($\alpha,\beta$) are compared.
-Table 2 gives an overview of the measured speedups in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning.
-Absolute runtimes and experiments with different pattern lengths are reported in Section B.5.
-The results show substantial runtime reductions.
-The most prominent improvements were observed for ARL with speedups up to $538.9$ on the Credit dataset, reducing runtime from $\sim 1.25$ hours to less than 10 seconds.
-For PR AUC, the highest speedup of $22.6$ was also on Credit with runtime reductions from $\sim 1.25$ hours to $\sim 4$ minutes.
-Finally, for ROC AUC, a speedup of $101.0$ on Census was the highest with an improvement from $\sim 20$ hours to $\sim 12$ minutes.
-Overall, larger datasets (both in terms of number of instances and features) benefit the most from optimistic estimates across all performance measures.
-For the smallest datasets (with generally small runtimes) like Approval, where the search split contains less than 250 instances, the speedups substantially decrease across all measures, leading to an overall slowdown in some cases.
-With regard to parameterization, a very low emphasis on size and class imbalance compared to the performance measure ($\alpha=\beta=0$) as well as a high emphasis ($\alpha=\beta=1$) seems to produce the highest speedups.
-Further investigations are required, however, an explanation for this may be that the search can more strongly focus on one or the other optimistic estimate factor
-Table 2: Speedups of median runtimes using optimistic estimates.
-Speedups are given in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning.
-Substantial runtime reductions were achieved on all but the smallest dataset.
+Table˜2 shows one of the result sets from this experiment. For more full result sets on the Adult dataset, see Section˜B.3.
+4.2 Runtime Speed Ups with Optimistic Estimates
+Table 3: Speedups of median runtimes using optimistic estimates, given in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning.
 (Cont.
-in Supp.
-Table 12, containing results for more datasets)
+in Table˜13, containing results for more datasets)
 Performance Measure
@@
 Approval n/m=690/15
-ARL 0    18.4 47.4 9.7 538.9 1.7
-0.1    18.6 39.4 7.7 385.6 0.8
-0.3    19.9 40.4 7.6 383.1 0.7
-1    23.3 44.7 7.6 388.4 0.6
-PR AUC 0    2.9 2.3 1.7 2.5 1.2
-0.1    1.9 1.2 0.7 1.1 0.4
-0.3    2.7 2.1 0.8 2.9 0.4
-1    11.4 19.9 2.2 22.6 0.4
-ROC AUC 0    101.0 0.9 2.1 83.8 0.8
-0.1    1.0 0.8 0.8 1.0 0.5
-0.3    1.3 1.0 0.9 1.1 0.5
-1    6.4 8.4 2.1 3.9 0.5
-4.3 Discovering Injected Subgroups
-To investigate the validity of SubROC and its ability to discover expected subgroups, we injected an artificially underperforming subgroup into a dataset based on the predictions of an xgboost classifier on the Adult dataset.
-The injected subgroup was selected randomly from all subgroups with pattern length $\leq 3$ in a cover size range from $0.4\%$ to $0.6\%$ of the search split.
-It covers 54 instances, has an NCR of 0.2, and an ARL of 1.07, PR AUC of 0.97, and ROC AUC of 0.90 before and 9.93, 0.62, and 0.1, respectively, after injection.
-For the injection, we multiplied the predicted scores of instances in the subgroup cover by $-1$.
-Thereby, the injected subgroup forms a known weakness in the predictions, and we can evaluate the subgroup search by whether this subgroup is included in the result set.
-The subgroup discovery was performed up to depth 4 under different configurations (baseline without the SubROC framework and three different weightings of cover size and class balance in the SubROC framework).
-Result sets were top-10 filtered to reflect practically interpretable result set sizes.
-Figure 2 shows the IoU as an overlap measure between found subgroups and the injected subgroup for the ROC AUC.
-In contrast to the baseline, the SubROC framework successfully discovers the injected subgroup and reports it at the highest ranks in the result set.
-Note that multiple patterns can describe the same set of instances, so the injected errors can be found several times.
-While ROC AUC robustly identifies the subgroup, similar experiments on ARL or the PR AUC did not reproduce these findings.
-The result sets contain only subgroups with IoU s with the injected subgroup below $0.1$.
-See Section B.4 for detailed results.
-Differences in terms of cover size and class balance skew (see Section 4.4) between the performance measures have a large influence on the results of these experiments.
-The injected subgroup contains more negatives than positives, which interacts with the class balance skews of ARL and PR AUC.
-Furthermore, the injected subgroup is small, making it inherently hard to capture with ARL, which has an embedded size skew.
-This highlights the importance of the size and class balance terms introduced in SubROC to counter these skews, but also shows that they do not completely remove the effects.
-Future work, for example, employing a normalized PR AUC [boyd_unachievable_2012] may be promising.
-Figure 2: Intersection over Union (IoU) of an injected subgroup with subgroups in a top-10 ROC AUC result set, ordered decreasingly by interestingness.
+ARL 0    9.1 11.6 482.0 10.2 49.3
+0.1    9.3 11.1 333.9 9.3 24.7
+0.3    10.4 13.0 337.6 9.5 24.6
+1    15.0 27.5 333.3 9.3 25.4
+PR AUC 0    2.0 1.5 460.7 3.3 43.8
+0.1    2.0 1.2 337.2 1.7 25.4
+0.3    2.0 1.2 338.0 1.6 24.6
+1    3.8 5.3 337.1 2.2 24.9
+ROC AUC 0    1.5 1.1 401.7 1.7 41.7
+0.1    1.5 1.0 332.8 1.6 17.0
+0.3    1.4 1.0 303.2 1.6 19.5
+1    2.8 4.4 307.4 1.8 24.2
+To investigate the efficiency improvements enabled by SubROC, we measure runtimes of subgroup discovery with and without pruning based on our optimistic estimates, carrying out 10 repetitions for each dataset (3 repetitions for Census due to long runtimes).
+The searches for the top-5 patterns run on predictions from an xgboost classifier, using the best-first-search enumeration strategy, with a maximum pattern length of 4.
+Different size and class balance weightings ($\alpha,\beta$) are compared.
+Table˜3 gives an overview of the measured speedups.
+Absolute runtimes are given in Section˜B.5.
+The results show substantial runtime reductions.
+Particularly large speedups were achieved on the Bank dataset.
+For ARL, pruning cut the runtime from $\sim 2.13$ min to <0.3<0.3 s ($\sim 482\times$ speedup).
+For PR AUC, the runtime dropped from $\sim 2.42$ min to $0.3$ s ($\sim 461\times$ speedup).
+For ROC AUC, pruning reduced the runtime from $\sim 2.67$ min to $0.4$ s ($\sim 402\times$ speedup).
+With regard to parameterization, both ignoring size and class balance altogether ($\alpha=\beta=0$) and weighting them equally to the performance measure ($\alpha=\beta=1$) appear to produce the highest speedups.
+Further investigations are required, however, an explanation for this may be that the search can more strongly focus on one or the other optimistic estimate factor.
+4.3 Discovering an Injected Subgroup
+(a) baseline
+(b) $\alpha=0,\beta=0$
+(c) $\alpha=0.3,\beta=0$
+(d) $\alpha=0,\beta=0.3$
+Figure 2: Intersection over Union (IoU) of subgroup covers, comparing an injected subgroup with subgroups in a top-10 ROC AUC result set, ordered decreasingly by interestingness.
 “baseline” was obtained searching in a base SubROC setting.
@@
 The full SubROC framework accurately found the injected subgroup for any weighting configuration by returning two highly overlapping subgroups in the highest ranks of all result sets.
+To investigate the validity of SubROC and its ability to discover expected subgroups, we injected an artificially underperforming subgroup into the Adult dataset based on the predictions of an xgboost classifier.
+The subgroup was randomly drawn from candidates with pattern length $\leq 3$ whose cover size fell between $0.4\%$$0.6\%$ of the search set.
+It covers 54 instances, has an NCR of 0.2, and an ARL of 1.07, PR AUC of 0.97, and ROC AUC of 0.90 before and 9.93, 0.62, and 0.1, respectively, after injection.
+For the injection, we multiplied the predicted scores in the subgroup by 1-1.
+Thereby, the injected subgroup forms a known weakness in the predictions, and we can evaluate the subgroup search by whether this subgroup is included in the result set.
+The subgroup discovery was performed up to depth 4 under different configurations (baseline without the SubROC framework and three different weightings of cover size and class balance in the SubROC framework).
+Result sets were top-10 filtered to reflect interpretable result set sizes.
+Figure˜2 shows the IoU as an overlap measure between found subgroups and the injected subgroup for the ROC AUC.
+In contrast to the baseline, the SubROC framework successfully discovers the injected subgroup and reports it at the highest ranks in the result set.
+Note that multiple patterns can describe the same set of instances, so the injected errors can be found several times.
+While ROC AUC robustly identifies the subgroup, similar experiments on ARL or the PR AUC did not reproduce these findings (all returned $\text{IoU}<0.1$; see Section˜B.4).
+The injected subgroup contains mostly negatives, which interacts with the class balance skews of ARL and PR AUC (see Section˜4.4.2).
+Furthermore, the injected subgroup is small, making it inherently hard to capture with ARL, which has an embedded size skew (see Section˜4.4.1).
+This highlights the importance of the size and class balance terms introduced in SubROC to counter these skews, but also shows that they do not completely remove the effects.
+Future work, for example, employing a normalized PR AUC [16] may provide further improvement.
 4.4 Importance of Cover Size and Class Imbalance Weighting
-To illustrate the effect of the interaction between model performance, size and class ratio on the proposed scoring functions, we conduct synthetic experiments (cf.
-[cook_when_2020]).
-Data was generated from a bivariate normal distribution with means 00 and variances $1$.
+To illustrate the effect of the interaction between model performance, size and class ratio on the proposed scoring functions, we conduct experiments on synthetic data (cf.
+[15]).
+Data was generated from a bivariate normal distribution with means 0 and variances $1$.
 The first variable $z$ was used to derive a ground truth label, the second $\hat{z}$ represents the prediction values.
 Varying their correlation thus varies the “prediction performance” (the $y$-axes).
-Dataset sizes (cover of a subgroup) are varied in Figures 3(a), 3(b), 3(e), 3(f), 3(i) and 3(j) ($x$ axis) and set to $100$ in Figures 3(c), 3(d), 3(g), 3(h), 3(k) and 3(l).
-Class balance (NCR) $q$ was varied by thresholding $z$ at the first $q$-quantile in Figures 3(c), 3(d), 3(g), 3(h), 3(k) and 3(l) ($x$ axis) and fixed at $q=0.5$ for Figures 3(a), 3(b), 3(e), 3(f), 3(i) and 3(j).
-Since scoring functions are relative to the overall dataset, we equivalently derive a dataset with correlation 00, size $100$ and $q=0.5$.
+Dataset sizes (cover of a subgroup) are varied in Figure˜3 ($x$ axis) and set to $100$ in Figure˜4.
+Class balance (NCR) $q$ was varied by thresholding $z$ at the first $q$-quantile in Figure˜4 ($x$ axis) and fixed at $q=0.5$ for Figure˜3.
+Since scoring functions are relative to the overall dataset, we equivalently derive a dataset with correlation 0, size $100$ and $q=0.5$.
 The generation process was repeated $20$ times followed by an averaging of scores to reduce noise.
-Note that this procedure allows to evaluate practically impossible subgroups.
-When the cover size or NCR of a subgroup is close to that of the overall dataset, then the exceptionality (i.e.
-performance difference) is limited in practice.
+Note that this procedure allows to evaluate practically impossible subgroups, since any subgroup with cover size or NCR close to that of the overall dataset can only have limited exceptionality (i.e.
+performance difference) in practice.
 4.4.1 Skew from Cover Size
-Figure 3(a) shows the dependence of $\varphi^{rasl}$ on the number of instances in its input, yielding larger values for larger inputs even if predictive performance is fixed constant (see the horizontal gradient at fixed correlations).
+Figure˜3(a) shows the dependence of $\varphi^{rasl}$ on the number of instances in its input, yielding larger values for larger inputs even if predictive performance is kept constant (see the horizontal gradient at fixed correlations).
 This can lead to low values of $\varphi^{rasl}$ on small subgroups, which do not necessarily correspond to a performance improvement over the complete dataset.
-Underperforming subgroups can therefore not be identified by a fixed threshold on the metric (e.g., $\geq 0$).
-This means that inherently smaller subgroups have to have a higher error rate to be considered important for ARL, and thus the approach returns results with a higher rate of passing the statistical test (Table 1).
-However, the strength of this effect can not be influenced by the user.
-Neither $\varphi^{rPRAUC}$ nor $\varphi^{rROCAUC}$ change with cover size (3(e), 3(i)).
+Underperforming subgroups therefore cannot be identified by a fixed threshold on the metric (e.g., $\geq 0$).
+Therefore smaller subgroups have to have a higher error rate to be considered important for ARL, and thus the approach returns results with a higher rate of passing the statistical test (Table˜1).
+However, the strength of this effect can not be controlled by the user.
+Neither $\varphi^{rPRAUC}$ nor $\varphi^{rROCAUC}$ change with cover size (3(c), 3(e)).
 Furthermore these figures show increased noise on the lower end of the cover size axes, indicating increased random effects on these scoring functions for subgroups with few instances which underlines the necessity of the statistical tests introduced by SubROC.
-Additionally, Figures 3(f) and 3(j) show that the introduced cover size weighting successfully leads to a dependence on input size, encoding a preference for larger inputs for both of these scoring functions.
-4.4.2 Skew from Class Ratio
-Figures 3(c) and 3(k) show that the ARL as well as the PR AUC depend on the class ratio.
-The PR AUC skew is known to be related to the unachievable area in PR space [boyd_unachievable_2012].
-This skew has a similar influence on the interpretation of the scores as has the cover size skew of the ARL.
-Figure 3(g) shows a more subtle effect for the ROC AUC through its contour lines, that was also found by [cook_when_2020].
-Figures 3(d), 3(h) and 3(l) show that the class balance weighting successfully introduces a preference towards subgroups with balanced classes.
-In particular, the score of subgroups with highly imbalanced classes is close to zero.
-Because cover size also has only a minor effect on the ROC AUC (cf.
-Figure 3(e)), searches utilizing the ROC AUC within SubROC can be expected to yield the most consistently underperforming subgroups over the widest range of cover sizes and class balances compared to the ARL and the PR AUC.
+Additionally, Figures˜3(d) and 3(f) show that the introduced cover size weighting leads to a dependence on input size, encoding a preference for larger inputs for both of these scoring functions.
 (a) $\varphi^{rasl}_{0,0}$
 (b) $\varphi^{rasl}_{1,0}$
-(c) $\varphi^{rasl}_{0,0}$
-(d) $\varphi^{rasl}_{0,1}$
-(e) $\varphi^{rROCAUC}_{0,0}$
-(f) $\varphi^{rROCAUC}_{1,0}$
-(g) $\varphi^{rROCAUC}_{0,0}$
-(h) $\varphi^{rROCAUC}_{0,1}$
-(i) $\varphi^{rPRAUC}_{0,0}$
-(j) $\varphi^{rPRAUC}_{1,0}$
-(k) $\varphi^{rPRAUC}_{0,0}$
-(l) $\varphi^{rPRAUC}_{0,1}$
-Figure 3: Skewed response of cover size and class balance weighted scoring functions depending on cover size, NCR and correlation (model performance) on synthetic data.
+(c) $\varphi^{rROCAUC}_{0,0}$
+(d) $\varphi^{rROCAUC}_{1,0}$
+(e) $\varphi^{rPRAUC}_{0,0}$
+(f) $\varphi^{rPRAUC}_{1,0}$
+Figure 3: Skewed response of cover size and class balance weighted scoring functions depending on cover size and correlation (model performance) on synthetic data.
 Contour lines show smoothed paths along which interestingness is equal.
+Unweighted scoring functions based on the AUC performance measures show increased noise on subgroups with small covers.
+Significant dependencies on the cover size are visible in the results for the unweighted scoring function based on the ARL performance measure as well as all weighted scoring functions.
+(a) $\varphi^{rasl}_{0,0}$
+(b) $\varphi^{rasl}_{0,1}$
+(c) $\varphi^{rROCAUC}_{0,0}$
+(d) $\varphi^{rROCAUC}_{0,1}$
+(e) $\varphi^{rPRAUC}_{0,0}$
+(f) $\varphi^{rPRAUC}_{0,1}$
+Figure 4: Skewed response of cover size and class balance weighted scoring functions depending on negative class ratio and correlation (model performance) on synthetic data.
+Contour lines show smoothed paths along which interestingness is equal.
+The unweighted scoring functions based on the ARL and the PR AUC both increase with the NCR for almost any correlation.
+The unweighted scoring function based on the ROC AUC depends on the difference between the NCR and $0.5$.
+All weighted scoring functions return the most extreme values close to NCR s of $0.5$, with a gradient from low to high scores as correlations decrease.
+4.4.2 Skew from Class Ratio
+Figures˜4(a) and 4(e) show that the ARL and PR AUC depend on the class ratio.
+The PR AUC skew is known to be related to the unachievable area in PR space [16].
+This skew has a similar influence on the interpretation of the scores as has the cover size skew of the ARL.
+Figure˜4(c) shows a more subtle effect for the ROC AUC through its contour lines, that was also found by [15].
+Figures˜4(b), 4(d) and 4(f) show that the class balance weighting introduces a preference towards subgroups with balanced classes.
+In particular, the score of subgroups with highly imbalanced classes is close to zero.
+Since ROC AUC exhibits the least prominent skews (cf.
+Figures˜3(c) and 4(c)), searches utilizing the ROC AUC within SubROC can be expected to yield the most consistently underperforming subgroups over the widest range of cover sizes and class balances compared to the ARL and PR AUC.
 5 Related Work
 We summarize work on producing descriptions for parts of a dataset where a soft classifier over- or underperforms and describe the differences to our approach.
-[sagadeeva_sliceline_2021] only consider residuals-based loss functions to rank subgroups in contrast to our ranking-based measures.
-Similarly, [zhang_identifying_2017] does not use AUC metrics or the subgroup discovery framework.
-Other research allows for user-specified performance measures for subgroup ranking, but they do not provide specific support such as optimistic estimates or generalization-awareness[chung_automated_2020, zhang_sliceteller_2017].
-[cabrera_fairvis_2019, chung_automated_2020, zhang_sliceteller_2017] propose automatic search approaches as part of generic interactive evaluation tools.
+Sagadeeva and Boehm [30] only consider residuals-based loss functions to rank subgroups in contrast to our ranking-based measures.
+Similarly, Zhang and Neill [31] use neither AUC metrics nor the subgroup discovery framework.
+Other research allows for user-specified performance measures for subgroup ranking, but they do not provide specific support such as optimistic estimates or generalization-awareness[6, 32].
+[7, 6, 32] propose automatic search approaches as part of generic interactive evaluation tools.
 None of these works applied the difference between the subgroup AUC to the overall dataset AUC for the ROC or PR curve in a subgroup discovery setting and investigated its properties in detail.
 The following works are concerned with the ROC curve of subgroups in detail.
-[gardner_evaluating_2019] propose to use the area between the ROC curves of a model on two subgroups in the context of fairness of data slicing but do not discuss the mining of respective patterns.
-[carrington_deep_2023] propose to evaluate specific parts of the ROC curve and call these parts subgroups.
+[33] propose to use the area between the ROC curves of a model on two subgroups in the context of fairness of data slicing but do not discuss the mining of respective patterns.
+[34] propose to evaluate specific parts of the ROC curve and call these parts subgroups.
 Considered parts in that work are intervals in the x-axis or the y-axis of the ROC space as well as intervals in the thresholds used to compute the ROC points from.
 In addition to this differing notion of subgroups, they do not investigate an automated search for subgroups.
-[tafvizi_attributing_2022] presents methods for computing the attribution of subgroups to the overall AUC.
+[35] presents methods for computing the attribution of subgroups to the overall AUC.
 They do this to identify parts of the data that contribute to a reduction in the overall AUC.
-The contribution of a subgroup to the overall dataset AUC is highly influenced by how the subgroup is embedded into the ordering of the overall dataset.
+The contribution of a subgroup to the overall dataset AUC is highly influenced by where its instances fall within the global ranking of predictions for the entire dataset.
 Therefore finding parts of the dataset with an exceptional AUC would simply not contribute to their goal.
-Our goal is to find under which circumstances a model is most likely to fail or succeed, so we take a different approach with AUC-based exceptionality.
-For subgroup discovery, many (tight) optimistic estimates (OE) have been proposed, primarily for settings with traditional target concepts (binary, numeric, or both) [grosskreutz_fast_2010, lemmerich_fast_2016].
-In a few settings, tight OE have also been derived for more complex model classes in the EMM framework [kalofolias_discovering_2019, belfodil_identifying_2020].
-An OE for a scoring function that computes the ROC AUC is given in [lemmerich_fast_2016].
-However, there the ROC curve is based on different attributes than what is studied in this work, such that the OE is not applicable here.
-To the best of our knowledge, none of the previously presented OEs can be applied to the scoring functions used in this work.
+Our goal is to find under which circumstances a model is most likely to fail, so we take a different approach with AUC-based exceptionality.
+For subgroup discovery, many (tight) optimistic estimates (OE) have been proposed, primarily for settings with traditional target concepts (binary, numeric, or both) [36, 37].
+In a few settings, tight OE have also been derived for more complex model classes in the EMM framework [38, 39].
+An OE for a scoring function that computes the ROC AUC is given in [37].
+However, there the ROC curve is based on different attributes compared to ours, such that the OE is not applicable here.
+To the best of our knowledge, no previously presented OE can be applied to the scoring functions used in this work.
 6 Conclusion
-We introduced SubROC, a novel framework for reliable and efficient discovery of interpretable subgroups with exceptional model performance with respect to soft classifiers.
+We introduced SubROC, a novel framework for reliable and efficient discovery of interpretable subgroups with exceptional model performance for soft classifiers.
 Our experiments demonstrate the necessity for such a framework as well as the effectiveness of SubROC when using the performance measures ARL, ROC AUC and PR AUC.
@@
 By providing optimistic estimates, we enable fast mining.
-To summarize, SubROC has the potential to enable routine evaluation of ML models, improve quality control procedures, inform model performance enhancements, and increase confidence in deployed ML models overall.
-In this work, we focused on ranking-based soft classifier evaluation measures due to their widespread use and their specific challenges to Exceptional Model Mining.
-Other popular performance measures, such as cross-entropy or the Brier score, can be integrated into classic subgroup discovery.
+To summarize, SubROC has the potential to enable routine evaluation of ML models, improve quality control procedures, inform model performance enhancements, and increase confidence in ML models.
+In this work, we focused on ranking-based soft classifier evaluation measures given their widespread use and the specific challenges they pose for integration into the Exceptional Model Mining framework.
+By contrast, other popular performance measures, such as cross-entropy or the Brier score decompose over individual instances and can therefore be more directly integrated into classic subgroup discovery.
 Their in-depth evaluation is an avenue for future research.
@@
 7 Acknowledgements
-The funding for this study was in part provided by the BMBF (01IS22077).
-[…]
+Funding was provided by the BMFTR (01IS22077, 01ED2507).
 Appendix A Proofs of Tight Optimistic Estimates
@@
 Here soft classifier performance measures are functions that assign a real value to ground truth and prediction pairs of soft classifiers.
-Our formalizations make use of standard multiset notation as presented in [syropoulos_mathematics_2001].
+Our formalizations make use of standard multiset notation as presented in [40].
 To handle values of instances in a dataset $D=(\mathcal{I},\mathcal{A})$ we define a database as the multiset $d(D)=[(A_{1}(c),\dots,A_{|\mathcal{A}|}(c))|c\in\mathcal{I}]$.
@@
 Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $m_{l}$ be a loss-type soft classifier performance measure. The soft classifier loss interestingness measure for $m_{l}$ is defined as
-$\varphi^{m_{l}}(p)=m_{l}(sg(p)^{\{y,\hat{y}\}})$
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 Let $D=(\mathcal{I},\mathcal{A})$ be a dataset. Let $m_{s}$ be a score-type soft classifier performance measure. The soft classifier score interestingness measure for $m_{s}$ is defined as
-$\varphi^{m_{s}}(p)=-m_{s}(sg(p)^{\{y,\hat{y}\}})$
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 The relative soft classifier loss interestingness measure is defined as
-$\varphi^{rm_{l}}(p)=\varphi^{m_{l}}(p)-m_{l}(\mathcal{I}^{\{y,\hat{y}\}})$
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 The relative soft classifier score interestingness measure is defined as
-$\varphi^{rm_{s}}(p)=\varphi^{m_{s}}(p)+m_{s}(\mathcal{I}^{\{y,\hat{y}\}})$
+[…]
 for any pattern $p$ that is defined on $D$.
-The main interestingness measure $\varphi^{rasl}$ from [duivesteijn_understanding_2014] is defined with alternative signatures (compared to the main text) for use in the proofs as
+The main interestingness measure $\varphi^{rasl}$ from [8] is defined with alternative signatures (compared to the main text) for use in the proofs as
 $PEN_{i}([c_{1},\dots,c_{m}]):=$ $\sum_{j=1}^{m}(\mathbbm{1}_{\{(y,\hat{y})\mid y=0\wedge\hat{y}>\hat{y}_{i}\}}(y_{j},\hat{y}_{j})$
@@
 $ARL([c_{1},\dots,c_{m}]):=$ $\frac{\sum_{i=1}^{m}\mathbbm{1}_{\{1\}}(y_{i})\cdot PEN_{i}([c_{1},\dots,c_{m}])}{\sum_{i=1}^{m}\mathbbm{1}_{\{1\}}(y_{i})}$
-$\varphi^{rasl}(p):=$ $ARL(sg(p)^{\{y,\hat{y}\}})-ARL(\mathcal{I}^{\{y,\hat{y}\}})$
+[…]
 for multisets of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$.
-$\varphi^{rasl}$ is a special case of Definition 3 with $m_{l}=ARL$.
+$\varphi^{rasl}$ is a special case of Definition˜3 with $m_{l}=ARL$.
 The definitions from the main text can be formalized on this basis as
-$PEN(c,p):=$ $PEN_{index(c)}(sg(p)^{\{y,\hat{y}\}})$
-$ARL(p):=$ $ARL(sg(p)^{\{y,\hat{y}\}})$
+[…]
+[…]
 for patterns $p$ and instances $c$.
-$\varphi^{ROCAUC}$ and $\varphi^{PRAUC}$ are both instances of Definition 2 with the ROC AUC and the PR AUC performance measures substituted in place of $m_{s}$ respectively.
-$\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ are the same for Definition 4.
+$\varphi^{ROCAUC}$ and $\varphi^{PRAUC}$ are both instances of Definition˜2 with the ROC AUC and the PR AUC performance measures substituted in place of $m_{s}$ respectively.
+$\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ are the same for Definition˜4.
 Note that when referring to the AUC of the PR curve in this work, we mean the area under the linearly interpolated PR curve. Therefore ROC AUC and PR AUC are defined as follows.
@@
 With that we define ROC AUC and PR AUC as follows.
-Definition 5 (ROCAUC [fawcett_introduction_2006]).
+Definition 5 (ROCAUC [41]).
 The ROC AUC is defined as
@@
 The ROC AUC and PR AUC functions as they are used in the main text have different signatures. They can be formalized based on the definitions here as
-$ROCAUC(I):=$ $ROCAUC(I^{\{y,\hat{y}\}})$
-$PRAUC(I):=$ $PRAUC(I^{\{y,\hat{y}\}})$
+[…]
+[…]
 for sets of instances $I$.
@@
 They allow to structure the construction of some of the tight optimistic estimates that are presented in this section into a statement about a performance measure, considering a bound for the performance measure, and an adaptation step of that bound to interestingness measures, yielding an optimistic estimate.
-The lower bound definition (Definition 8) requires for a soft classifier performance measure an additional soft classifier performance measure that does not return smaller values than the reference performance measure for any input.
-The tightness criterion of Definition 9 requires a lower bound to be equal to the reference performance measure for at least one input.
+The lower bound definition (Definition˜8) requires for a soft classifier performance measure an additional soft classifier performance measure that does not return smaller values than the reference performance measure for any input.
+The tightness criterion of Definition˜9 requires a lower bound to be equal to the reference performance measure for at least one input.
 The performance measure lower bound definition and the bound tightness definition are defined analogously to optimistic estimates and tight optimistic estimates respectively.
@@
 A tight optimistic estimate for $\varphi^{m_{l}}$ is given by
-$oe_{\varphi^{m_{l}}}(p)=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}})$
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 $\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi^{m_{l}}(p^{\prime})=oe_{\varphi^{m_{l}}}(p).$ (2)
-Proof for Equation 1
+Proof for Equation˜1
 By definition of $\varphi^{m_{l}}$ and $oe_{\varphi^{m_{l}}}$ this equation is equivalent to
-$\forall p^{\prime}\supset p:m_{l}(sg(p^{\prime})^{\{y,\hat{y}\}})\leq b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (3)
+[…]
 The right side of this equation is constant since $D$ and $p$ are fixed.
-[lemmerich_novel_2014] states that for all two rules $p_{gen}$ and $p_{spec}$ the following statement holds
+[12] states that for all two rules $p_{gen}$ and $p_{spec}$ the following statement holds
 $p_{gen}\subset p_{spec}\Rightarrow sg(p_{gen})\supseteq sg(p_{spec}).$
@@
 holds for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
-Equation 3 follows from Equation 4 and Equation 5, concluding the proof for Equation 1.
-Proof for Equation 2
+Equation˜3 follows from Equation˜4 and Equation˜5, concluding the proof for Equation˜1.
+Proof for Equation˜2
 By definition of $\varphi^{m_{l}}$ and $oe_{\varphi^{m_{l}}}$ this equation is equivalent to
-$\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge m_{l}(sg(p^{\prime})^{\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$
+[…]
 The equivalence of $sg(p^{\prime})$ and $sg^{\prime}$ as well as the triviality of $sg^{\prime}=sg^{\prime}$ imply that the following equation is equivalent:
-$\exists sg^{\prime}\subseteq sg(p):m_{l}(sg^{\prime\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (6)
+[…]
 Because $b_{m_{l}}$ is defined to be a tight soft classifier performance measure upper bound of $m_{l}$ we know that
@@
 By fixing $C=sg(p)^{\{y,\hat{y}\}}$ we get
-$\exists C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}:m_{l}(C^{\prime})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (7)
+[…]
 The definition of databases states that for any dataset $D^{\prime}=(\mathcal{I}^{\prime},\mathcal{A}^{\prime})$ every instance in $\mathcal{I}^{\prime}$ creates exactly one element in $d(D^{\prime})$.
@@
 $\forall C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}\exists\mathcal{I}^{\prime}\subseteq sg(p):C^{\prime}=\mathcal{I^{\prime}}^{\{y,\hat{y}\}}$. (8)
-From Equation 7 and Equation 8 follows
-$\exists C^{\prime}\subseteq sg(p)^{\{y,\hat{y}\}}\exists\mathcal{I}^{\prime}\subseteq sg(p):C^{\prime}=\mathcal{I^{\prime}}^{\{y,\hat{y}\}}\wedge m_{l}(C^{\prime})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$ (9)
-The indifference of the order in a series of the same quantifiers, the equality of $C^{\prime}$ and $\mathcal{I^{\prime}}^{\{y,\hat{y}\}}$ as well as simplification rules imply that the following is equivalent to Equation 9
-$\exists\mathcal{I}^{\prime}\subseteq sg(p):m_{l}(\mathcal{I^{\prime}}^{\{y,\hat{y}\}})=b_{m_{l}}(sg(p)^{\{y,\hat{y}\}}).$
-This is trivially equivalent to Equation 6, concluding the proof for Equation 2. ∎
+From Equation˜7 and Equation˜8 follows
+[…]
+The indifference of the order in a series of the same quantifiers, the equality of $C^{\prime}$ and $\mathcal{I^{\prime}}^{\{y,\hat{y}\}}$ as well as simplification rules imply that the following is equivalent to Equation˜9
+[…]
+This is trivially equivalent to Equation˜6, concluding the proof for Equation˜2. ∎
 Lemma 8 (Result Transition from Score to Interestingness Measure).
@@
 A tight optimistic estimate for $\varphi^{m_{s}}$ is given by
-$oe_{\varphi^{m_{s}}}(p)=-b_{m_{s}}(sg(p)^{\{y,\hat{y}\}})$
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 $m_{s}^{\prime}$ is a loss-type soft classifier performance measure by definition and $b_{m_{s}^{\prime}}$ is a tight soft classifier performance measure upper bound for $m_{s}^{\prime}$.
-It follows from Lemma 7 that the soft classifier loss interestingness measure $\varphi^{m_{s}^{\prime}}$ based on $m_{s}^{\prime}$ has the following tight optimistic estimate
-$oe_{\varphi^{m_{s}^{\prime}}}(p)=b_{m_{s}^{\prime}}(sg(p)^{\{y,\hat{y}\}})$
+It follows from Lemma˜7 that the soft classifier loss interestingness measure $\varphi^{m_{s}^{\prime}}$ based on $m_{s}^{\prime}$ has the following tight optimistic estimate
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 $\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi^{\prime}(p^{\prime})=oe_{\varphi^{\prime}}(p).$ (11)
-Proof for Equation 10
+Proof for Equation˜10
 By definition of $\varphi^{\prime}$ and $oe_{\varphi^{\prime}}$ this equation is equivalent to
@@
 $\forall p^{\prime}\supset p:\varphi(p^{\prime})\leq oe_{\varphi}(p)$ (13)
-Equation 12 is true if and only if Equation 13 is true, concluding the proof for Equation 10.
-Proof for Equation 11
+Equation˜12 is true if and only if Equation˜13 is true, concluding the proof for Equation˜10.
+Proof for Equation˜11
 By definition of $\varphi^{\prime}$ and $oe_{\varphi^{\prime}}$ this equation is equivalent to
@@
 $\exists sg^{\prime}\subseteq sg(p):sg(p^{\prime})=sg^{\prime}\wedge\varphi(p^{\prime})=oe_{\varphi}(p).$ (15)
-Equation 14 is true if and only if Equation 15 is true, concluding the proof for Equation 11. ∎
-All in all by sequentially applying Lemma 7 (respectively Lemma 8) and then Lemma 9 to a soft classifier performance measure upper (respectively lower) bound for a performance measure $m$ we obtain a tight optimistic estimate for the relative soft classifier loss (respectively score) interestingness measure based on $m$.
+Equation˜14 is true if and only if Equation˜15 is true, concluding the proof for Equation˜11. ∎
+All in all by sequentially applying Lemma˜7 (respectively Lemma˜8) and then Lemma˜9 to a soft classifier performance measure upper (respectively lower) bound for a performance measure $m$ we obtain a tight optimistic estimate for the relative soft classifier loss (respectively score) interestingness measure based on $m$.
 This proving scheme is used to prove tight optimistic estimates for the relative ARL, ROC AUC and PR AUC based interestingness measures.
@@
 $\exists C^{\prime}\subseteq C:ARL(C^{\prime})=b_{ARL}(C).$ (17)
-Proof for Equation 16
+Proof for Equation˜16
 By definition of the average ranking loss (see main text) and $b_{ARL}$ this equation is equivalent to
@@
 $\forall C^{\prime}\subseteq C:\max_{i\in\{1,\dots,|C^{\prime}|\}}\{RL_{i}(C^{\prime})\}\leq\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ (19)
-It follows that the following expression implies Equation 18.
+It follows that the following expression implies Equation˜18.
 $\forall C^{\prime}\subseteq C:\frac{\sum_{i=1}^{|C^{\prime}|}RL_{i}(C^{\prime})}{\sum_{i=1}^{|C^{\prime}|}\mathbbm{1}_{\{1\}}(y^{\prime}_{i})}\leq\max_{i\in\{1,\dots,|C^{\prime}|\}}\{RL_{i}(C^{\prime})\}$ (20)
-It remains to show that Equation 20 holds.
-It is trivial that the average of a multiset $[m_{1},\dots,m_{|M|}]=M$ of real numbers $m_{i}\in\mathbb{R}$ is not greater than the maximum of $M$:
+It remains to show that Equation˜20 holds.
+It is trivial that the average of a multiset $m_{1},\dots,m_{|M|}]=M$ of real numbers $m_{i}\in\mathbb{R}$ is not greater than the maximum of $M$:
 $\frac{\sum_{i=1}^{|M|}m_{i}}{|M|}\leq\max M$ (21)
@@
 $\frac{\sum_{i=1}^{|C^{\prime}|}RL_{i}(C^{\prime})}{\sum_{i=1}^{|C^{\prime}|}\mathbbm{1}_{\{1\}}(y^{\prime}_{i})}=\frac{\sum_{i=1}^{|M_{C^{\prime}}|}m_{i}}{|M_{C^{\prime}}|}$ (22)
-Equation 20 follows from Equation 21 and Equation 22, concluding the proof for Equation 16.
-Proof for Equation 17
+Equation˜20 follows from Equation˜21 and Equation˜22, concluding the proof for Equation˜16.
+Proof for Equation˜17
 By definition of the average ranking loss and $b_{ARL}$ this equation is equivalent to
@@
 The ranking loss $RL_{i}$ for the instance at position $i$ in any given multiset of instance values $c_{j}\in\{0,1\}\times\mathbb{R}$ defines a weighted sum of the negative instances with rank at least as high as $i$ with weights that do not change for sub-multisets of the input.
-Let $i=\arg\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ and $C^{\prime}=(C\circledast\{(y,\hat{y})\in C\mid y=0\wedge\hat{y}\geq\hat{y}_{i}\})\cup[c_{i}]$.
-$C^{\prime}$ consists of the positive instance $c_{i}$ and exactly the elements of $C$ that define the non-zero summands in the definition of $PEN_{i}$ (see main text).
-By definition of the ranking loss, $C^{\prime}$ demonstrates the existence of a sub-multiset of $C$ for which Equation 23 holds.
-This concludes the proof for Equation 17. ∎
+Let $i=\arg\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ and $^{\prime}=(C\circledast\{(y,\hat{y})\in C\mid y=0\wedge\hat{y}\geq\hat{y}_{i}\})\cup[c_{i}].
+CC^{\prime}$ consists of the positive instance $c_{i}$ and exactly the elements of $C$ that define the non-zero summands in the definition of $PEN_{i}$ (see main text).
+By definition of the ranking loss, $C^{\prime}$ demonstrates the existence of a sub-multiset of $C$ for which Equation˜23 holds.
+This concludes the proof for Equation˜17. ∎
 See 1
@@
 Let $D=(\mathcal{I},\mathcal{A})$ be a dataset.
-Lemma 10 states that a tight soft classifier performance measure upper bound for the (loss-type) average ranking loss is given by the function
+Lemma˜10 states that a tight soft classifier performance measure upper bound for the (loss-type) average ranking loss is given by the function
 $b_{ARL}(C)=\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$
 for any $C=[c_{1},\dots,c_{n}]$ (multiset of instance values $c_{i}\in\{0,1\}\times\mathbb{R}$).
-It follows from Lemma 7 that a tight optimistic estimate for $\varphi^{asl}$ is given by
-$oe_{\varphi^{asl}}(p)=\max_{i\in\{1,\dots,|sg(p)|\}}\{RL_{i}(sg(p)^{\{y,\hat{y}\}})\}$
+It follows from Lemma˜7 that a tight optimistic estimate for $\varphi^{asl}$ is given by
+[…]
 for any pattern $p$ that is defined on $D$.
-$c=-ARL(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rasl}$ is given by
-$oe_{\varphi^{rasl}}(p)=\max_{i\in\{1,\dots,|sg(p)|\}}\{RL_{i}(sg(p)^{\{y,\hat{y}\}})\}-ARL(\mathcal{I}^{\{y,\hat{y}\}})$
+[…]
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 Let $C^{\prime}$ be such a sub-multiset.
-We continue with a proof by contradiction, showing that the premise $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that Equation 24 is not true.
+We continue with a proof by contradiction, showing that the premise $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that Equation˜24 is not true.
 Firstly the diagonal ROC curve with only supporting points $(0,0)$ and $(1,1)$ has an AUC of $\frac{1}{2}$.
@@
 $\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}\cup\{-\infty\}:$ $\frac{TP(C^{\prime},t)}{P(C^{\prime})}$ $<\frac{FP(C^{\prime},t)}{N(C^{\prime})}$
-$\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}:$ $\frac{1}{P(C^{\prime})}\cdot|[(y,\hat{y})$ $\in C^{\prime}\mid y=1\wedge\hat{y}>t]|<$
-$\frac{1}{N(C^{\prime})}\cdot|[(y,\hat{y})$ $\in C^{\prime}\mid y=0\wedge\hat{y}>t]|.$
+[…]
+[…]
 Because the FDR and TPR are both limited to the range $[0,1]$, it follows
@@
 $\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}:y=1\wedge y^{\prime}=0\wedge\hat{y}\leq t<\hat{y}^{\prime}$
-If there exist instances $(y,\hat{y})$ and $(y^{\prime},\hat{y}^{\prime})$ in $C^{\prime}$ that fulfill this expression, then they also fulfill the following expressions:
-$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime}$
-$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $(y<y^{\prime}\wedge\hat{y}>\hat{y}^{\prime})\vee(y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime})$
-The negation of Equation 24 is equivalent:
-$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $\neg(y\geq y^{\prime}\wedge\hat{y}\leq\hat{y}^{\prime})\vee\neg(y\leq y^{\prime}\wedge\hat{y}\geq\hat{y}^{\prime})$
-$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $\neg((y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime}))$
-$\Leftrightarrow$ $\neg\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$
-Since in case $b_{ROCAUC}(C)=\frac{1}{2}$ we know that Equation 24 holds, we can conclude that $ROCAUC(C^{\prime\prime})\geq\frac{1}{2}$ holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined).
+If there exist instances (y,y^)($\hat{y}$) and $(y^{\prime},\hat{y}^{\prime})$ in $C^{\prime}$ that fulfill this expression, then they also fulfill the following expressions:
+$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime}$
+$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $(y<y^{\prime}\wedge\hat{y}>\hat{y}^{\prime})\vee(y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime})$
+The negation of Equation˜24 is equivalent:
+$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $\neg(y\geq y^{\prime}\wedge\hat{y}\leq\hat{y}^{\prime})\vee\neg(y\leq y^{\prime}\wedge\hat{y}\geq\hat{y}^{\prime})$
+$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $\neg((y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime}))$
+$\Leftrightarrow$ $\neg\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$
+Since in case $b_{ROCAUC}(C)=\frac{1}{2}$ we know that Equation˜24 holds, we can conclude that $ROCAUC(C^{\prime\prime})\geq\frac{1}{2}$ holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined).
 Therefore $b_{ROCAUC}$ is a soft classifier performance measure lower bound for ROC AUC in this case.
-The case $b_{ROCAUC}(C)=0$ is present if and only if $C$ contains erroneous predictions, meaning that there exists at least two instances $(y,\hat{y})$ and $(y^{\prime},\hat{y}^{\prime})$ in $C$ such that $y=0$, $y^{\prime}=1$ and $\hat{y}>\hat{y}^{\prime}$.
-The sub-multiset of $C$ that only contains these two instances (once each) has an ROC AUC of 00.
+The case $b_{ROCAUC}(C)=0$ is present if and only if $C$ contains erroneous predictions, meaning that there exists at least two instances (y,y^)($\hat{y}$) and $(y^{\prime},\hat{y}^{\prime})$ in $C$ such that $y=0$, $y^{\prime}=1$ and $\hat{y}>\hat{y}^{\prime}$.
+The sub-multiset of $C$ that only contains these two instances (once each) has an ROC AUC of 0.
 This is the overall minimum of the ROC AUC.
@@
 Let $D=(\mathcal{I},\mathcal{A})$ be a dataset.
-Lemma 2 states that a tight soft classifier performance measure lower bound for the (score-type) area under the ROC curve is given by $b_{ROCAUC}$.
-It follows from Lemma 8 that a tight optimistic estimate for $\varphi^{ROCAUC}$ is given by
-$oe_{\varphi^{ROCAUC}}(p)=-b_{ROCAUC}(sg(p)^{\{y,\hat{y}\}})$
+Lemma˜2 states that a tight soft classifier performance measure lower bound for the (score-type) area under the ROC curve is given by $b_{ROCAUC}$.
+It follows from Lemma˜8 that a tight optimistic estimate for $\varphi^{ROCAUC}$ is given by
+[…]
 for any pattern $p$ that is defined on $D$.
-$c=ROCAUC(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rROCAUC}$ is given by
-$oe_{\varphi^{rROCAUC}}=ROCAUC(\mathcal{I}^{\{y,\hat{y}\}})-b_{ROCAUC}(sg(p)^{\{y,\hat{y}\}})$
+[…]
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 A.3 Area Under the Precision-Recall Curve
-000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111recallprecisionborderapproximation 1approximation 2
+00.20.20.40.40.60.60.80.81100.20.20.40.40.60.60.80.811recallprecisionborderapproximation 1approximation 2
 Figure 5: Lower border of the achievable area in PR space and approximations for a positive class ratio $\frac{1}{2}$.
@@
 Define
-$T(C)$ :=$:=\{(-\infty,\hat{y}_{1})\}\cup\{(\hat{y}_{i-1},\hat{y}_{i})\mid i\in\{2,\dots,|C|\}\wedge\hat{y}_{i-1}\neq\hat{y}_{i}\}$
-$T_{<\hat{y}_{minpos}}(C)$ :=$:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}<\hat{y}_{minpos}\}$
-$T_{\geq\hat{y}_{minpos}}(C)$ :=$:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}\geq\hat{y}_{minpos}\}$
+$T(C)$ $:=\{(-\infty,\hat{y}_{1})\}\cup\{(\hat{y}_{i-1},\hat{y}_{i})\mid i\in\{2,\dots,|C|\}\wedge\hat{y}_{i-1}\neq\hat{y}_{i}\}$
+$T_{<\hat{y}_{minpos}}(C)$ $:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}<\hat{y}_{minpos}\}$
+$T_{\geq\hat{y}_{minpos}}(C)$ $:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}\geq\hat{y}_{minpos}\}$
 From the definition of $\hat{y}_{minpos}$ follows
-$\forall t\in\{\hat{y}\mid(y,\hat{y})\in C\}:$ $t<\hat{y}_{minpos}$
+$\forall t\in\{\hat{y}\mid(y,\hat{y})\in C\}:\$$t<\hat{y}_{minpos}$
 $\Rightarrow TP(C,t)=P(C)$
@@
 Define $\mathcal{C}_{\hat{y}\subseteq\hat{y}_{minpos}}$ similarly for any multiset $\mathcal{C}$ of instance values $(y,\hat{y})\in\{0,1\}\times\mathbb{R}$ that contains at least one instance with $y=1$.
-In particular, Equation 30 also holds for any sub-multiset $C^{\prime}$ of $C$ with corresponding $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$. Therefore
+In particular, Equation˜30 also holds for any sub-multiset $C^{\prime}$ of $C$ with corresponding $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$. Therefore
 $\forall C^{\prime}\subseteq C:PRAUC(C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime})=PRAUC(C^{\prime})$
@@
 $\left\{(0,1),(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
-=$=\left\{(0,1),(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
+$=\left\{(0,1),(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
 Let
-$\mathfrak{P}_{worst}^{\prime}$ =$=\left\{(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
-=$=\left\{(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
-Clearly the AUC between the points $(0,1)$ and $(0,0)$ is 00, so
+$\mathfrak{P}_{worst}^{\prime}$ $=\left\{(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
+$=\left\{(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
+Clearly the AUC between the points $(0,1)$ and $(0,0)$ is 0, so
 $PRAUC(C_{worst})=AUC(\mathfrak{P}_{worst}^{\prime}).$
-$\mathfrak{P}_{worst}^{\prime}$ are the supporting points of the minimum PR curve described in [boyd_unachievable_2012].
+$\mathfrak{P}_{worst}^{\prime}$ are the supporting points of the minimum PR curve described in [16].
 As $\mathfrak{P}_{worst}^{\prime}$ shows, the minimum PR curve only depends on the positive class ratio.
@@
 Also no point on the linear interpolation of $\mathfrak{P}_{worst}^{\prime}$ is strictly above the correct PR interpolation of $\mathfrak{P}_{worst}^{\prime}$ as shown in the following.
-The points on the correct interpolation between the two points in $\mathfrak{P}_{worst}^{\prime}$ according to [boyd_unachievable_2012] are
+The points on the correct interpolation between the two points in $\mathfrak{P}_{worst}^{\prime}$ according to [16] are
 $\left(r_{i},\frac{\pi r_{i}}{(1-\pi)+\pi r_{i}}\right)$
@@
 $=$ $\frac{1+N(C_{worst})+N(C_{worst})^{2}}{1+2N(C_{worst})+N(C_{worst})^{2}}$
-$\leq$ $1$
+$\leq$ $1.$
 Therefore
@@
 Let $D=(\mathcal{I},\mathcal{A})$ be a dataset.
-Lemma 4 states that a tight soft classifier performance measure lower bound for the (score-type) area under the PR curve is given by $b_{PRAUC}$.
-It follows from Lemma 8 that a tight optimistic estimate for $\varphi^{PRAUC}$ is given by
-$oe_{\varphi^{PRAUC}}(p)=-b_{PRAUC}(sg(p)^{\{y,\hat{y}\}})$
+Lemma˜4 states that a tight soft classifier performance measure lower bound for the (score-type) area under the PR curve is given by $b_{PRAUC}$.
+It follows from Lemma˜8 that a tight optimistic estimate for $\varphi^{PRAUC}$ is given by
+[…]
 for any pattern $p$ that is defined on $D$.
-$c=PRAUC(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rPRAUC}$ is given by
-$oe_{\varphi^{rPRAUC}}=PRAUC(\mathcal{I}^{\{y,\hat{y}\}})-b_{PRAUC}(sg(p)^{\{y,\hat{y}\}})$
+[…]
+[…]
 for any pattern $p$ that is defined on $D$.
@@
 $\Leftrightarrow$ $\frac{P(C)-1+N(C)}{P(C)+N(C)}$ $\geq\left(\frac{N(C)}{P(C)}\right)^{\frac{\beta}{\alpha}}\cdot\left(\frac{P(C)-1}{N(C)}\right)^{\frac{\beta}{\alpha}}$
-$\Leftrightarrow$ 11 11 $\geq\left(1-\frac{1}{P(C)}\right)^{\frac{\beta}{\alpha}}+\frac{1}{P(C)+N(C)}$ (32)
+$\Leftrightarrow$ $1$ $\geq\left(1-\frac{1}{P(C)}\right)^{\frac{\beta}{\alpha}}+\frac{1}{P(C)+N(C)}$ (32)
 In case $\alpha=\beta$ follows
-11 11 $\geq\left(1-\frac{1}{P(C)}\right)^{1}+\frac{1}{P(C)+N(C)}$
+$1$ $\geq\left(1-\frac{1}{P(C)}\right)^{1}+\frac{1}{P(C)+N(C)}$
 $\Leftrightarrow$ $N(C)$ $\geq 0$
@@
 for any constant $c\geq 1$.
-Therefore Equation 32 holding in case $\alpha=\beta$ implies that it also holds in case $\alpha\leq\beta$. We conclude that the weighting $w$ does not decrease when $c$ is removed from $C$.
+Therefore Equation˜32 holding in case $\alpha=\beta$ implies that it also holds in case $\alpha\leq\beta$. We conclude that the weighting $w$ does not decrease when $c$ is removed from $C$.
 Note that the case $P(C)=N(C)+1$ works out here even though $cb(C\ominus[c])$ is not defined by the first case of the minimum in the $cb$ definition anymore.
@@
 $w(C^{\prime})=(2\cdot\min\{P(C),N(C)\})^{\alpha}$
-Assume $N(C)=0\vee P(C)=0$ (Equal to the case in the $cb$ definition returning 00).
+Assume $N(C)=0\vee P(C)=0$ (Equal to the case in the $cb$ definition returning 0).
 In this case
@@
 $\forall C^{\prime}\subseteq C:w(C^{\prime})=|C^{\prime}|^{\alpha}\cdot 0^{\beta}=0.$
-$b_{w}$ is a tight upper bound in this case since it also returns 00:
+$b_{w}$ is a tight upper bound in this case since it also returns 0:
 $b_{w}(C)=(2\cdot\min\{P(C),N(C)\})^{\alpha}=(2\cdot 0)^{\alpha}=0$
@@
 B.2 Quality Parameters of Results Across Multiple Datasets
-Table 4: Overview of result set properties for combinations of datasets and performance measures for an xgboost classifier. (Continuation of Table 1 in the main text)
-Mean Exceptionality
+Table 5: Overview of result set properties for combinations of datasets and performance measures for an xgboost classifier. (Continuation of Table˜1 in the main text)
+Mean Exception- ality
 Mean Cover Size
@@
 Significant/ Top-5
-Top-5/ Significant/ Original
-Dataset Metric     b f     b f     b f     b f     b f
-Mushroom n=8124 NCR=0.52 ROC AUC     0.00 -    203 -    0.76 -    0.28 -    0/5 0/0/100
-PR AUC     0.00 0.00    35 57    0.17 0.49    0.14 -    0/5 1/1/100
+Filtered-5/ Significant/ Top-100
+Dataset
+Metric
+b
+f
+b
+f
+b
+f
+b
+f
+b
+f
+Mushroom n=8124 NCR=0.52 ROC     0.00 -    203 -    0.76 -    0.28 -    0/5 0/0/100
+PR     0.00 0.00    35 57    0.17 0.49    0.14 -    0/5 1/1/100
 ARL     - -    - -    - -    - -    0/0 0/0/0
-Statlog n=1000 NCR=0.30 ROC AUC     0.74 -    23 -    0.04 -    0.36 -    0/5 0/0/100
-PR AUC     0.50 -    21 -    0.63 -    0.95 -    0/5 0/0/100
+Statlog n=1000 NCR=0.30 ROC     0.74 -    23 -    0.04 -    0.36 -    0/5 0/0/100
+PR     0.50 -    21 -    0.63 -    0.95 -    0/5 0/0/100
 ARL     0.16 -    241 -    0.28 -    - -    0/1 0/0/1
-Cancer n=699 NCR=0.66 ROC AUC     0.03 -    41 -    1.00 -    0.17 -    0/5 0/0/5
-PR AUC     0.55 -    78 -    1.00 -    0.80 -    0/5 0/0/7
+Cancer n=699 NCR=0.66 ROC     0.03 -    41 -    1.00 -    0.17 -    0/5 0/0/5
+PR     0.55 -    78 -    1.00 -    0.80 -    0/5 0/0/7
 ARL     0.18 -    83 -    1.00 -    - -    0/1 0/0/1
@@
 Result Set Size (Top-5/ Significant/ Original)
-Dataset Interestingness Measure
+Dataset
+Interestingness
+Measure
 UCI Census-Income (KDD) ROC AUC 0.95 0.46 33 0.03 5/11/100
@@
 B.3 Case Study: Adult
-The following are full result sets of the searches on the Adult dataset, that are also reported in Table 1 in the main text in terms of aggregated statistics.
+The following are full result sets of the searches on the Adult dataset, that are also reported in Table˜1 in the main text in terms of aggregated statistics.
 Note in particular the amount of overlap between patterns in the result sets from the base SubROC setting compared to the subgroups in the results of the full SubROC framework.
@@
 Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
-[…]
-[…]
-[…]
-[…]
-[…]
-$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
+159.2829
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{capital-loss}\in[0.0,213.0)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}$
+372.3026
+0.8409
+0.7780
+4309
+0.3892
+157.6086
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}$
+370.6283
+0.8473
+0.8053
+4628
+0.4114
+142.0611
+$\text{capital-loss}\in[0.0,213.0)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}$
+355.0808
+0.8611
+0.8346
+4961
+0.4328
+141.2313
+$\text{marital-status}=\text{Married-civ-spouse}$
+354.2510
+0.8660
+0.8508
+5280
+0.4496
+138.0929
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{capital-loss}\in[0.0,213.0)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}\>\wedge$
+$\text{native-country}=\text{United-States}$
+351.1126
+0.8275
+0.7748
+3894
+0.4004
+0.0000
+$\emptyset$
+213.0197
+0.9728
+0.9241
+11305
+0.2482
 Table 8: Result set of basic SubROC setting with ROC AUC.
 Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
-[…]
-[…]
-[…]
-[…]
-[…]
-$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
+0.9241
+$\text{education}=\text{Some-college}\>\wedge$
+$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
+$\text{marital-status}=\text{Divorced}\>\wedge$
+$\text{workclass}=\text{Private}$
+1.0000
+0.8439
+0.0000
+21
+0.0476
+0.9241
+$\text{education}=\text{Some-college}\>\wedge$
+$\text{fnlwgt}\in[178811.0,196338.0)\>\wedge$
+$\text{relationship}=\text{Not-in-family}\>\wedge$
+$\text{sex}=\text{Male}$
+1.0000
+0.8857
+0.0000
+32
+0.0312
+0.9241
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{education}=\text{Some-college}\>\wedge$
+$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
+$\text{marital-status}=\text{Divorced}$
+1.0000
+0.8770
+0.0000
+29
+0.0345
+0.9241
+$\text{education-num}\in[10,11)\>\wedge$
+$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
+$\text{marital-status}=\text{Divorced}\>\wedge$
+$\text{race}=\text{White}$
+1.0000
+0.8703
+0.0000
+27
+0.0370
+0.9241
+$\text{education-num}\in[10,11)\>\wedge$
+$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
+$\text{marital-status}=\text{Divorced}\>\wedge$
+$\text{workclass}=\text{Private}$
+1.0000
+0.8439
+0.0000
+21
+0.0476
+0.0000
+$\emptyset$
+213.0197
+0.9728
+0.9241
+11305
+0.2482
 Table 9: Result set of basic SubROC setting with PR AUC.
 Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
-[…]
-[…]
-[…]
-[…]
-[…]
-$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
+0.9395
+$\text{education}=\text{Bachelors}\>\wedge$
+$\text{hours-per-week}\in[44,50)\>\wedge$
+$\text{occupation}=\text{Prof-specialty}\>\wedge$
+$\text{relationship}=\text{Husband}$
+14.0000
+0.0333
+0.4400
+26
+0.9615
+0.9395
+$\text{education-num}\in[13,14)\>\wedge$
+$\text{hours-per-week}\in[44,50)\>\wedge$
+$\text{occupation}=\text{Prof-specialty}\>\wedge$
+$\text{relationship}=\text{Husband}$
+14.0000
+0.0333
+0.4400
+26
+0.9615
+0.9371
+$\text{education-num}\in[14,15)\>\wedge$
+$\text{hours-per-week}\in[50,55)\>\wedge$
+$\text{occupation}=\text{Exec-managerial}\>\wedge$
+$\text{sex}=\text{Male}$
+13.0000
+0.0357
+0.5357
+29
+0.9655
+0.9371
+$\text{education}=\text{Masters}\>\wedge$
+$\text{hours-per-week}\in[50,55)\>\wedge$
+$\text{occupation}=\text{Exec-managerial}\>\wedge$
+$\text{sex}=\text{Male}$
+13.0000
+0.0357
+0.5357
+29
+0.9655
+0.9311
+$\text{education}=\text{Masters}\>\wedge$
+$\text{hours-per-week}\in[50,55)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}\>\wedge$
+$\text{occupation}=\text{Exec-managerial}$
+11.0000
+0.0417
+0.5600
+26
+0.9615
+0.0000
+$\emptyset$
+213.0197
+0.9728
+0.9241
+11305
+0.2482
 Table 10: Result set of full SubROC framework with ARL.
 Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
-[…]
-[…]
-$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
-Table 10: Result set of full SubROC framework with ROC AUC.
+609186.2021
+$\text{marital-status}=\text{Married-civ-spouse}$
+354.2510
+0.8660
+0.8508
+5280
+0.4496
+389481.0168
+$\text{relationship}=\text{Husband}$
+315.8451
+0.8619
+0.8485
+4638
+0.4495
+34597.0291
+$\text{sex}=\text{Male}$
+223.1345
+0.9541
+0.9053
+7581
+0.3109
+9394.5568
+$\text{capital-gain}\in[0.0,114.0)$
+216.3884
+0.9711
+0.9015
+10331
+0.2126
+4270.2607
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{sex}=\text{Male}$
+228.3650
+0.9505
+0.8764
+6821
+0.2708
+0.0000
+$\emptyset$
+213.0197
+0.9728
+0.9241
+11305
+0.2482
+Table 11: Result set of full SubROC framework with ROC AUC. (Extended version of Table˜2 in the main text)
 Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
-[…]
-[…]
-[…]
-[…]
-[…]
-$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
+316.1949
+$\text{marital-status}=\text{Married-civ-spouse}$
+354.2510
+0.8660
+0.8508
+5280
+0.4496
+286.2383
+$\text{relationship}=\text{Husband}$
+315.8451
+0.8619
+0.8485
+4638
+0.4495
+67.9166
+$\text{capital-gain}\in[0.0,114.0)\>\wedge$
+$\text{marital-status}=\text{Married-civ-spouse}$
+370.6283
+0.8473
+0.8053
+4628
+0.4114
+64.1438
+$\text{sex}=\text{Male}$
+223.1345
+0.9541
+0.9053
+7581
+0.3109
+63.0874
+$\text{capital-gain}\in[0.0,114.0)$
+216.3884
+0.9711
+0.9015
+10331
+0.2126
+0.0000
+$\emptyset$
+213.0197
+0.9728
+0.9241
+11305
+0.2482
 Table 12: Result set of full SubROC framework with PR AUC.
 Interestingness Pattern ARL PR AUC ROC AUC Cover NCR
-[…]
-[…]
-[…]
-[…]
-[…]
-$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
+460.8622
+$\text{marital-status}=\text{Married-civ-spouse}$
+354.2510
+0.8660
+0.8508
+5280
+0.4496
+420.2511
+$\text{relationship}=\text{Husband}$
+315.8451
+0.8619
+0.8485
+4638
+0.4495
+87.2251
+$\text{occupation}=\text{Exec-managerial}$
+71.7942
+0.9121
+0.9027
+1518
+0.4862
+79.5800
+$\text{occupation}=\text{Prof-specialty}$
+69.6873
+0.9104
+0.8978
+1467
+0.4649
+74.2654
+$\text{hours-per-week}\in[50,55)$
+58.9793
+0.8939
+0.8844
+1113
+0.4582
+0.0000
+$\emptyset$
+213.0197
+0.9728
+0.9241
+11305
+0.2482
 B.4 Searching for an Injected Subgroup
-For details on the experimental setup, see Section 4.3 in the main text.
+For details on the experimental setup, see Section˜4.3 in the main text.
 Figure 6: IoU of an injected subgroup with subgroups in a top-10 ARL result set, ordered decreasingly by interestingness.
@@
 B.5 Optimistic Estimates
-Table 12: Speedups of median runtimes using optimistic estimates. (Continuation of Table 2 in the main text)
+Table 13: Speedups of median runtimes using optimistic estimates. (Continuation of Table˜3 in the main text)
 Performance Measure
@@
 Cancer n/m=699/9
-ARL 0    381.7 2.6 2.8
-0.1    292.1 1.7 1.5
-0.3    283.2 1.7 1.1
-1    280.6 1.6 1.2
-PR AUC 0    11.8 0.6 2.2
-0.1    5.2 0.5 0.6
-0.3    5.2 0.5 0.6
-1    5.2 0.5 0.6
-ROC AUC 0    29.8 0.8 0.8
-0.1    3.6 0.6 0.5
-0.3    3.7 0.6 0.5
-1    3.6 0.7 0.5
-Table 13: Median runtimes in seconds for ARL with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
-Subgroups were required to cover at least 20 instances and the search stopped at depth 3.
-10 measurements were performed for each configuration.
-0 0.1 0.3 1
-Dataset Off On Off On Off On Off On
-UCI Census-Income (KDD) 6899 723.8 6795 712.9 6877 667.3 6872 592.3
-OpenML Adult 177.7 11.6 178.0 13.9 177.9 12.9 178.8 11.5
-UCI Bank Marketing 28.8 5.9 28.9 6.9 29.0 6.9 29.2 6.8
-UCI Credit Card Clients 523.3 7.5 525.2 10.1 530.3 10.0 531.4 10.0
-UCI Mushroom 17.7 0.2 17.8 0.3 17.9 0.3 17.8 0.3
-Statlog (German Credit Data) 3.3 1.6 3.3 2.2 3.3 2.2 3.3 2.2
-UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.1 0.2 0.1 0.1 0.1
-UCI Credit Approval 0.7 0.5 0.7 0.9 0.7 1.0 0.7 1.0
-Table 14: Median runtimes in seconds for ROC AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
-Subgroups were required to cover at least 20 instances and the search stopped at depth 3.
-10 measurements were performed for each configuration.
-0 0.1 0.3 1
-Dataset Off On Off On Off On Off On
-UCI Census-Income (KDD) 5469 4785 5499 5391 5609 4814 5477 1429
-OpenML Adult 187.2 192.5 187.8 197.7 186.4 182.9 187.6 44.1
-UCI Bank Marketing 39.6 21.2 39.8 43.3 39.9 40.6 40.0 26.6
-UCI Credit Card Clients 633.1 18.8 647.0 658.6 641.0 633.9 645.6 367.6
-UCI Mushroom 20.5 1.0 20.7 7.7 20.6 7.6 20.8 7.7
-Statlog (German Credit Data) 5.7 6.1 5.6 7.0 5.6 7.0 5.6 6.9
-UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.3 0.2 0.3 0.2 0.3
-UCI Credit Approval 1.0 1.2 1.0 1.6 1.0 1.6 1.0 1.6
-Table 15: Median runtimes in seconds for PR AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
-Subgroups were required to cover at least 20 instances and the search stopped at depth 3.
-10 measurements were performed for each configuration.
-0 0.1 0.3 1
-Dataset Off On Off On Off On Off On
-UCI Census-Income (KDD) 7007 3645 6923 4149 7014 3502 7147 1522
-OpenML Adult 192.4 121.7 194.2 173.6 192.2 114.4 191.8 23.9
-UCI Bank Marketing 35.8 28.2 35.9 39.5 35.5 37.1 35.6 23.2
-UCI Credit Card Clients 595.4 457.1 599.8 578.2 598.9 405.8 597.4 110.5
-UCI Mushroom 24.2 3.3 24.3 7.4 24.3 7.4 24.3 7.2
-Statlog (German Credit Data) 4.7 6.0 4.7 7.1 4.8 7.1 4.8 6.7
-UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.3 0.2 0.3 0.2 0.3
-UCI Credit Approval 0.9 0.9 0.9 1.6 0.9 1.7 0.9 1.7
-Table 16: Speedups of median runtimes reported in Tables 13, 14 and 15.
-Performance Measure
-Weight ($\alpha=\beta$)
+ARL 0    617.8 132.4 5.0
+0.1    410.1 85.5 2.8
+0.3    405.6 84.2 2.6
+1    403.0 85.6 2.7
+PR AUC 0    699.8 84.4 4.8
+0.1    211.9 62.3 2.9
+0.3    89.7 85.9 2.9
+1    78.6 85.7 2.9
+ROC AUC 0    337.7 87.2 8.3
+0.1    268.6 54.8 5.0
+0.3    277.1 53.4 4.3
+1    249.6 68.9 4.3
+Table 14: Median runtimes in seconds for ARL without optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
+Subgroups were required to cover at least 20 instances and the search stopped at depth 4.
+10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
+Dataset 0 0.1 0.3 1
+UCI Census-Income (KDD) 114771.7 115310.2 114248.1 114822.9
+OpenML Adult 1058.0 1056.0 1044.8 1061.8
+UCI Bank Marketing 127.8 128.6 128.7 128.9
+UCI Credit Card Clients 4826.9 4803.8 4867.5 4800.9
+UCI Mushroom 119.2 119.4 118.5 119.1
+Statlog (German Credit Data) 10.6 10.7 10.6 10.8
+UCI Breast Cancer Wisconsin 0.1 0.1 0.1 0.1
+UCI Credit Approval 1.9 1.8 1.8 1.9
+Table 15: Median runtimes in seconds for ROC AUC without optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
+Subgroups were required to cover at least 20 instances and the search stopped at depth 4.
+10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
+Dataset 0 0.1 0.3 1
+UCI Census-Income (KDD) 84443.9 85060.0 85223.8 84915.9
+OpenML Adult 1055.2 1048.9 1036.1 1051.5
+UCI Bank Marketing 160.0 159.6 160.2 161.4
+UCI Credit Card Clients 5304.0 5305.0 5293.7 5275.4
+UCI Mushroom 118.4 117.7 117.4 118.9
+Statlog (German Credit Data) 16.2 16.5 16.4 16.5
+UCI Breast Cancer Wisconsin 0.3 0.3 0.3 0.3
+UCI Credit Approval 2.5 2.4 2.4 2.4
+Table 16: Median runtimes in seconds for PR AUC without optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
+Subgroups were required to cover at least 20 instances and the search stopped at depth 4.
+10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
+Dataset 0 0.1 0.3 1
+UCI Census-Income (KDD) 118601.6 118895.8 118370.9 118797.9
+OpenML Adult 1123.3 1134.3 1127.3 1124.1
+UCI Bank Marketing 145.4 146.4 147.7 146.9
+UCI Credit Card Clients 5139.3 5138.0 5117.3 5122.6
+UCI Mushroom 161.0 162.8 162.8 162.6
+Statlog (German Credit Data) 14.2 14.8 14.2 14.1
+UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.2
+UCI Credit Approval 2.2 2.2 2.2 2.2
+Table 17: Share of remaining evaluated subgroups in the same searches reported in Tables˜14, 15 and 16
 UCI Census-Income (KDD)
@@
 UCI Credit Approval
-ARL 0 9.5 15.3 4.9 69.4 98.9 2.1 2.5 1.4
-0.1 9.5 12.8 4.2 52.2 64.3 1.5 1.6 0.8
-0.3 10.3 13.8 4.2 52.8 64.4 1.5 1.1 0.7
-1 11.6 15.6 4.3 52.9 64.8 1.5 1.1 0.7
-PR AUC 0 1.9 1.6 1.3 1.3 7.3 0.8 1.7 1.0
-0.1 1.7 1.1 0.9 1.0 3.3 0.7 0.6 0.6
-0.3 2.0 1.7 1.0 1.5 3.3 0.7 0.6 0.6
-1 4.7 8.0 1.5 5.4 3.4 0.7 0.5 0.5
-ROC AUC 0 1.1 1.0 1.9 33.7 21.6 0.9 0.9 0.9
-0.1 1.0 0.9 0.9 1.0 2.7 0.8 0.6 0.6
-0.3 1.2 1.0 1.0 1.0 2.7 0.8 0.6 0.6
-1 3.8 4.3 1.5 1.8 2.7 0.8 0.6 0.7
-Table 17: Share of remaining evaluated subgroups in the same searches reported in Tables 13, 14 and 15
-Performance Measure
-Weight ($\alpha=\beta$)
-UCI Census-Income (KDD)
-OpenML Adult
-UCI Bank Marketing
-UCI Credit Card Clients
-UCI Mushroom
-Statlog (German Credit Data)
-UCI Breast Cancer Wisconsin
-UCI Credit Approval
-ARL 0 4.51% 2.47% 6.34% 0.41% 0.21% 28.23% 20.76% 48.52%
-0.1 4.10% 2.37% 6.34% 0.41% 0.21% 28.23% 20.76% 53.70%
-0.3 3.74% 2.13% 6.34% 0.41% 0.21% 28.23% 20.76% 61.64%
-1 3.12% 1.81% 6.34% 0.41% 0.21% 28.23% 20.76% 67.35%
-PR AUC 0 42.38% 48.05% 45.50% 59.99% 8.48% 100.00% 20.76% 51.82%
-0.1 48.67% 76.41% 89.06% 87.05% 16.68% 100.00% 51.18% 88.98%
-0.3 39.40% 42.41% 75.48% 47.02% 16.70% 100.00% 51.18% 94.74%
-1 11.98% 4.73% 24.25% 7.79% 15.56% 89.35% 62.44% 91.12%
-ROC AUC 0 83.00% 97.49% 47.80% 3.10% 2.74% 100.00% 71.65% 94.74%
-0.1 91.62% 97.49% 98.05% 99.03% 22.29% 100.00% 71.65% 94.74%
-0.3 79.95% 87.01% 82.80% 94.49% 22.29% 100.00% 71.65% 94.74%
-1 13.92% 11.33% 28.98% 36.66% 22.29% 94.27% 71.65% 92.11%
-Figure 7: Histogram of standard deviations divided by medians for all runtimes of Tables 13, 14 and 15.
+Performance
+Measure Weight ($\alpha=\beta$)
+ARL 0 6.62% 3.82% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
+0.1 6.07% 3.48% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
+0.3 5.20% 2.82% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
+1 3.21% 1.07% 0.02% 4.55% 0.03% 0.13% 6.68% 0.48%
+PR AUC 0 42.47% 47.13% 0.02% 17.65% 0.03% 0.13% 6.68% 0.48%
+0.1 42.47% 60.95% 0.02% 42.01% 0.07% 0.13% 6.68% 0.48%
+0.3 42.47% 63.18% 0.02% 42.77% 0.20% 0.13% 6.68% 0.48%
+1 19.63% 7.85% 0.02% 25.62% 0.25% 0.13% 6.68% 0.48%
+ROC AUC 0 61.26% 73.32% 0.02% 42.99% 0.04% 0.13% 7.15% 0.48%
+0.1 61.26% 73.32% 0.02% 42.99% 0.04% 0.13% 7.15% 0.48%
+0.3 61.29% 73.32% 0.02% 42.99% 0.04% 0.13% 7.15% 0.48%
+1 26.83% 9.89% 0.02% 36.11% 0.04% 0.13% 7.15% 0.48%
+Figure 8: Histogram of standard deviations divided by medians for all runtimes of Tables˜14, 15 and 16.
 Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.
-Table 18: Median runtimes in seconds for ARL with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
-Subgroups were required to cover at least 20 instances and the search stopped at depth 4.
-10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
-0 0.1 0.3 1
-Dataset Off On Off On Off On Off On
-UCI Census-Income (KDD) 98310 5346 98206 5265 98876 4958 98284 4223
-OpenML Adult 978.1 20.6 995.1 25.2 969.7 24.0 983.5 22.0
-UCI Bank Marketing 119.5 12.4 119.2 15.5 119.9 15.7 119.4 15.6
-UCI Credit Card Clients 4510 8.4 4479 11.6 4449 11.6 4491 11.6
-UCI Mushroom 109.5 0.3 110.8 0.4 109.8 0.4 109.9 0.4
-Statlog (German Credit Data) 9.8 3.8 10.0 5.9 9.9 5.9 9.8 6.0
-UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.2 0.2 0.2 0.1 0.1
-UCI Credit Approval 1.7 1.0 1.8 2.2 1.8 2.5 1.8 2.8
-Table 19: Median runtimes in seconds for ROC AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
-Subgroups were required to cover at least 20 instances and the search stopped at depth 4.
-10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
-0 0.1 0.3 1
-Dataset Off On Off On Off On Off On
-UCI Census-Income (KDD) 73835 731.2 72673 73204 74144 57059 74759 11694
-OpenML Adult 961.8 1076 965.9 1138 971.8 1014 976.0 116.3
-UCI Bank Marketing 150.1 69.9 150.5 199.2 149.8 170.8 148.0 72.0
-UCI Credit Card Clients 4907 58.6 5000 5256 4954 4569 4967 1267
-UCI Mushroom 107.2 3.6 107.6 29.7 108.3 29.6 106.9 29.9
-Statlog (German Credit Data) 15.4 18.2 15.4 25.0 15.2 25.2 15.3 23.4
-UCI Breast Cancer Wisconsin 0.2 0.2 0.2 0.4 0.2 0.4 0.2 0.4
-UCI Credit Approval 2.3 3.1 2.3 4.8 2.3 4.8 2.3 4.7
-Table 20: Median runtimes in seconds for PR AUC with (“On”) versus without (“Off”) optimistic estimate pruning, with equal weighting parameters $\alpha$ and $\beta$ in $\{0,0.1,0.3,1\}$.
-Subgroups were required to cover at least 20 instances and the search stopped at depth 4.
-10 measurements were performed for each configuration, except for the dataset “UCI Census-Income (KDD)” for which it were 3 measurements.
-0 0.1 0.3 1
-Dataset Off On Off On Off On Off On
-UCI Census-Income (KDD) 99395 34225 101272 52252 101602 37638 101372 8896
-OpenML Adult 1044 444.9 1060 917.9 1032 483.9 1031 52.0
-UCI Bank Marketing 137.4 79.2 137.4 189.0 137.7 162.6 138.8 62.3
-UCI Credit Card Clients 4779 1899 4823 4199 4842 1681 4731 209.3
-UCI Mushroom 146.4 12.5 149.2 29.0 149.0 28.9 148.7 28.5
-Statlog (German Credit Data) 13.2 20.8 13.3 28.1 13.3 27.9 13.4 25.3
-UCI Breast Cancer Wisconsin 0.2 0.1 0.2 0.3 0.2 0.3 0.2 0.4
-UCI Credit Approval 2.1 1.8 2.1 5.1 2.1 5.4 2.1 5.2
-Table 21: Share of remaining evaluated subgroups in the same searches reported in Tables 18, 19 and 20
-Performance Measure
-Weight ($\alpha=\beta$)
-UCI Census-Income (KDD)
-OpenML Adult
-UCI Bank Marketing
-UCI Credit Card Clients
-UCI Mushroom
-Statlog (German Credit Data)
-UCI Breast Cancer Wisconsin
-UCI Credit Approval
-ARL 0 2.39% 0.68% 2.66% 0.04% 0.03% 19.99% 18.97% 27.64%
-0.1 2.17% 0.66% 2.66% 0.04% 0.03% 19.99% 18.97% 32.30%
-0.3 1.97% 0.61% 2.66% 0.04% 0.03% 19.99% 18.97% 39.50%
-1 1.54% 0.53% 2.66% 0.04% 0.03% 19.99% 18.97% 47.98%
-PR AUC 0 25.80% 27.77% 22.54% 22.70% 5.31% 99.60% 18.97% 30.55%
-0.1 42.09% 66.38% 84.34% 65.46% 10.43% 99.60% 47.47% 83.66%
-0.3 27.81% 28.28% 65.27% 17.37% 10.49% 98.95% 47.47% 92.61%
-1 4.32% 1.57% 11.72% 1.23% 9.85% 87.02% 57.76% 88.14%
-ROC AUC 0 0.99% 96.32% 38.06% 1.05% 2.10% 99.60% 68.17% 93.91%
-0.1 90.86% 96.32% 97.18% 92.96% 16.28% 99.60% 68.17% 93.91%
-0.3 67.43% 83.35% 74.42% 75.78% 16.28% 99.60% 68.17% 93.91%
-1 8.41% 4.76% 15.70% 11.74% 16.28% 92.08% 68.17% 89.20%
-Figure 8: Histogram of standard deviations divided by medians for all runtimes of Tables 18, 19 and 20.
-Most series of measurements did not experience noticeable runtime fluctuations compared to their median runtime.
-For the speedups of median runtimes reported in Tables 18, 19 and 20, see Table 2 in the main text.